A Complete Bibliography of Publications in Biometrika for the decade 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

& [Sun02]. α [KM09, WJ00]. χ^{2} [WW06]. D [JG09, LM09]. G [Moo09]. G_{2} [Tan01]. J [Tan01]. K [SR01a, YSZ06, Zha06]. M [CT06, DM01, DM02a, Ste06]. P [AEB07, FR08, GRW06, PWB08, TG07]. R_{∗} [BBL04]. t [AR02, DJ00, FFM08, GC04, HSW00, KCD08, WCK09]. U [SR01a].

-Aberration [Tan01]. -Characteristics [Tan01]. -designs [WJ00]. -Estimands [DM02a]. -Estimation [DM01]. -estimators [Moo09, Ste06]. -link [KCD08]. -mixture [WCK09]. -optimal [JG09, LM09]. -permanent [KM09]. -quantile [CT06]. -Sample [SR01a, YSZ06, Zha06]. -Statistic [SR01a]. -test [DJ00]. -tests [AR02]. -type [HSW00]. -value [GRW06]. -Values [FR08, AEB07, PWB08, TG07]. -variates [GC04].

84 [MO05, WO04]. 86 [Kos01a]. 87 [MBM05]. 88 [Jef03, KM05].

90 [QZ05a]. 91 [SR05]. 92 [GG06a, HNPE07].

Aalen [MS07]. Aberration [But03a, But03b, But05, FM00, Tan01, JAW09]. above [JL04]. Absolute [CP05, CYZZ08, LL05, LL08a, PY03, TCGW07, WLT07]. Accelerated [BRT01, PW03, Str05, TC06b, THW05, Zho05, JLWY03, JS09]. Accounting [CYCJ07, Wan03]. Accounts [PGP01]. accumulating [BS09]. accuracy [JP09a]. Accurate [DM02a, FR08, Ste06]. across [AF06]. Adaptation [LS04]. Adaptive [BB01, BCMR09, BKY06, BM05b, CS05, CB07b, FMT04, FW08, JT06, NK05, PSH06, TM03, WJT02, WSG09, ZL07, Guo09]. added [AR02]. added-variable [AR02]. Adding [MO97, MO05]. Additive [CRW02, JZ07, KL00, LTR^{+}08, LK07, MS02a, MS02b, MS07, SK05, YC04, CMMY09, SMLG09]. Additive-Index [LK07]. Additivity [CKS03]. Adequacy [FLF02]. Adjust [CK08]. Adjusted [AM06, ŞM05, WH03, JP09a, ŞM09]. Adjusting [JP09a, PK04b, UF07]. Adjustment [CDP08]. adjustments [CH09]. Affine [HR02]. after [HL02a, KR09]. against [CT05, GC02]. Age [KNN08a, KNN08b, WCMR07]. Age-at-Onset [WCMR07]. Age-Period-Cohort [KNN08a, KNN08b]. Aggregate [DHO02]. AIC [LWZ08, Yan05]. Akaike [VB05]. Algorithm [BB02, CD04, EL01, GY03, LZ08, Mar01, RS01, Sun02]. Algorithms [SS00]. Aliasing [HL02b]. Alignment [GM06, MSV07]. Allowing [FM06]. Alternating [Kuk04]. Alternative [CF08b, Dat05, Dav02, PL01, RS05]. Amendments [Ano04a, Ano06a, Ano07a, DRS12, GG06a, HL06a, HNPE07, HMXZ09b, KM05, Kos01a, LY09a, MBM05, MO05, Pal10, QZ05a, SR05, SL06a]. Among [FK07, BKW05, Wan04]. Analogue [Han08]. Analysing [BRS05, HWZ06, JS02, SK00]. Analysis [AV07, AAD00, CWW00, CC04, Car05, CW08, CL00, CJY02, CJ05, CC05b, CYZZ08, CSB03, CF01, CH00b, Cuz01, DG04, DP02, DKLW08, DAV03, DK02, Gab02, GWS07, GZH06, Joh01, KB01, LL06, LKLK06, LKA03, LJ03, MBH00, MN07, NSW02, Nie01, NG00, Oak01, PK04a, PK01, PH07, Pre07, Rei08, Riv01, SR01b, SR04, SA00, SW05, SL02, TLR03, WZ06, XZ07, YP05, Zho05, Zhu06, vLMZ01, CM04, FFM08, Hal00, HLL09, HWZ02, RDG09, SR05, WW09, WO04]. Analytic [OW04]. Analyzing [SCS07, SB07]. Applicability [HJPR02]. Application [BBM00, GF00b, KP04, LMKP07, ML02, MO97, MR01, ONL07, SCS07, TWG07, vdLA00, GDRG09, MO05]. Applications [AKR03, CIS06, FSW06, FH04, GM06, HY08, LL06, LOS07, MN07, MC00, QZ05b, Sie04, LM09, WCK09]. Applying [Fat06]. Approach [Bea08, BRS05, BPR02, CW08, CL00, CI05, DLS01, DKLW08, GZH06, HQ05, LL05, MKB03, MN07, OO07, SL02, SK00, TL04, HMXZ09b, HMXZ09a, KR09, LCM09, LT09]. Approaches [DL05, QS02]. Approximate [IZ00, Kim02, PLF05, BCMR09]. Approximating [KM09]. Approximation [BP07, FF07, LZ08, Sev00, DM09]. Approximations [DW02, DM01, KO00, KW05, YJS02, HWZ02]. ARCH [CP05, PY03]. arcsinh [JP09b]. Area [DRS05, DRS12, GMR08, LMKP07, YL08a, CT06, LR09, GM04]. Arising [CM08, DL01]. Arnason [KB03]. Arnason-Schwarz [KB03]. Arrays [But05, Tan06]. Ascertainment [Cla03, PGP01]. Assays [HN01]. Assessing [ALS03, BKW05, BSCC01, QS04, WD00b]. Assessment [CDI04, MBM00, PK04b, MBM05]. Assisted [BCO05]. Association [AST05, CF01, Cox03, FK07, FJ00, OW03, WD00b, Whi04]. Associations [BRL02, BRN08, MB05]. assumption [O’Q03]. Aster [GWS07]. Asymmetric [BJ02]. Asymptotic [CC07, CKO09, CHN03, HN01, Hel00, HN00, LL08b, YJS02, ZZ05, dRG01]. Asymptotically [GTW04]. Asymptotics [FM06, LR08, ZFH08, JS09]. Attributable [Ros01]. Augmentation [Cla03]. Augmented [LSW06]. Autocorrelated [LL04]. Autocorrelation [DDW00]. Automatic [RS07, Uek09]. Automodels [HY08]. Autoregressions [BM05a, RW02]. Autoregressive [HN00, IWLA03, LLW01, LL08a, SS03, WL01, WCK09]. Auxiliaries [SR01a]. Auxiliary [FSW06, JZ07, LT08, YL08a, ZR00]. Available [KO00]. Average [GL08, LL08a, Pre07, Rød06, SCW07, CHIM09, WJ00].

B [Jef03]. Back [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d]. Backfitting [HWC04]. Balanced [DT04a, RF03, TF04]. Bands [CHT06b, HN00, NP08]. Bartlett [CC06b]. Based [AST05, BL03, CDI04, CD04, CH00b, CHN03, Cuz01, Del06, FM05, HL02a, HCR06, IWLA03, JSH07, JHHB01, LMKP07, LB06, TLZW04, TCGW07, WR02, WZ06, WD08, Zha01, ZL05, ZZ04, Zou08, vLMZ01, Bra03, CTSW08, GM04, Ste06, JLWY03, JY03]. Bases [CI05]. Basic [DRS05, DRS12]. Basis [KAI04, HWZ02]. Bayes [And07, BLP00, BSM06, CSXB08, Dun09, GF00a, GMR08, GS04, SMLG09, Swe01, WW06]. Bayesian [SR05, And07, BE01, BCMR09, Bla03, BB03, CS09, CM08, CDI04, CC08, CS05, DP02, DFR06, DGK01, DP08b, FFM08, Gib00, GM06, Han09, Hof03, JAW09, Kim02, KB01, KB02, KAI04, Laz03, LLM04, LMP07, LMP01, LS00, MR01, NG00, NK05, OO02, Pad02, PB01, PLF05, PCK06, RDG09, Sch05a, SR01b, SR04, SDH04, Sev07, SIC03, SS03, Tar02, Ver00, Wal03, WLP05, WLS06, WW09, Wan09, WJT02]. Bayesian-inspired [JAW09]. Be [ES01, RF03, Yan05]. Benjamini [Wes08]. Benjamini-Hochberg [Wes08]. Bernoulli [CJ06]. Best [BM05b, DL05]. Beta [IZ00]. between [DK04, FM00, GC04, Yan05]. Bias [DT04b, DL01, KF09, KC06, PLF05, Swe01]. biased [LT09, Tsa09]. biased-sampling [Tsa09]. Biases [Ros06]. BIC [Yan05]. Bilateral [KM01, MBM00, KM05, MBM05]. Binary [CL06a, CHB06, CW02, Cox03, FK07, Kuk04, LSW06, OZ01, Qaq03, WL03, Jos04, KCD08]. Bingham [KW05]. Binomial [Llo00, SZG08, TS03, DW09, FJ09]. Bioequivalence [SKP00]. Bioinformatics [GM06]. Biometrika [HNPE07, Jef03, GG06a, KM05, Kos01a, MBM05, MO05, QZ05a, SR05, WO04, AB01, Cox01, Dav01, Hal01, Oak01, Smi01, TC01, Ton01]. Biplot [Gow04]. Biplots [Gab02]. Bivariate [BRN08, CF08a, DR04, HN06, JVL05, LPL08, Nie01, OR00, WD00b]. Block [PP01, RF03, WW06, CT09, Gua09, LL09]. Blocking [But06, MDC02]. Bone [SK05]. Bonferroni [Ano04a, Guo09, SC96]. Boolean [MC00]. Boosting [DT04b]. Bootstrap [AH00, BHY01, DY08, Hel00, Kim02, LY03, PP01, PLF05, Sam03, SKP00, LL09]. Both [HN06]. Boundary [ONL01]. Bounding [Gib00]. Bounds [CT01, MB05]. Box [CTW05]. Branching [BBO02]. Breakdown [HSW00, MU01]. Breaking [DP08a]. Breslow [Che09]. Breslow-type [Che09]. Bridge [WL03, HMXZ09b, HMXZ09a]. Brownian [BBO02]. Bubbles [TWG07]. Building [CL06a]. Bürmann [CKS03].

Calculations [Shi01]. Calibrated [HL05]. Calibration [CK08, GF00a, Hel00, Wu03, Pal09, Pal10]. Can [Yan05]. Canonical [Nie01]. Capture [BE01, GJA07, HH05, KB02, Tar02, WLS06]. Capture-Recapture [BE01, HH05, KB02, Tar02, WLS06]. Carcinogenicity [HN01]. Carlo [AKM05, HJPR02, IZ00, JSH07, JK07, LR01, LT05, LS00, MPRB06, PR08, ZN00]. Carry [KFD01]. Carry-Over [KFD01]. Carryover [BK06]. Case [Bon07, CC05a, Che01, Dav02, JV04, KCS04, KL00, LSW06, LT06, NYK06, PK04a, RSC02, SR01b, SR04, Sev00, SA00, XL03, Zha01, ZLZ01, ZLWZ01, HP09, KC09, SR05]. Case-Augmented [LSW06]. Case-Cohort [Che01, KCS04, KL00, LT06, NYK06, SA00, KC09]. Case-Control [Bon07, CC05a, JV04, PK04a, RSC02, SR01b, SR04, XL03, Zha01, HP09, SR05]. Case-Deletion [ZLWZ01]. Categorical [CDI04, SR01b]. Causal [HV08, Kur07, RSSW03, CSTT09]. Cause [CF08a, GT05, TDM02, VR08, JF09]. Cause-Specific [CF08a, JF09]. Caused [WC08]. Causes [CD04]. Censored [BT00, BRT01, CC04, CTW05, CJY02, CC05b, CC06c, DAV03, EMY03, JLY06, JT05, KCS04, NYK06, SB07, TC06b, WT06, WCMR07, Zho05, vdLA00, LT09, YSZ06]. Censoring [AM06, Bet00, DK03, HN01, LB06, Oak08, OGC07, SR02, VYZ01, ZLZ01, ZL05]. Censorship [CHT06b]. Centenary [Dav01, Hal01, Oak01, Smi01]. Central [CN06, YC03, Zen08]. Chain [DHO02, HJPR02, IZ00, JSH07, KNN08a, KNN08b, LR01, MPRB06, PR08]. Chain-Ladder [KNN08a, KNN08b]. Change [KFD01, MS07]. Change-Over [KFD01]. Change-Point [MS07]. Changepoint [BBM00, Gua04, WWM01]. Characteristic [CC02, HHF04, QZ03, JP09a, QZ05a]. Characteristics [Tan01]. Checking [LL05, Xia09]. Chi [Sch03]. Chi-Squared [Sch03]. choice [HR09]. Cholesky [DG04, Pou07, Rov00]. Choosing [CSB03]. Chordless [CW00b].

Circulant [KDA00]. Circular [DM02b, SHD02]. Claim [LTLE03]. Class [CJ01, EC02, GL08, vdLA00, AEB07, Bra03, Ste06]. Classification [HA03, He09, JP09a]. Classifiers [WSL08, CH09]. Clinical [BM05b, CSB03, CS05, CB07b, Cuz01, LS04, TL04, TM03, WT06, LSS09]. Clipped [WLT07]. Closed [KB03, MKHT09]. Closed-form [KB03]. Cluster [BRS05, BY08, FMT04, HSW01, LNO07]. Cluster-Correlated [BRS05, LNO07]. Clustered [CWW00, CB07a, CJ05, CHB06, Gli07, HWC04, Hug06, LL00, LC01, LWWC04, SL02, GK04, GC02, JS09, LY09a, LY09b]. Clustered/Longitudinal [LWWC04]. Clustering [KTV06, vLMZ01]. Cochran [Cox07]. Code [KO00, Xu05]. codes [CGOO09]. Coefficient [ŞM08, XZT04, HWZ02]. Coefficients [GT05, WW06]. Coherence [Che05]. Cohort [Che01, KCS04, KNN08a, KNN08b, KL00, LT06, NYK06, SA00, WCMR07, KC09]. Column [BW07]. Columns [CT01]. Combination [TF04]. Combinatorial [RF03]. Combined [HN01, SL02]. Combining [ES01, Qin00]. Comments [WO04]. Common [BPR02, PY08, Sch03, Zhu06, BRVW09]. Community [BBL04]. Comparing [CI05, FW08, Fin02, HZ00]. Comparison [DK04, DL05, JHHB01]. Comparisons [BDS06, Bra03]. Compatibility [Wan04]. Competing [BRL02, BRN08, BS08, CF08a, CD04, FJC01, GT05, JvdLH03, Lee06, SA00, PF07]. Complete [Sch05b, GC02]. Complex [BS08, BCO05, Cla03, KO00, Fat06]. Complexity [CW02]. Compliance [RR04]. Component [Boi03, CH00b, Fra03, JHS00, Rei07, Sch03]. Components [BW07, Bla03, BPR02, CHN03, Ger08, HN06, LL02, LMR01, MU01, ZHC08, Zhu06, Jef03, WD00a]. Componentwise [ST05]. Composite [VV05, GDRG09, MKHT09]. Computation [CIS06, BCMR09, LS00]. Computationally [BRT01, CHB06]. Computer [But01, KO00, OO02, BST09, CGOO09, Fat06]. computer-intensive [Fat06]. Computing [FR07, AKM05]. Concentration [DP04]. concerning [HE04]. Concordance [GH05]. Conditional [Cla03, Cox03, DY08, EL01, FY04, LLW01, LL05, LL08a, LWZ08, LOS07, Nie01, QZ05b, RW02, Sev00, VB05, WL03, YJS02, YC07]. Conditionally [WI08]. Conditioning [Han03, LT05]. Conditions [AMMC07, Kos99, Kos01b, RF03, VR08, Wal03, Kos01a]. Confidence [BBO02, Bot03, CM08, CQ00, CHT06b, DM02a, FM05, HHF04, HL05, HN00, LL06, NP08, Sha09, UTW05, AEB07, LSS09, LL09]. Confirmatory [PK01]. Conflict [Yan05]. Confounding [WC08]. Conjecture [NCGD03]. conjectured [Pal09, Pal10]. Connection [FM00]. Connections [QZ05b]. Consequences [DC08]. Conservative [Ano04a, SC96]. Considerations [LS04]. Consistency [Oak08, RSSW03, Sam03, SKP00, Wal03, WLP05]. Consistent [CNFC00]. Constant [OGC07]. Constant-Sum [OGC07]. Constants [KW05, MPRB06]. Constrained [MH07]. Constraints [AH00]. Constructed [CR04]. Constructing [But03b]. Construction [But03a, But05, LMT09, SL06a, SL06b, SLL09]. Context [BP07, DM01]. Contiguity [CGR04]. Continuous [CI05, Vid01, WI08, SZK09]. contrasts [Wes08]. contributions [BB03, Hal00]. Control [BKY06, Bon07, BDS06, CC05a, JV04, PK04a, RSC02, SR01b, SR04, XL03, Zha01, GRW06, HP09, SR05]. Conventional [BGD07]. Converge [SS00]. Convergence [Gib00, Sun02]. Convex [Mey03]. Coordination [TF04]. Copula [FJ00, OR00, OW03, PCK06]. Corners [HH07]. Correction [CC06b, DL01]. Corrections [Ano04a, Ano06a, Ano07a, DRS12, HL06a, HNPE07, Kos01a, MBM05, MO05, Pal10, QZ05a, SR05, SL06a, GG06a, HMXZ09b, KM05, LY09a]. Correlated [AKR03, BRS05, GWC04, HS03, HCR06, KKPS04, LNO07, Lin08, MP03, OZ01, PDT05, Qaq03, QLL08, Rao04, Vic00, YL08b]. Correlation [Boi03, CJ06, CH00b, LLM04, Nie01, Pou07, QLL08, Wan03, WC03, GG06b]. Correlations [Mar00, Qaq03]. Correspondence [Gab02]. Costs [BT00]. Count [CP02a, HS03, HWZ06, JS02, LZH07, SF03, Zha02, Zha06]. Counterexample [LTLE03]. Counterfactual [VR08]. Counting [HS01, ZL06]. Counts [FSW06, CL09, DDS03, DW09, HLL09]. Covariables [WR02]. Covariance [AAD00, Boi02, BW05, CDR07, CF08b, CNFC00, CH00b, DP02, DAV03, HLPL06, JW05, KC06, LGS08, ND02, PM03, Pou00, Pou07, WCK03, WP03, YP06]. Covariances [CN06]. Covariate [LL00, PH07, PK04b, RSC02, ŞM05, ŞM09, Waa08, WC03, JP09a]. Covariate-Adjusted [ŞM05, ŞM09, JP09a]. Covariates [CIS06, GS01, JZ07, LWC07, LT08, PGP01, RSG07, SR01b, TD01, LTX09]. Coverage [BBL04, ML02, MC00, Swe01]. Cox [CTW05, CIS06, Fin02, GS01, Lin00, LTX09, MKB03, SW03, SIC03, WNZZ09, ZL07]. Credible [FM06]. Criteria [CC08, KAI04, Kur07]. Criterion [And07, CDI04, Fat06, Wan09]. Cross [CF08a, FK07, MF09]. Cross-Product [FK07]. cross-sectional [MF09]. Crossover [BK06]. Crossvalidation [FY04, ORSV01, HR09]. Cube [DT04a]. Cumulative [SZG08]. Cure [LX05, LY04]. Current [Bet00, JvdLH03, JV04, JVL05, LX05, MS02a, TC06b, WD00b, vdLA00]. Curve [CC02, DGK01, TM08, JP09a]. Curve-Fitting [DGK01]. Curved [RS05]. Curves [Ano06a, GG05, HHF04, QZ03, VYZ01, GG06a, QZ05a]. Cutting [HH07]. Cycle [CW00b]. Cyclical [HW00].

D [Jef03]. Dabrowska [OW03]. Dantzig [JR09]. Data [BT00, BRS05, Bet00, BRT01, CWW00, CC04, CFLZ05, CP02a, CW08, CT04, CB07a, Che00, Che01, CJY02, CDI04, CJ05, CC05b, CL06a, CC06c, CW00a, CF08a, Cla03, CHB06, CL08, DP02, DHO02, DAV03, EMY03, ES01, FJC01, FLF02, Gli07, HN06, HZF02, HS03, HWC04, HWZ06, Hug06, HH05, ICL01, JHS00, JvdLH03, JV04, JVL05, JLY06, JS02, KB02, KCS04, KP04, Kuk04, KDL07, LX05, LNO07, LL00, LPL08, Lin00, LC01, LWWC04, LH06, Lin08, LZH07, MS02a, ONL07, Pad02, PK04a, PK01, PDT05, PF06, QS02, QLL08, RSC02, Rei07, SC04, SB07, ŞM08, SK05, SL02, SF03, SK00, TLR03, TM08, TC06b, THW05, Waa08, WT06, WLP05, WD00b, WR02, WCMR07, Whi04, WP03, XL03, XZ07, YY01, YP05, YP06, YC04, YL08b, YLK07, Zha01, Zha02, ZHC08, ZLWZ01, ZFH08, vdLA00, CTD09, CMMY09]. data [CPQ09, GK04, GDRG09, HS09, JS09, Jos04, JY03, KCD08, KR09, LL09, LY09a, LY09b, LT09, Nor09, PF07, RDG09, Tsa09, WCK09, YSZ06, Zha06, CM04]. Data-informed [CM04]. Database [CC06a]. Dealing [CHIM09]. death [MF09]. Decision [CSB03]. Deck [KF04]. Decomposability [FD03]. Decomposition [DG04, JW05, LKLK06, Rov00]. Decompositions [Pou07]. Deconvolution [HQ05]. Deficiency [CM01]. Degrees [HS01, Rei07, CH02a]. Degrees-of-freedom [CH02a]. Delayed [GM01]. Deletion [PQP08, ZLWZ01]. Densities [BD06, CR04, FY04, Wan04]. Density [Del06, DT04b, FZ04, HN06, JHHB01, JH07, NP08, Ren02, RDG09]. Density-Based [JHHB01]. Dependence [BE01, CP02b, Gli07, HZF02, LMP01, MB05, ST03, Guo09, KMT03, NGCD09]. Dependent [CJ06, CRA06, CT04, SHD02, VYZ01, LL09, SZK09]. Depicting [HH02]. Derivation [Dat05]. Deriving [Cox08, FR08]. Design [AB01, BC02, Che01, KCS04, LS04, Ros04, SA00, TM03, Ver00, WC03, Ful09, JG09]. Designs [BK06, BW07, BB01, BDS06, But01, But03a, But03b, But06, CT01, CS05, CB07b, DK04, Edm02, HL02b, KFD01, MP03, RF03, SL06a, SL06b, Tan01, TGM00, WT06, Xu05, BST09, CT09, DT09, Fat06, JAW09, LM09, PWB08, Qia09, QW09, SLL09, WJ00]. detectability [FJ09]. Detection [BE01, BPR02, KKPS04]. Determining [Zen08]. Deviation [CYZZ08, LL05, LL08a, WLT07]. Deviations [CP05, FH04, PY03]. Diagnosis [SXZ08]. Diagnostic [BL03, HCR06, LL05, QLL08, SW03, TC06a, ZZ04, ZITZ08]. Diagnostics [Car05, Gli07, PQP08]. Diagonal [JL04]. Dichotomous [RR04]. Difference [MBH00, SW03, GC04]. Differences [TDM02]. Different [Ren02]. Differential [Ros06]. Diffusion [Bla03, DFR06, LR01, RS01, Sho02]. Dimension [AMMC07, CLC07, Li07, Peñ09, YC07, Zen08, CPQ09, Xia09]. Dimensional [DKLW08, GTW04, HT02, KP04, Lin08, SXZ08, CH09]. Dimensions [Riv01, Sch05b]. Direct [BK06, KFD01, SZG08]. Direction [YC05, BRVW09]. Directional [Riv01]. Directions [Lue04]. Dirichlet [DGG07, IZ00, KTV06, PR08]. Discovering [LMP07]. Discovery [BKY06, LZ08, GRW06, KL09]. Discrepancy [FM05]. Discrete [Bla03, CL06b, HQ05, Ros01, Sev00, Vid01, KL09]. Discrete-Transform [HQ05]. Discretely [TWG07]. Discriminant [EC02, LJ03, Zhu06]. Discrimination [AS06, BC02, DK04, KB02, Ren02]. Discriminatory [GH05]. Disease [AST05, KC09]. Dispersions [CHN03, LN01]. Disseminated [Rei08]. Distance [Bon05]. Distortion [WC08]. distributed [Hal00]. Distribution [AV07, AM06, BL03, BP07, CW02, Dat05, HH02, HN01, HCR06, HB06, Kom01, LBS02, OO02, PK04b, SR02, Sev00, TC06a, TCGW07, WL03, WW06, GC04]. Distributions [CMW07, CI05, CH02b, CHN03, Dat05, HT02, JL04, LF05, LB06, MH07, MO97, NCGD03, Qaq03, TDM02, WT06, WCMR07, WI08, YJS02, JP09b, KL09, MO05, MPRB06, PB02, ZN00]. Divergence [JHHB01]. DNA [BBM00, CC06a]. domains [Nor09]. Dose [Che05, Imb00]. Dose-Finding [Che05]. Dose-Response [Imb00]. Double [FMT04, LL09, LL08b]. Doubly [CC04, vdLA00, CTD09, Moo09]. doubly-robust [Moo09]. Down [HH07]. Driven [FSW06, GS07, DDS03]. Duration [AV07, CYCJ07, JT05]. Duration-Response [JT05]. Dynamic [DP02, FLF02, CGOO09, Peñ09].

ECM [SS00]. Edge [SSM05]. Editorial [Tit01]. Edwards [Sun02]. Effect [CYCJ07, CDP08, HH05, LN01, MDC02, ZL05, AR02, BRVW09]. Effectiveness [CRW02, CYCJ07]. Effects [BK06, CPQ09, CHB06, CH02b, KFD01, Kur07, LS03a, LS03b, LWZ08, MKB03, PH07, PGP01, PB06, RSG07, RVM08, RR04, Ros01, Ros06, TGM00, VB05, WC08, CSTT09, CHIM09, Dun09, JP09a]. Efficacy [BB04]. Efficiencies [CH00b]. Efficiency [LT08, MU01, CTD09, WJ00]. Efficient [AST05, BSM06, BC02, BDS06, CSTT09, DT04a, DK02, FH04, HOR00, Jos04, LH06, MCW06, MS02a, PCK06, RP04, TM04, WT06, WCK03, XZT04, ZL06, MPRB06, SZK09]. Elicitations [OO07]. Empirical [And07, CBD02, CM08, CQ00, CSW02, CC06b, DM01, FM05, FM06, GF00a, GMR08, LMKP07, Laz03, Qin00, QZ05b, Sch05a, SMLG09, VR08, WR02, WW06, XZ07, ZR00, Zho05, ZL05, ZITZ08, ZFY02, ZF02, CPQ09, Nor09]. Empirical-Likelihood-Based [ZL05]. Empirical-Type [CM08, FM06]. Empirically [CHP00]. Empty [BBO02]. emulation [CGOO09]. Endpoints [TL04, LSS09]. entire [WSG09]. Environment [CC05a]. Epidemics [BBO02, BBL04]. Equal [HN01]. Equality [HT02]. Equation [BRS05, QS02]. Equations [KCS04, Lin08, Pan01, PQP08, QLL00, QS04, SL02, SK00, WC03, YP06, YJS02, YL08b, ZITZ08, Uek09]. Equivalence [SR04, SR05, TL04]. Equivalent [LWWC04]. Equivariant [HR02]. Ernest [BB03]. errata [Pal10]. Error [BE01, CW08, HSD06, LWC07, Lue04, Mar03, TCGW07, TD01, TM04, Ver00, YL08a, KL09, LR09]. Error-Prone [LWC07]. Errors [BW05, CP05, DM00, Hel00, KKPS04, LL04, LL00, LTR^{+}08, MP03, PK04b, Rao04, WR02]. Errors-in-Covariables [WR02]. Estimands [DM02a]. Estimate [Dem00, FM01, HL06a, HL06b]. Estimated [HYE07, LL05, TCGW07]. Estimates [DT04b, HK01, XL03]. Estimating [AM06, BT00, BB04, BRS05, CRA06, CYCJ07, DHO02, FY04, GTW04, HH02, HN06, Han03, HE04, Imb00, KCS04, LZ08, Lin08, Pan01, PW03, PQP08, QZ03, QLL00, QS02, QS04, Sev02, SND05, SL02, SK00, TS03, TLZW04, TW05, Waa08, WH03, WC03, WO04, Whi04, YP06, YC03, YL08b, ZITZ08, Fat06, HLL09, QZ05a, Uek09]. Estimation [AST05, Ano06a, BL03, BBL04, BRN08, BSM06, BE01, BC02, Bon05, BCO05, CW08, CP05, CC05a, CDR07, CF08a, CH00a, CN06, CH02b, CDP08, Del06, DM01, DCNZ05, GG05, Ger08, GMR08, GF00b, GJA07, HRR00, HZF02, HLPL06, HB06, Hug06, HOR00, HH05, IWLA03, JvdLH03, JZ01, JH07, JK07, KN05, LMKP07, LS03b, LLW01, LL08a, LPL08, Lie01, LLM04, LMP07, LH06, LH01, Llo00, LZH07, MBH00, MS02a, MC00, MU01, NT08, NP08, ND02, PDT05, PY03, PF06, Pou00, Pou07, Pre07, RR04, RS07, SR02, SR01a, SDH04, SCW07, Sho02, VRR07, WT06, Wal00, WC03, WCMR07, WSL08, WCK03, WP03, XZT04, XL03, Yan05, YL08a, YC05, YL07, ZL06, Zha02, CT06, Che09, CSTT09, CHIM09, FSR06, GG06a, GM04, GDRG09, Gua09, KR09, LY09a, LY09b]. estimation [LR09, LT09, MF09, MKHT09, RDG09, SMLG09]. Estimative [UF07]. Estimator [Ber07, BD06, LS03a, LL02, MCW06, OMS04, Pan01, PLF05, Sev00, TH05, TD01, Zho05, vdLA00, CTD09]. Estimators [BW05, CSW02, CNFC00, CH00b, Dat05, DRS05, DRS12, GT05, HL02a, JHHB01, MH07, Ren02, RW00, TM04, Wu03, CFVG09, CKO09, HSW00, LTX09, Moo09, Ste06]. evaluating [HP09]. Evaluation [And07, TCGW07, O’Q03]. Event [Cuz01, YY01]. Event-Based [Cuz01]. Event-Time [YY01]. Events [AM06, Ebr06, FH04]. Exact [CRCW05, Mar00, SMG02]. Examination [HH05]. Excess [FM01]. Exchangeability [Kel04]. Exclusion [SSM05]. Expansion [CKS03]. Expansions [Hel00, dRG01]. Expectancy [CC06c]. Expected [FM05, PB02]. Expected-posterior [PB02]. Experiments [AB01, But01, CL00, ES01, MP03, NG00, Ros01, BST09, JG09]. Exploiting [CC05a, ST05]. Exponential [AS01, BP07, DY08, ELP03, MO97, RS05, GM04, KF09, MKHT09, MO05, Tan09]. Exponentially [Sch05a]. Extended [CC08, KNN08a, KNN08b]. Extending [BGD07]. Extension [TF04]. extensions [JS09]. Extra [HN06]. Extremal [CP02b]. Extreme [ST03].

Factor [BSM06, GWC04, Kel04, LKA03, Tan06, Vic00, Peñ09, WJ00]. Factorial [But03a, But03b, But06, NG00, Tan01, CT09]. Factorials [FM00]. Factorisable [RP04]. Factorizable [FR07]. Factors [BB01, BLP00, PY08]. Failure [BL03, BRL02, BRN08, BRT01, CWW00, CFLZ05, CD04, GT05, Gli07, Lee06, PW03, SR02, SC04, TC06b, THW05, TDM02, YC04, YL08b, ZLZ01, Zho05, JLWY03, JS09, LY09a, LY09b]. Failure-Time [Gli07]. False [BKY06, GRW06, LZ08, MB05, KL09]. Families [AS01, DY08, MO97, RS05, MKHT09, MO05]. Family [AST05, CSXB08, MO97, NSW02, PGP01, Qaq03, Whi04, YJ00, GM04, GG06b, KF09, MO05]. Family-Based [AST05]. familywise [KL09]. Fast [Gua09, KO00]. Featureless [KDA00]. fiducial [HL09]. Fields [BB02, LGS08, TWG07]. Fieller [BLP00]. filling [QW09]. Filter [Cho02]. financial [WCK09]. Find [BM05b]. Finding [Che05]. Finite [PL01, LCM09]. First [BM05a, Cox01]. First-Order [BM05a]. Fisher [Bot03, KW05, LBS02, LCM09]. Fisher-Bingham [KW05]. Fisz [FSR06]. Fit [ACH00, Bon07, BSCC01, Gab02, Gua08, GC02]. Fitting [BBM00, Che01, DGK01, LSW06, Lin00]. Flexible [KCD08, MS02b, ZHC09]. Following [LH01, HY03]. Forecasting [KNN08a, Pal09, Pal10]. Forensic [CC06a]. form [KB03]. Forming [WLS06]. Formula [Joh01, GK04]. Forward [AR02]. Four [CW00b, JAW09]. Four-Cycle [CW00b]. four-level [JAW09]. Fractal [FF07]. Fractional [But03a, But03b, KF04, Lie01, RF03, Tan01, CT09]. Fractionally [LL08a]. Fractions [FM00]. Fragments [MSV07]. Frailty [DLS01, DL05, GZH06, HS03, LL00, ZCI09, ZHC09]. Frameworks [ZZ05].

Free [DGK01, FR07]. Free-Knot [DGK01]. Freedom [HS01, Rei07, CH02a]. frequency [Nor09]. Frequentist [CM08, FM06, LF05, LC03, PB06]. Front [Ano00e, Ano00f, Ano00g, Ano00h, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano03h, Ano04f, Ano04g, Ano04h, Ano04i, Ano05e, Ano05f, Ano05g, Ano05h, Ano06f, Ano06g, Ano06h, Ano06i, Ano07f, Ano07g, Ano07h, Ano07i, Ano08e, Ano08f, Ano08g, Ano08h]. Full [GZH06]. Fully [SS03]. Function [DM01, EL01, HRR00, Kag01, KN05, KAI04, LZH07, Mey03, MR01, Pre07, SND05, WH03, WL03, ZLZ01, GM04, HWZ02]. Functional [CL08, GS01, Ger08, HCR06, JHS00, TM08, TM04, ZHC08, RDG09]. Functional-Based [HCR06]. Functionals [ELP03, GDRG09]. Functions [AS01, BKW05, BPR02, CH00a, CH00b, DM02a, EC02, GMR08, GTW04, HH02, Han03, HZ00, HE04, Imb00, PW03, QLL00, QS04, SC04, Sev02, Sev07, TLZW04, Waa08, CFVG09, HLL09]. future [CTSW08]. Fuzzy [TG07, KL09].

G [Cox07]. Gamma [HS03, ZCI09]. Gap [AM06, SC04, Str05]. GARCH [PY03]. Gaussian [AKM05, CS09, CGR04, CGOO09, CW02, Dat05, DP04, JW05, JK07, Lie01, Mar00, NCGD03, PCK06, SSM05, TH05, WD00a, WO04, YL07]. Gene [CC05a]. Gene-Environment [CC05a]. General [BGD07, CJ01, CC07, Dav01, FR07, GZH06, Han03, Kos99, Kos01b, MN07, MB05, RP04, ZITZ08, Kos01a]. Generalised [AKR03, But05, Edm02, FZ04, HP01, HK01, ICL01, JZ01, KN05, LS03a, LN01, LL02, LS00, MN03, NCGD03, ORSV01, Pan01, Pou00, QLL00, QS04, SK00, TC06a, Vid01, WC03, YP06]. generalization [Cox07]. Generalized [DGG07, HL09, LY09a, LY09b, SCS07, ŞM08, JR09, KCD08, ŞM09, Wan09]. Generated [OW03]. Generating [OZ01]. Generation [CW00b]. Generic [Ros06]. Genetic [PGP01, Whi04]. Genomic [ML02]. Geometric [AMMC07]. Geometry [Gow04, Mar02, Mar03]. Gibbs [Gib00, LS00, LOS07]. Goodness [Bon07, BSCC01, Gab02, Gua08, GC02]. Goodness-of-Fit [Bon07, Gua08, GC02]. Graphical [CMW07, CL06b, FD03, GS04, JW05, Kur07, Mat06, SSM05, YL07, AKM05, CS09, O’Q03, WW09]. Graphs [DP04, RW02]. Group [BJ02, JT06, LS04, LL06, TS03, DJ00, HMXZ09b, HMXZ09a, PWB08]. grouped [WNZZ09]. Grouping [LH06]. Growth [VYZ01]. Guaranteed [GHJK00].

Haar [FSR06]. Haar-Fisz [FSR06]. Haplotype [AST05]. Haplotype-Disease [AST05]. Hastings [GM01, RS01]. Hazard [CHT06b, CF08a, JZ07, MS02b, PK01, SC04, YP05]. Hazards [CJ01, CRW02, DL01, Ebr06, GH05, KL00, Lin00, MS02a, MS07, SK05, TD01, WS01, YY01, YC04, Zou08, KC09, O’Q03, SMLG09, Tsa09, ZL07]. heavy [WCK09]. heavy-tailed [WCK09]. Hessian [Lue04]. Heterogeneity [AV07, BE01, HH05, MBH00, Tar02]. Heteroscedastic [IWLA03, LKA03, MCW06]. Heteroscedasticity [CNFC00, LLW01, LL05, LL08a]. Heteroscedasticity-Consistent [CNFC00]. Hidden [SW05]. Hierarchical [And07, BKW05, DLS01, DL05, GM06, HS01, IZ00, LN01, Mei08, YLK07, PR08]. Hierarchically [WNZZ09]. High [AMMC07, HT02, KP04, MU01, Sch05b, SXZ08, dRG01, CH09]. High-Dimension [AMMC07]. High-Dimensional [HT02, KP04, SXZ08, CH09]. High-Order [dRG01]. Higher [FM06]. Higher-Order [FM06]. Hilferty [Ter03]. History [GWS07, Bin00, BB03, GG06b, Hal00]. Hochberg [HH07, TL08, Wes08]. Holds [AMMC07]. Holm [Guo09, HH07]. Homogeneity [HP01, LCM09]. Homogeneous [GJA07]. Horvitz [Fat06]. Horvitz-Thompson [Fat06]. Hot [KF04]. Households [BBL04]. Hundred [AB01, TC01]. Hybrid [Kuk07, LS03a]. Hyper [CMW07, Rov00]. Hyper-Inverse [CMW07]. Hypercube [But01, SL06a, SL06b, Qia09, SLL09]. hypercubes [LMT09]. Hypotheses [BGD07, KP04, MB05, Ros08, BRVW09]. Hypothesis [Dav02].

Identifiability [EMY03, Kur07, LK07, OGC07]. Identification [CC06a, GWC04, KNN08b, Lee06, SW05, Vic00]. ignorability [HS09]. Ignorable [Pad02, QS02]. illness [MF09]. illness-death [MF09]. Illustrated [SK05]. Image [Gib00]. Imperfect [BL03]. Implementation [Swe05]. Implementing [LC03]. Implications [Pre07, WC03]. Importance [FH04, HYE07, Mei08]. Improve [YJ00]. Improved [Ano04a, CN06, CNFC00, HW00, LBS02, SC96, Vid01]. Improving [CTD09, CFVG09, LT08, QLL00]. Imputation [ACHM02, Che00, GY03, Kim02, KF04, Pad02, PLF05, Rei07, Rei08, RW00, TDM02, KR09]. Incidence [SZG08]. Included [CC06a]. inclusion [Fat06]. Incomplete [Edm02, RF03, Waa08, ZLWZ01, CTD09]. Indentification [Yan05]. Independence [CB07a, CC05a, DM00, RW02, Sch05b]. Index [KX07, LK07, NT01, YC05]. Indirect [WC08]. Individual [SKP00]. Individually [XL03]. Individuals [CC06a]. Induced [JS09]. Inefficiency [TM03]. Inference [Ano07a, Bea08, BS08, Bla03, CB07a, CC07, Cla03, CP02b, CD04, DY08, DP08b, FF07, FJ09, FR07, GS07, GTW04, HY03, HNPE05, Hel00, HA03, HV08, JT05, KP04, LB06, LC03, LMP01, OO02, PB01, PGP01, PCK06, QLL00, QS04, RSC02, RS01, RW00, RSSW03, Ros01, ST03, Sev07, ST05, TLZW04, VV05, WR02, WD08, ZR00, ZL05, HNPE07, HL09, JLWY03, LL08b, PF07, SZK09, Ste06]. Inferences [CW00a, PB06]. Infinite [GTW04]. Infinite-Dimensional [GTW04]. infinitely [Wes08]. Influence [ALS03, BPR02, CH00b, GMR08, Pre07, ZZ04, CM04]. Information [And07, Ano00i, Ano01i, Ano02i, Ano03i, Ano04j, Ano04k, Ano05i, Ano05j, Ano06j, Ano06k, Ano07j, Ano07k, Ano08i, Ano08j, Ano08k, Bot03, CC08, KAI04, VB05, YL08a, Zha01, ZR00, LCM09, Wan09]. Informative [HWZ06, PDD06, QLL08, SR02]. informed [CM04]. Inhomogeneous [BSCC01, Gua08, Waa08, Gua09]. inspired [JAW09].

Integrated [LL08a, Sev07]. Intensity [CH00a, RSG07]. intensive [Fat06]. Interactions [Tan06, VR08, DT09]. Intercept [WL03]. Intercorrelation [FLF02]. Interest [Swe05]. Internal [Cox08]. Interpolated [HL05]. Interpolating [DHO02]. Interval [BRT01, JH07, LB06, OGC07, SB07, TC06b, ZLZ01, CFVG09, YSZ06]. Intervals [CM08, CQ00, FM05, HHF04, HL05, LL06, LF05, SDH04, AEB07, LSS09, LL09, Sha09]. Intervention [HV08]. intractable [MPRB06]. Intraslice [CN06]. Intrinsic [BM05a, SS03]. Invariance [Kuk04]. Invariant [SMG02]. Inverse [CMW07, Dat05, LCT07, NCGD03, NCT05, Pre07, Rov00, YC03]. Inversion [CLC07]. Irregular [Sie04]. Isotonic [AS06, WWM01]. Issue [MN07]. item [KR09]. Iteration [LY03].

Jackknife [Ber07, LR09]. James [TH05]. Joint [THW05, ZHC08]. Journey [Ton01]. Jump [BSM06, GM01, JSH07]. Justification [SIC03, Tan01].

Kazakhstan [SCS07]. Kendall [Oak08]. Kernel [DT04b, DP08a, HWC04, JH07, LWY06, Wan03, YL08b]. Kernel-Type [JH07]. Kernels [LWWC04]. Khmaladze [O’Q03]. Khmaladze-type [O’Q03]. Kin [WCMR07]. Kin-Cohort [WCMR07]. Kinematics [GS04]. Knot [DGK01]. Kolmogorov [Bin00].

Lack [ACH00]. Ladder [KNN08a, KNN08b]. Lag [CRW02]. Landmark [ALS03, KDL07]. Large [CC08, OMS04, WSL08, WP03]. Large-Margin [WSL08]. Large-Sample [OMS04]. lasso [Han09, ZL07]. Latent [BP07, CRW02, HSD06, Lee06, HS09, TG07]. Latent-Model [HSD06]. Latin [SL06a, But01, LMT09, Qia09, SL06b, SLL09]. Lauritzen [Sun02]. Layout [BDS06]. Least [CP05, CYZZ08, Dem00, JLY06, LL05, LL08a, PY03, TW05]. Least-Squares [JLY06]. Lebesgue [Bin00]. Left [PF06]. Left-Truncated [PF06]. length [LT09]. length-biased [LT09]. Lengths [FM05]. Level [But03a, But06, DRS05, DRS12, FM00, GHJK00, PDD06, TGM00, CT09, JAW09]. Lévy [GS07, WD00a]. Lhoste [BB03]. Life [AM06, CC05b, CC06c, GWS07, JF09]. Lifetime [LB06]. Likelihood [Ano06a, CT05, CC05a, CQ00, CSW02, CC06b, Cla03, CRCW05, Dat05, Del06, DM01, DLS01, DL05, DR04, EL01, FZ04, Fra03, GG05, GZH06, HL02a, HK01, HA03, HLPL06, Kuk07, LBS02, Laz03, LPL08, LL02, Lie01, Llo00, MH07, Mar00, Pou00, QZ05b, RS05, Sch05a, Sev00, Sev04, Sev07, SIC03, ST05, Swe01, VV05, WR02, XZ07, YLK07, ZR00, Zho05, ZL05, ZITZ08, ZFY02, ZF02, vdLA00, AKM05, Bra03, CPQ09, Che09, GG06a, LT09, LTX09, MKHT09, Nor09, Tan09, Tsa09, WO04]. Likelihood-Based [WR02, Bra03]. Likelihoods [CM08, FM06, Qin00, Sar03, KB03, MKHT09]. Limit [HL02b]. limited [CHIM09]. Limits [DM02a, MR01, UF07, Vid01]. Linear [AKR03, BGD07, BKY06, BC02, CHT06a, CQ00, CC06c, Dav02, DL05, GT05, GS04, HP01, HK01, HCR06, HWC04, ICL01, JZ01, KN05, KP04, LS03a, LN01, LL04, LWC07, LTR^{+}08, LWZ08, LL02, MCW06, Mey03, NYK06, NT08, Pou00, TC06a, Ver00, Vid01, WR02, LY09a, LY09b, ŞM09, Wes08]. linearization [GDRG09, KR09]. Link [Kag01, KCD08]. Local [ACHM02, CQ00, CJ05, Del06, MH07, Mar02, Swe05, YL08b, ZZ04, Dun09]. Locally [AST05, TM04, FSR06, Ter03]. Locally-Efficient [AST05]. Loci [Sie04]. Log [FW08, LT08, AEB07, GK04]. Log-Rank [FW08, LT08, AEB07, GK04]. Logistic [AS06, Bon05, Bon07, CC02, EC02, Kag01, Kuk04, QI06, QZ03, Shi01, WL01, Zha01, QZ05a]. Logistic-Type [EC02]. Loglikelihood [CB07a]. Long [LMP01, TH05, YP05]. Long-Memory [TH05]. Long-Range [LMP01]. Long-Term [YP05]. Longitudinal [DP02, HZF02, HS03, HWC04, JS02, LWWC04, PM03, PK04a, QS02, ŞM08, SK00, THW05, TD01, WP03, XZ07, YP06, ZFH08]. Longitudinal/Clustered [HWC04]. Look [WWM01, CL09]. Low [AMMC07, CH09]. Low-Sample-Size [AMMC07]. Lower [MB05].

Machines [LKLK06]. Main [MDC02, TGM00]. many [Wes08]. Mapping [Sie04]. Margin [WSL08]. Marginal [CF01, Cox03, CR04, Ebr06, HLL09, HB06, KC09, Qaq03, QZ05b, RJT00, SCW07, WL03, Wan03, Wan04, YY01, ZFH08, AKM05]. Mark [CL00, HB06]. Mark-Recapture [CL00]. markers [HP09]. Markov [BB02, BS09, Bla03, DHO02, HJPR02, IZ00, JSH07, LR01, MPRB06, PR08, SK05]. Marrow [SK05]. Martingale [SW03]. Masked [CD04, AR02]. Mass [FM01]. Matched [RSC02, XL03]. Matching [GL08, LC03, MR01, SMG02, Swe05, WL03]. Matérn [GG06b]. Matrices [Boi02, Boi03, BW05, CJ06, CF08b, DP02, WP03]. Matrix [CDR07, CLC07, CNFC00, CH00b, HLPL06, KC06, Pou00, Pou07, Rov00, Zha01, WW09]. Matter [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h]. Maxima [ST05, TWG07, NGCD09]. Maximising [EL01]. Maximum [Ano06a, CC05a, Dat05, GG05, HL02a, HK01, LPL08, LL02, Lie01, Llo00, MKHT09, Mar00, Pou00, Sev00, vdLA00, Che09, GG06a]. Maximum-Likelihood [HL02a]. May [ES01]. Mean [Ano06a, BRS05, CC05b, GG05, GHJK00, LZH07, PM03, RR04, SR01a, SND05, VRR07, YC07, Zen08, CTD09, GG06a, LR09, Sha09]. Mean-Covariance [PM03]. Means [ELP03, PDT05, Qaq03]. Measure [CGR04, Cox08, OW03, ST03, Bin00]. Measured [TD01, YL08a, CMMY09].

Measurement [CW08, GF00b, HSD06, Kel04, LL00, LTR^{+}08, Lue04, Mar03, TM04, DT09, WO04]. Measurements [CYCJ07, HWZ02, JY03]. Measures [MP03, WZ06, ZLWZ01, ZITZ08, vLMZ01]. Measuring [DRS05, DRS12]. Mechanism [ICL01]. Median [BHY01, Ger08, HR02]. Medical [BT00]. Memory [TH05]. Mendell [Jef03]. Meta [MBH00]. Meta-Analysis [MBH00]. Method [Che00, Cho02, Cox08, CL08, DT04a, FY04, HH07, JYW01, Kuk07, MO97, ORSV01, SB07, SL06a, SL06b, Tar02, WLT07, Zha02, AKM05, LY09a, LY09b, MO05, MPRB06, Wes08]. Methodology [Dav01]. Methods [AAD00, AH00, BE01, CSW02, Fue02, Hof03, LR01, MC00, PL01, Sam03, SC04, TDM02, vLMZ01, AEB07, HP09, PR08]. Metropolis [GY03, GM01, RS01]. Metropolis-Hastings [GM01, RS01]. Microarray [CW08]. Mild [AMMC07]. Minimal [RF03]. Minimand [JYW01]. Minimum [Bon05, But03a, But03b, But05, JHHB01, Tan01, JAW09]. mis [CTSW08]. mis-specified [CTSW08]. Misclassification [Llo00]. Mises [LBS02]. Missing [CIS06, GT05, ICL01, LWC07, RSC02, Rei07, SR01a, TLR03, TDM02, WD08, YLK07, HS09, LTX09]. missing-data [HS09]. Missingness [QS02]. Misspecification [GS01, RVM08, WC03]. Misspecified [DL01, HK01, PK01]. Mixed [HP01, HK01, ICL01, JZ01, LS03a, LS03b, LWZ08, LL02, SCS07, TC06a, VB05, LR09]. Mixed-Effects [LWZ08, VB05]. Mixing [TC06a]. Mixture [CL06a, FSW06, KTV06, LMR01, Mar02, NG00, PL01, WL01, WJT02, ZFY02, DM09, Jef03, Tan09, WCK09, KMT03]. Mixtures [Ano07a, HNPE05, RJT00, ZHC09, GC02, HNPE07]. Mode [FM01]. Model [BC02, BPR02, Bon05, BCO05, CK08, CDI04, CIS06, CC08, CW00a, CHP00, CC07, CHB06, CH02b, CD04, DRS05, DRS12, DK03, Dav02, DDW00, DFR06, DR04, DP04, FLF02, GT05, GS01, GZH06, Gua04, HZF02, HS03, HCR06, HSD06, Joh01, KDA00, KB02, KX07, KNN08a, KNN08b, LX05, Lee06, LJ03, ML02, MS02b, MS07, NT01, NT08, OR00, OW03, OO02, PW03, QLL08, Riv01, SCS07, SW03, Sie04, SHD02, Str05, SXZ08, TC06a, TC06b, TCGW07, TGM00, TD01, VV05, Ver00, Vic00, WD08, WS01, WL01, Xia09, Yan05, YC04, YL07, Zho05, ZFY02, Zou08, AR02, CS09, DW09, DT09, FFM08, JLWY03, JS09, KC09, LL08b, LTX09, MF09, Peñ09, PB02, WCK09, ZL07]. Model-Assisted [BCO05]. Model-Based [WD08]. Modelling [AF06, BRL02, CF08b, FSW06, MKB03, Mat06, NGCD09, NG00, OZ01, PY08, PDD06, RSG07, THW05, YP06, ZHC08, PM03, ZHC09]. Models [AAD00, AKR03, And07, BGD07, BKW05, BS08, BC02, Boi02, Boi03, CTW05, Car05, CMW07, CM01, Che01, CJ01, CRW02, CJY02, Cho02, CP02b, CF01, CL06b, CF08b, CW02, DP02, DFR06, DK04, DL01, DLS01, DL05, DGG07, Ebr06, Fin02, FD03, FR07, FSW06, GS07, GWS07, GS04, GM06, GWC04, HP01, HK01, HS01, HA03, HWC04, HSD06, HV08, ICL01, IWLA03, IZ00, JHS00, JZ01, JW05, JK07, Kel04, KTV06, KN05, KCS04, LS03a, LS03b, LB06, LN01, LSW06, LL00, LLW01, LL05, LL08a, LPL08, LWC07, LTR^{+}08, LWZ08, LL02, LR01, Lin00, LK07, LY04, LT06, LR00, MKB03, MCW06, MN03, Mar02, Mar03, MC00, PK01, PY03, PL01, PCK06, Pou00, QI06, Rao04, RP04, RVM08, RS01, RR04, RJT00, SSM05, Sar03, SR04, SN00, ŞM08, Sev04, Shi01]. Models [SK05, SW05, Tar02, TM04, VB05, VRR07, Vid01, WR02, WL03, WLS06, WO04, WCK03, XZT04, YY01, ZL06, Zha01, ZLWZ01, vdLA00, AKM05, BS09, CTSW08, CS09, CT06, Che09, DDS03, DM09, Dun09, FJ09, GM04, HWZ02, KMT03, KCD08, KB03, KF09, LM09, LR09, PR08, SR05, ŞM09, SMLG09, Tan09, Tsa09, WW09, Wes08, ZCI09]. Moderate [FH04]. Modified [Che01, Sar03, Sev02, Sev04]. Moment [GL08, YJS02]. Moments [YC03, LY09a, LY09b]. Monitoring [JVL05, TM03]. Monotone [LZH07, CFVG09]. Monotonicity [GHJK00]. Monro [Jos04]. Monte [AKM05, HJPR02, IZ00, JSH07, JK07, LR01, LT05, LS00, MPRB06, PR08, ZN00]. Motion [BBO02]. Move [WO04]. Movement [Riv01]. Moving [LL08a]. MR0368317 [He09]. MR1603936 [MO05]. MR1603940 [WO04]. MR1782479 [MBM05]. MR1844846 [KM05]. MR1859408 [Jef03]. MR2006837 [QZ05a]. MR2050457 [SR05]. MR2158619 [DRS12]. MR2202653 [HNPE07]. MR2234187 [GG06a]. MR2507144 [LY09a]. MR2507147 [HMXZ09b]. MR2767288 [Pal10]. Multi [AKR03, DKLW08, GWC04, HY08, Lin08, PDD06, Rei07, WO04]. Multi-Component [Rei07]. Multi-Dimensional [DKLW08, Lin08]. Multi-Factor [GWC04]. Multi-Level [PDD06]. Multi-Move [WO04]. Multi-Parameter [HY08]. Multi-State [AKR03]. Multicategory [LKLK06]. multigrid [LS00]. Multimodality [CT04, DR04]. Multinomial [SN00]. Multiple [ACH00, ACHM02, BBM00, CSXB08, GY03, HOR00, KB02, MB05, MSV07, Pad02, PY08, Rao04, Rei07, Rei08, TL04, TDM02, YY01, BRVW09, KC09]. Multiple-Imputation [GY03]. Multiplicative [MS02b]. Multiresolution [FF07]. Multiscale [KN05]. Multistate [BS08]. Multitaper [Wal00]. Multitype [BBL04]. Multivariate [Ano07a, BRL02, CFLZ05, CHT06a, CHB06, CW02, DG04, FD03, HNPE05, HR02, JL04, Kom01, Kuk04, LNO07, LLW01, LGS08, Mat06, Pou00, Qaq03, QI06, RS07, SC04, ST03, TLR03, TPP00, Wal00, YC04, YC07, HNPE07, ZCI09, ZN00]. mutations [BS09].

N [Jef03]. Natural [AS01, GM04]. Nearest [HA03]. Nearest-Neighbour [HA03]. nearly [BST09, LMT09]. negative [DW09]. Neighbour [HA03]. Nested [BW07, DK04, Qia09]. Networks [KAI04, PB01]. Neural [PB01]. Neutral [ELP03]. Neutral-to-the-Right [ELP03]. no [GG06a, Jef03, KM05, Kos01a, MBM05, MO05, QZ05a, SR05, WO04]. Noise [LR00]. Non [DK04, JK07, LCM09, MDC02, RR04, Sev07, TLZW04, WO04]. Non-Bayesian [Sev07]. Non-Compliance [RR04]. Non-finite [LCM09]. Non-Gaussian [JK07, WO04]. Non-Orthogonal [MDC02]. Non-Sequential [DK04]. Non-Smooth [TLZW04]. Nonadaptive [JT06]. Noncentral [WW06]. noncompliance [CSTT09]. nondecomposable [AKM05]. Nonidentifiability [Bet00]. Nonignorable [BRS05, ICL01, TLR03, VRR07]. Noninformative [Bet00]. Nonlinear [AAD00, HP01, JK07, LS03a, LS03b, RS01, SCS07, SND05, KF09, LM09, LR09]. Nonlinearity [TC00, TC02]. Nonmonotone [VRR07]. Nonnegligible [Tan06]. Nonnested [Fin02]. Nonnormal [CI05, CH02b, SCS07]. Nonparametric [AAD00, Ano06a, Ano07a, BRN08, BS08, CT05, CW08, CMMY09, CL00, CF08a, DM00, DAV03, DP08b, Dun09, GG05, GG06a, GF00b, HRR00, HHF04, HNPE05, HL06a, HL06b, HNPE07, Hug06, JvdLH03, KKPS04, KN05, KP04, LS03b, LMP07, LWWC04, LT09, MBH00, MF09, OO07, Pad02, PF06, PF07, PK04b, PL01, Sho02, SF03, Tar02, TW05, Wan03, WCMR07, WJT02, WP03, YL08b, ZLZ01, Zha06, ZN00, vdLA00, Bra03, CSTT09, RDG09]. Nonparametrics [Hal01]. Nonregular [But03a, Sev04, Xu05]. Nonresponse [CK08, Pad02, TLR03, VRR07, KR09]. Nonstationary [Fue02, LLW01, MC00, ND02, Rao04, LL08b]. Nordstrom [Xu05]. Nordstrom-Robinson [Xu05]. Normal [BB01, HLPL06, Kel04, Kom01, LPL08, LMR01, MH07, Pou00, SND05, Ver00, Jef03, WW09]. Normalising [KW05, MPRB06]. Normality [YJ00]. normally [Hal00]. Note [BS08, CHT06a, CR04, Kag01, Kim02, LWZ08, LWY06, NCT05, PLF05, PQP08, Sam03, TC02, VV05, Ver00, WS01, ZF02, Zou08, Guo09, He09, Jef03, JF09, Moo09, Pal09, PWB08, SZK09, Tan09, Uek09]. Nuisance [CC06b, Dav02, Han03, HE04, RS05, Sar03]. Null [MB05]. Number [CT01, LMR01, MB05, Jef03].

Objective [CS09, FFM08, Swe01]. Objects [KDA00]. Observables [Kom01]. Observation [HWZ06, DDS03]. Observation-driven [DDS03]. Observational [JT05, Ros01, Ros04, Ros06]. Observations [AKR03, CTW05, Mar00, Vid01, Hal00]. Observed [DFR06, LR00, Pad02, RSG07, RS01, TC00, TC02]. Obtain [CSW02, LOS07]. Occurrence [ST05]. Odds [DK03]. Off [HH07, AB01]. One [AB01, BJ02, BDS06, HN06, Kel04, SW05, TL04, Ste06, TC01]. One-Factor [Kel04]. One-Sided [BJ02, BDS06, TL04, Ste06]. One-Way [BDS06]. Only [Dav02, RS05]. Onset [WCMR07]. Operating [CC02, HHF04, QZ03, JP09a, QZ05a]. Optimal [BK06, BW07, BJ02, BRVW09, But01, But06, CB07b, CT09, DT09, KFD01, MDC02, NT08, UTW05, Wu03, YC07, JG09, LM09]. Order [BM05a, FM06, HP01, MR01, PDT05, Ros08, dRG01, Bra03]. Order-Restricted [HP01, PDT05]. Ordered [SC04, TS03, He09]. Ordering [CT05, DP08b, HL02a]. Orderings [Hof03]. Ordinal [CF01, SK00]. Orthodox [DL05]. Orthogonal [BST09, But01, But05, CI05, MDC02, SL06a, SL06b, Tan06, LMT09, SLL09]. Orthogonality [Pou07]. Other [HS01]. Outbreaks [BB04]. Outcome [Che00, RR04, SZK09]. outcome-dependent [SZK09]. Outcomes [HCR06, SR01a, VRR07, KC09]. Outlier [BPR02]. Outliers [TPP00, AR02]. Output [BSM06, KO00]. Outputs [OO02]. Overdispersion [JS02]. Overestimation [CC02]. overlap [CHIM09]. Overrunning [HL02a]. Overstratification [DC08].

Pages [AB01]. Paired [ZHC08]. Pairwise [Gli07, Kuk07, TM08, NGCD09]. Panel [CW00a, FLF02, HWZ06, LZH07, SF03, Zha02, HLL09, Zha06]. Paradox [HE04]. Parallel [CW02]. Parameter [BBL04, BC02, CM01, Dav02, DM01, FSW06, HY08, IZ00, KAI04, MO97, RVM08, UTW05, WLT07, XL03, HR09, MO05]. Parameter-Driven [FSW06]. Parameter-Redundant [CM01]. Parameterised [HS01]. Parameters [CC06b, Dat05, Fra03, GTW04, Han03, HE04, LH01, Pou07, RS05, Sar03, Swe05, Whi04]. Parametric [AF06, CHP00, DY08, MR01, Qin00, SXZ08]. Partial [HS09, Hof03, LCT07, LTR^{+}08, Sch03, SIC03, ZF02, DT09, LT09, LTX09, Tsa09]. Partially [DFR06, HWC04, LLW01, LWC07, MCW06, Pad02, RSG07, RS01, TC00, TC02]. Particle [Cho02, GF00b]. partition [Dun09]. Partner [vdLA00]. Path [SW05, Zou08]. Path-Based [Zou08]. Pattern [KMT03]. Pattern-mixture [KMT03]. Penalised [BCO05, CRCW05, HO05, HLPL06, Lie01]. Penalized [LR08, Zou08, CKO09, WNZZ09]. Performance [WC03]. Period [HL06a, HL06b, KNN08a, KNN08b]. Periodic [HRR00]. Periodogram [HL06a, HL06b, OMS04, ORSV01]. permanent [KM09]. Permutation [HT02, Kuk04, AEB07]. Persistent [OMS04]. Personal [Ton01]. perspective [Fat06]. Perturbing [JYW01]. Peto [HN01]. Pitt [WO04]. Pivot [Ros01]. Pivotal [GTW04]. Pizzetti [Hal00]. Plague [SCS07]. Planar [KDL07]. Plans [MDC02]. Plant [GJA07]. Plant-Capture [GJA07]. Plausible [CK08]. plot [CT09, JG09]. Point [BY08, LR00, MS07, MU01, Waa08, CFVG09, GC02, Gua09]. points [HSW00]. Pointwise [CL08]. Poisson [BSCC01, DDW00, DDS03, Gua08, MKB03, ML02, RSG07, Shi01]. Poles [CH00a]. Polya [ZHC09]. Polygons [LWY06]. Polynomial [CJ05, LZH07, ONL01]. Population [BE01, GJA07, HL05, HV08, HH05, JSH07, KB01, KC06, ZR00, CTD09, FJ09]. Population-Based [JSH07]. Positive [CF01, FK07, NCGD03]. Possibly [JT05, CTSW08]. Post [WLS06]. Post-Strata [WLS06]. Posterior [CIS06, FM06, SDH04, YJS02, PB02]. Power [CTW05, GH05, LS04, Ren02, SSM05, YJ00, Bra03]. pp [HNPE07]. Practical [HR02]. Predictability [MN07]. Predicting [CTSW08, KO00, SZG08]. Prediction [BY08, LMKP07, LF05, TCGW07, UF07, Vid01, HP09]. Predictive [And07, Kom01, LF05]. Predictor [DL05]. predictors [LR09]. Prepivoting [LY03]. Presence [BRL02, BRN08, CC06b, KC06, SR02, Tar02, WD08]. Present [Dav02, RS05]. Prevalence [HOR00]. Principal [BPR02, Boi03, CH00b, CHN03, Ger08, JHS00, Lue04, Sch03, ZHC08, Zhu06]. Principle [Swe01, Pal09, Pal10]. Principles [Che05]. Prior [OO07, PB02]. Priors [BGD07, CL06b, ELP03, GL08, LC03, MR01, SMG02, SS03, Swe05]. Probabilistic [SHD02, Pal09, Pal10]. Probabilities [Ber07, Fat06]. Probability [GH05, LMP07, MR01, SZG08, SDH04, SMG02, Swe01, Swe05, WSL08, Bin00, BB03, GG06b, Hal00, MF09]. Problem [BLP00, CT05, ML02, WWM01, ZL05]. Problems [CC07, HQ05, Sie04, YLK07, HS09, TG07]. Procedure [Ano04a, Rød06, SC96, SKP00, ZZ04, Ful09, Jos04]. Procedures [BKY06, BM05b, CHP00, CSXB08, DY08, QZ03, TL08, Gua09, Guo09, KL09, QZ05a]. Process [BM05a, CP05, DGG07, FYK04, HY03, IZ00, KTV06, RSG07, CGOO09, PR08]. Processes [BBO02, Bla03, BSCC01, CHT06a, DP08a, FF07, Fue02, GS07, Gua08, Lie01, LR00, MC00, OMS04, Sho02, TH05, Waa08, ZL06, GC02, Gua09, WD00a]. Procrustes [KM05, KM01]. Product [FK07]. Profile [DM01, HWC04, Sar03, WH03, Tan09]. Profile-Kernel [HWC04]. Profiles [CC06a]. Prognostic [BB01, Han08]. Projected [HE04]. Projection [Ren02, Tan01]. Projections [KM01, KM05]. Projective [TGM00]. Prone [LWC07]. Proof [NCGD03]. Propensity [Han08, Imb00]. proper [KMT03]. Properties [DY08, OMS04, RW02, ST03, Xu05, CKO09, Ful09]. Property [OGC07, SMG02]. Proportional [BK06, DK03, DL01, Ebr06, GH05, KFD01, Lin00, PK01, TD01, WS01, YY01, Zou08, O’Q03, Tsa09, ZL07]. Proportions [GF00b, TS03]. Propriety [CIS06]. Prospective [GZH06, SR04, SR05]. Protein [GM06]. Pseudo [AKR03, GZH06, LTX09, ONL07, Tsa09, LT09]. Pseudo-Observations [AKR03]. Pseudo-partial [LTX09, Tsa09, LT09]. Pseudolikelihood [CR04, SB07, Zha02]. Pursuit [Ren02].

Quadratic [AS01, BP07, QLL00, QS04, GM04]. Qualitative [AH00]. Quality [AM06]. Quality-of-Life-Adjusted [AM06]. Quantile [BS08, DCNZ05, Kos99, Kos01b, CT06, Kos01a, PF07]. Quantiles [HL05]. Quantitative [Sie04]. Quasi [FD04]. Quasi-variances [FD04]. Quasilikelihood [BBM00].

R [Jef03]. Radial [KAI04]. Random [BB02, CJ06, CHT06b, CHB06, CH02b, Cox03, DK03, HH02, LL06, LN01, LGS08, MKB03, PGP01, RVM08, TWG07, UTW05, WL03, ZR00, Dun09]. Random-Effect [LN01]. Randomisation [WT06]. Randomised [Cuz01, RR04]. Randomization [CL08]. Randomized [CB07b, CSTT09]. randomness [GC02]. Range [CJ06, CSW02, LMP01]. Rank [BW05, FW08, LT08, Nie01, WZ06, Zho05, AEB07, GK04, JLWY03, JY03]. Rank-Based [WZ06, JLWY03, JY03]. Rare [Hel00, HOR00]. Rate [BKY06, LZ08, SS00, XL03, KL09, Sun02]. Rates [CHT06b, Llo00]. Ratio [CT05, CF08a, CRCW05, DCNZ05, FZ04, LBS02, RS05, Sev00, Sev04, XL03, Sha09]. Ratios [FK07, YP05]. rearrangement [CFVG09]. Recapture [BE01, CL00, HH05, KB02, Tar02, WLS06]. Receiver [CC02, HHF04, QZ03, JP09a, QZ05a]. Reconciling [ZZ05]. Records [Rei08]. Recovering [KC06]. Recurring [Ebr06]. Recursions [RP04]. Recursive [BB02, FR07]. Reduce [Han03]. Reducing [AH00, CF08b, HR09, PLF05]. Reduction [CLC07, DT04b, Li07, YC07, KF09, Peñ09, Xia09]. Redundant [CM01]. Reference [CL06b, PB06]. Regenerative [HJPR02]. Region [UTW05]. Regions [Bot03]. Regression [ACH00, AF06, BRT01, Bon05, CWW00, CC04, CP02a, Car05, Che01, CJ01, CSW02, CJ05, CC05b, CC06c, CIS06, CF01, CN06, CLC07, CC02, DDW00, DM00, DK04, DL01, DM02b, Dow03, FYK04, GT05, GS01, GHJK00, GH05, HO05, HL06a, HL06b, HZ00, JZ07, JLY06, Kuk04, KL00, LX05, LL04, LSW06, LCT07, LC01, LWWC04, Llo00, Lue04, MS02a, Mey03, NYK06, NCT05, OR00, ONL01, ONL07, PCK06, Pre07, QZ03, Rao04, SC04, SZG08, ŞM05, Shi01, SK05, SXZ08, TW05, VRR07, Wan03, WZ06, WJT02, WWM01, XZ07, Yan05, YC07, YL08b, Zha01, ZFH08, CMMY09, FFM08, HL09, Han09, HSW00, JY03, LY09a, LY09b, LTX09, QZ05a, SMLG09, WNZZ09, Xia09]. Regressions [DR04, YC05]. Regular [FM00, CT09, GC02]. regularization [WSG09]. Regularly [Mar00]. Regularly-Spaced [Mar00]. Rejection [CBD02, GM01]. rejective [Ful09]. Related [JHHB01]. Relatives [CC06a]. Rendering [CHP00]. Repeated [CYCJ07, VRR07, WZ06, DT09, HWZ02, JY03, PWB08]. repeatedly [CMMY09]. Representation [AMMC07, HW00]. Representations [WD00a]. Resampling [HSW01, JYW01, LMKP07]. Resampling-Based [LMKP07]. Residual [CC05b, TW05, JF09]. Residuals [Car05, GWC04, SW03, SN00, Vic00, WS01]. Response [CF01, Imb00, JT05, Kur07, LWC07, KCD08]. Responses [BB01, ICL01, WD08, CTSW08]. Restoration [Gib00]. Restoring [Sam03]. Restricted [CSW02, HP01, LL02, Mey03, PDT05]. result [Cox07]. Results [But03a, But05, LM09]. Retrospective [NSW02, PR08, SR04, SR05]. Reversibility [CHT06a]. Reversible [BSM06, GM01, JSH07]. Revisiting [LL04]. Richly [HS01]. Richly-Parameterised [HS01]. Right [ELP03, JT05, SMG02, VYZ01, LT09]. Right-Censored [JT05, LT09]. Right-Invariant [SMG02]. Risk [BRL02, BRN08, HP09]. Risks [BS08, CF08a, CD04, FJC01, GT05, JvdLH03, Lee06, PF06, SA00, PF07]. Robbins [Jos04]. Robbins-Monro [Jos04]. Robinson [Xu05]. Robust [AST05, BPR02, Che00, CHP00, CH00b, CHN03, Ger08, GMR08, JZ01, LH06, ONL07, Pan01, Tan06, TGM00, XL03, dRG01, CTD09, Moo09]. Robustness [HSD06, QS04, CTD09]. Role [Imb00, ONL07]. Row [BW07]. Row-Column [BW07]. Rubin [Jef03].

Saddlepoint [DW02, DM09, KW05, AEB07, Ter03]. Sample [AMMC07, Ano06a, Bea08, CT05, CSB03, FW08, GG05, Kos99, Kos01b, LS04, OMS04, Rei07, SR01a, Shi01, Smi01, VYZ01, YP05, ZLZ01, ZL05, CH09, DJ00, GG06a, GDRG09, Kos01a, YSZ06, Zha06, GK04]. Sample-size [GK04]. Sampled [TWG07]. Sampler [Gib00, HYE07, LOS07, WO04, LS00]. Samples [Ber07, LSW06, TF04]. Sampling [BY08, CBD02, DT04a, FMT04, FSW06, FH04, HYE07, NT08, NK05, PDD06, RW02, TCGW07, Wu03, ZR00, Ful09, MF09, SZK09, Tsa09]. Scalar [DM01, DM02a]. Scale [CH09]. Scaled [WS01]. Scales [AF06].

Scaling [DKLW08, Gow04]. Schoenfeld [WS01]. Schwarz [KB03]. Score [DL01, HP01, Han08, Imb00]. search [AR02]. Seasonally [OMS04]. Second [MR01, Bra03]. Second-Order [MR01, Bra03]. Secondary [HY03, LH01, LSS09]. sectional [MF09]. Secure [BBL04]. Seemingly [DR04]. Segmentation [BBM00]. Selecting [QLL08]. Selection [CFLZ05, CC08, CC07, DFR06, DP04, FD03, GF00a, HLPL06, KTV06, KX07, KAI04, KC06, NK05, ORSV01, Sie04, VV05, WCK03, YL07, Zou08, AR02, CS09, HMXZ09b, HMXZ09a, PB02, Uek09]. Selections [NT01]. selector [JR09]. Selectors [WLT07]. Semi [FJC01]. Semi-Competing [FJC01]. Semicompeting [PF06]. Semiparametric [CWW00, CC04, CTW05, CP02a, CC05a, Che01, CJ01, CJY02, CC05b, CW00a, CC07, GT05, GZH06, Gua04, HZF02, HB06, HP09, JT05, KCS04, LX05, LPL08, LC01, LJ03, LMP01, LY04, LT06, MCW06, MH07, RSC02, TD01, TM04, WT06, WD08, XZ07, YP05, ZL06, Zha02, ZL05, ZFY02, Che09, JS09, SZK09]. Semivarying [XZT04]. Semivarying-Coefficient [XZT04]. Sensitivity [Han03, Ros04]. Sequence [BBM00, CL06a]. Sequential [BJ02, Cho02, DK04, HL02a, HY03, JT06, LS04, LL06, LH01, DJ00, LSS09, PWB08]. Sequentially [SC04]. Serial [DM00]. Serially [HS03]. Series [CGR04, DK02, FD03, FSW06, HW00, LL05, LL08a, Mat06, MSV07, PY08, SCS07, Ton01, TC00, TC02, TPP00, WO04, CL09, DW09, Nor09, Peñ09]. Set [PB06]. Sets [BBO02, FM06]. Settings [HT02, CH09]. Shape [ALS03, DKLW08, KM01, KM05, KDL07, Mey03]. Shape-Restricted [Mey03]. Shape-Space [KDL07]. Shapes [HZ00, MBM00, WL03, MBM05]. Shared [GZH06, RVM08, Yan05]. sharpness [Pal09, Pal10]. Shephard [WO04]. Short [TH05, YP05]. Short- [TH05]. Short-Term [YP05]. Shrinkage [AS01, BD06, Kom01, NCT05, WW06, JR09]. Sided [BJ02, BDS06, TL04, Ste06]. Sign [LNO07]. Significance [Rei07]. Simes [Ano04a, Rød06, SC96]. Similar [ES01]. Simple [BRT01, CH02b, DK02, JYW01, LL04, ORSV01, Ste06]. Simulating [Qaq03]. Simulation [CMW07, DK02, FR07, GS07, HJPR02, LB06, Mar00]. Simulation-Based [LB06]. Simulation-Free [FR07]. Simulations [LTLE03]. Simultaneous [HN00, NP08]. Single [KX07, LK07, NT01, Vic00, YC05]. Single-Factor [Vic00]. Single-Index [KX07, LK07, NT01, YC05]. Sinh [JP09b]. Sinh-arcsinh [JP09b]. Size [AMMC07, BE01, CSB03, FW08, GF00b, GJA07, HH05, KB01, LS04, Shi01, CH09, DJ00, FJ09, GK04]. Sizes [BRS05, BY08, LL06]. Skew [MH07]. Skew-Normal [MH07]. Skewness [NCGD03]. Sliced [NCT05, Pre07, QW09, SCW07]. Small [BB04, DRS05, DRS12, GMR08, LMKP07, Rei07, YL08a, CT06, LR09, GM04]. Small-area [GM04]. Small-Sample [Rei07]. Smooth [Del06, DCNZ05, TLZW04, Uek09]. smooth-threshold [Uek09]. Smoothed [BHY01, BW05]. Smoother [DK02]. Smoothers [CQ00, DW02]. Smoothing [KAI04, KDL07, LWWC04, MC00, ONL07, ORSV01, PSH06, RS07, SS03, CH02a, HR09, JS09]. smoothing-parameter [HR09].

Smoothly [WLT07]. solution [WSG09]. Some [But03b, CW02, Ful09, LM09, Sev04, Xu05]. Space [DK02, FSW06, HW00, JK07, KDL07, Lin08, vLMZ01, NGCD09, QW09]. space-filling [QW09]. Spaced [Mar00]. Spaces [CC08, SXZ08]. Span [ORSV01]. Sparse [JHS00, Li07, ZHC08]. Spatial [DGG07, Fue02, Gua08, ND02, ST03, Waa08, ZZ05, GC02, Gua09, ZHC09]. Spatially [BSCC01, Lin08, PSH06, WJT02]. Species [LMP07]. Specific [CF08a, PW03, JF09]. Specified [BW07, Qaq03, WI08, CTSW08]. Spectra [HN00, RS07]. Spectral [Boi02, BD06, CF08b, DG04, FZ04, Fue02, NP08, Wal00, Sha09]. Spherical [Dow03, Ger08]. Spline [HO05, SS03, CKO09]. Splines [BCO05, CRCW05, DGK01, KDL07, LR08, LWWC04, LZH07, PSH06, RS07, WJT02, ZFH08, CH02a]. split [CT09, JG09]. split-plot [CT09]. split-split-plot [JG09]. Squared [Sch03, LR09]. Squares [Dem00, JLY06, TW05]. Stage [BM05b, BY08, CL00, FW08, WT06, SZK09]. Standard [BW05]. Standardised [MBH00]. State [AKR03, DK02, FSW06, HW00, JK07, Sho02]. Static [Cho02]. stationary [FSR06]. Statistic [LT05, Mat06, SR01a, Sev04]. Statistical [GTW04, MBM00, MBM05, Riv01, TLZW04, Xu05, Hal00]. Statistics [FM05, ZZ05, Bin00, BB03, GK04, GG06b, Hal00, Sha09]. Status [Bet00, JvdLH03, JV04, JVL05, LX05, MS02a, TC06b, WD00b, vdLA00]. Stein [TH05]. Step [BKY06, HH07]. Step-Down [HH07]. Step-Up [HH07, BKY06]. Stick [DP08a]. Stick-Breaking [DP08a]. Stochastic [BBL04, DP08b, GS07, Hof03, KP04, LZ08, LTLE03, Mar01]. Strata [KB02, WLS06]. Strategies [FW08, NT08]. Stratified [Ber07, ZR00]. Stratum [Sar03]. Strengths [Yan05]. Structural [Ano06a, GG05, HSD06, RW02, RR04, GG06a]. Structure [HH02, HZF02, LGS08, ND02, QLL08, WC03]. Structured [LN01, LKLK06, MN03]. Structures [MB05, PM03, YP06]. Student [FFM08, WCK09]. Student- [FFM08]. Studentization [FR08]. Studies [AST05, Bin00, Bon07, BB03, CC05a, Che05, GG06b, Hal00, JT05, KL00, Kur07, NYK06, NSW02, PM03, PK04a, PGP01, RSC02, Ros01, Ros04, Ros06, SR01b, SR04, vdLA00, HP09, KC09, SR05]. Study [LT06]. Subject [AM06, PW03, Wan03]. Subject-Specific [PW03]. Subspace [CN06, Zen08]. Subspaces [Sch03, YC03]. Successive [AM06]. Sufficient [Li07, LT05, RF03, VR08, Wal03, YC07]. Sum [OGC07]. Summary [Cox08]. Superiority [TL04]. Superiority-Equivalence [TL04]. Supersaturated [CT01]. Support [JL04, LKLK06]. Supremum [CBD02, FW08]. surface [WSG09]. Surrogate [Che00]. Survey [Lin00, SK00, Wu03, GM04, GDRG09, KR09]. Surveys [Bea08, BCO05, CSW02, Smi01]. Survival [CTW05, CRA06, CC05b, GZH06, LL06, LL00, LPL08, MN07, Oak01, PW03, PK01, PH07, TDM02, WT06, YP05, GK04, ZCI09, ZHC09]. Switching [LR01]. Symmetric [Car05]. Symmetrical [But05]. Symmetry [KDA00, KM01, MBM00, YJ00, KM05, MBM05]. Synchronization [TM08]. Synthesis [LN01].

tailed [WCK09]. Tangent [KM01, KM05]. Tapered [Nor09, PP01]. Tau [Oak08]. Technique [DT04b, FSR06]. Techniques [MC00]. Temporal [FYK04, PH07]. Term [YP05]. Test [BM05b, CKS03, DM00, FW08, Gua08, HN01, LNO07, LH01, LT08, Mat06, Mey03, Nie01, SF03, TC06a, Zha01, ZLZ01, DJ00, GC02, YSZ06]. Testing [ACH00, BGD07, Bon07, CT04, CSXB08, CL08, DK03, Dav02, FM01, FLF02, IWLA03, Jef03, LGS08, LMR01, MB05, Mei08, QS02, Ros08, Sch05b, TS03, TC00, TC02, TDM02, BRVW09]. Tests [BL03, BJ02, CT05, CRCW05, DL01, FZ04, GHJK00, HP01, HT02, HL02a, JT06, Kos99, Kos01b, LSS09, LBS02, LOS07, QLL08, Rei07, RS05, SSM05, Sch03, SCW07, TL04, VYZ01, dRG01, AEB07, AR02, Bra03, CH02a, Kos01a, Zha06, ZN00]. Their [HY08, Xu05]. Theory [But03b, CC07, Dav01, HO05, Tan01]. There [HN06]. Third [YC03]. Thompson [Fat06]. Three [BDS06, FK07, Riv01, TGM00]. Three-Level [TGM00]. Threshold [BBL04, IWLA03, SCS07, Uek09]. thresholding [OW04]. Thresholds [AF06, IWLA03]. tilt [Tan09]. Tilted [Sch05a]. Tilting [CHP00]. Time [AM06, BL03, BRL02, BRN08, BRT01, CWW00, CFLZ05, CHT06a, CRW02, CGR04, DK02, FJ00, FD03, FSW06, Gib00, Gli07, HW00, LL05, LL08a, Mat06, MSV07, PY08, PW03, SCS07, SR02, SC04, TC06b, Ton01, TC00, TC02, TPP00, THW05, WO04, YY01, YC04, YL08b, Zho05, CL09, DW09, He09, JLWY03, JS09, KMT03, LY09a, LY09b, Nor09, Peñ09]. time-ordered [He09]. Time-Reversibility [CHT06a]. Times [Cuz01, HWZ06, JVL05, Lee06, ST05, Str05, ZCI09]. TM [EL01, Sun02]. Tolerance [MR01]. Tool [SW03]. Tracking [WLP05]. Tractable [CHB06]. Trait [Sie04]. Traits [HOR00]. Transform [HQ05]. Transformation [CTW05, Che01, CJY02, GT05, KCS04, LPL08, LY04, LT06, ZL06, Che09, Ter03, ZCI09]. Transformations [FJ00, Lee06, YJ00]. Transplant [SK05]. Treatment [BM05b, CRW02, CYCJ07, FW08, Kur07, ONL01, RF03, RR04, Ros01, TDM02, ZL05, CHIM09]. Treatment/Response [Kur07]. Treatments [BDS06, KFD01]. Treed [WLS06]. Trees [CL06a, ZHC09]. Trial [CSB03, HY03]. Trials [BM05b, CS05, CB07b, Cuz01, LS04, LL06, RR04, TL04, TM03, WT06, CSTT09, LSS09]. Trojan [Edm02]. True [PK04b]. Truncated [PF06]. Truncation [CRA06, Del06]. Tuning [WLT07, JR09]. Two [BM05b, BY08, BDS06, But03a, But06, CT05, CL00, FM00, FW08, IZ00, Kos99, Kos01a, Kos01b, SHD02, Tan06, VYZ01, WT06, YP05, ZLZ01, ZZ05, ZL05, CT09, GC04, GDRG09, JAW09, SZK09]. two- [JAW09]. Two-Factor [Tan06]. Two-Level [But03a, But06, FM00, CT09]. Two-Parameter [IZ00]. Two-Sample [CT05, Kos99, Kos01b, VYZ01, YP05, ZLZ01, ZL05, Kos01a, GDRG09]. Two-Stage [BM05b, BY08, CL00, FW08, WT06, SZK09]. Type [CM08, EC02, FM06, JH07, LL02, Che09, HSW00, Wan09, O’Q03].

Unbiased [DL05, LH01]. Uncertainty [CP02b, CDP08, OO02, OO07, TGM00]. Unconditional [DY08]. Understanding [Hug06, WLP05]. Undirected [JW05]. Unequal [Ber07, HN01]. Unified [Wal00, KR09]. Uniform [HL02b, RSSW03]. Uniformity [FM00]. Unistage [Ber07]. Unit [JH07]. Univariate [JVL05]. Unknown [BY08, Lee06, Ver00]. Unobserved [AV07, Kur07]. Unrelated [DR04]. Unspecified [HZF02]. Updating [Joh01, WJ00, DJ00]. Upper [CT01]. Use [GDRG09, LOS07, TS03]. Used [MBH00, Rei08]. Using [AH00, BCO05, CB07a, CK08, CSW02, CSB03, Cla03, CN06, CL08, DG04, FF07, Ger08, GM06, HL06a, HL06b, Hel00, HH05, KDA00, LZ08, LZH07, LT08, Mey03, QZ03, QZ05a, QLL00, RS01, RR04, TW05, WL03, YL08b, YLK07, ZR00, ZHC08, AEB07, GK04, JP09a, MKHT09, Uek09, WSG09].

Vaccination [BBL04]. Vaccine [BB04]. Valid [Rød06]. Validation [WR02]. Validity [FM06]. value [GRW06]. Values [FR08, ST03, AEB07, PWB08, TG07]. Variability [AH00, BKW05, DRS05, DRS12, HR09]. Variable [BP07, CFLZ05, GF00a, KTV06, KX07, Kur07, Mei08, NK05, SW05, Zou08, AR02, HMXZ09b, HMXZ09a, TG07, Uek09]. Variables [CJ06, CF01, CW02, Cox03, FK07, LWC07, OZ01, Qaq03, SHD02, WNZZ09]. Variance [AS01, Ber07, CW08, Joh01, LKLK06, LL02, MBH00, MU01, Pan01, PLF05, Pou07, Pre07, SCW07, TW05, Ver00, XL03, GM04, Gua09, KR09, Moo09]. variances [FD04]. variates [GC04]. Variation [ALS03]. Variational [vLMZ01]. Varying [CYCJ07, ŞM08, HWZ02]. Varying-coefficient [HWZ02]. Vector [DM02a, HN00, LKLK06, RW02]. Version [GY03]. Versions [Mar01]. versus [DL05, HWC04, Mey03]. Via [HYE07, BBM00, FM01, GTW04, HLPL06, KTV06, PY08, RS07, WW06, WLS06, Xia09, YJS02, YC03]. View [BY08, Wal00]. Volatility [GS07, FSR06]. Volume [Ano00i, Ano01i, Ano02i, Ano03i, Ano04j, Ano04k, Ano05i, Ano05j, Ano06j, Ano06k, Ano07j, Ano07k, Ano08i, Ano08j, Ano08k].

W. [Cox07]. Wavelet [AS01, AF06, ONL01, ONL07, SDH04, WW06, HL09, OW04]. Wavelets [IWLA03]. Way [BDS06]. Weibull [MO05, MO97]. Weighted [CP05, Che09, FW08, GF00b, KCS04, LNO07, LY03, Sch03, TL08, AEB07, GK04]. Weighting [Bea08, CK08, GRW06]. Weights [CSW02, GM04]. Westfall [CL08]. Westfall-Young [CL08]. Whittle [CGR04]. Wijs [BM05a]. Wilcoxon [Wan09]. Wilcoxon-type [Wan09]. Wilson [Ter03]. Wilson-Hilferty [Ter03]. Wishart [CMW07, Rov00]. Within [HSW01, Wan03]. Within-Cluster [HSW01]. Within-Subject [Wan03]. without [CLC07, WD00a]. Working [WC03, CTSW08].

XLIX [GG06b]. XLVI [Bin00]. XLVII [Hal00]. XLVIII [BB03].

Years [AB01, Cox01, TC01]. Young [CL08].

Zero [Bot03]. Zeros [CDR07].

References

Akritas:2000:NMM

[AAD00] Michael G. Akritas, Steven F. Arnold, and Yunling Du. Nonparametric models and methods for nonlinear analysis of covariance. Biometrika, 87(3):507–526, September 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/3/507;
 http://www.jstor.org/stable/2673627.

Atkinson:2001:OHY

[AB01] Anthony C. Atkinson and R. A. Bailey. One hundred years of the design of experiments on and off the pages of Biometrika. Biometrika, 88(1):53–97, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673675.

Aerts:2000:TLF

[ACH00] Marc Aerts, Gerda Claeskens, and Jeffrey D. Hart. Testing lack of fit in multiple regression. Biometrika, 87(2):405–424, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/405;
 http://www.jstor.org/stable/2673472.

Aerts:2002:LMI

[ACHM02] Marc Aerts, Gerda Claeskens, Niel Hens, and Geert Molenberghs. Local multiple imputation. Biometrika, 89(2):375–388, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/375;
 http://www.jstor.org/stable/4140583.

Abd-Elfattah:2007:WLR

[AEB07] Ehab F. Abd-Elfattah and Ronald W. Butler. The weighted logrank class of permutation tests: P -values and confidence intervals using saddlepoint methods. Biometrika, 94(3):543–551, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/3/543;   http://www.jstor.org/stable/20441395.

Antoniadis:2006:PMT

[AF06] Anestis Antoniadis and Piotr Fryzlewicz. Parametric modelling of thresholds across scales in wavelet regression. Biometrika, 93(2): 465–471, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/93/2/465;   http://www.jstor.org/stable/20441296.

Azzalini:2000:RVU

[AH00] Adelchi Azzalini and Peter Hall. Reducing variability using bootstrap methods with qualitative constraints. Biometrika, 87(4):895– 906, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/87/4/895;   http://www.jstor.org/stable/2673617.

Atay-Kayis:2005:MCM

[AKM05] Aliye Atay-Kayis and Hélène Massam. A Monte Carlo method for computing the marginal likelihood in nondecomposable Gaussian graphical models. Biometrika, 92(2):317–335, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/2/317;
 http://www.jstor.org/stable/20441191.

Andersen:2003:GLM

[AKR03] Per Kragh Andersen, John P. Klein, and Susanne Rosthøj. Generalised linear models for correlated pseudo-observations, with applications to multi-state models. Biometrika, 90(1):15–27, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/1/15;   http://www.jstor.org/stable/30042016.

Albert:2003:ALI

[ALS03] Michael H. Albert, Huiling Le, and Christopher G. Small. Assessing landmark influence on shape variation. Biometrika, 90(3):669–678, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/669;   http://www.jstor.org/stable/30042074.

Andrei:2006:EQL

[AM06] Adin-Cristian Andrei and Susan Murray. Estimating the qualityof-life-adjusted gap time distribution of successive events subject to censoring. Biometrika, 93(2):343–355, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/343;
 http://www.jstor.org/stable/20441286.

Ahn:2007:HDL

[AMMC07] Jeongyoun Ahn, J. S. Marron, Keith M. Muller, and Yueh-Yun Chi. The high-dimension, low-sample-size geometric representation holds under mild conditions. Biometrika, 94(3):760–766, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/3/760;   http://www.jstor.org/stable/20441411.

Ando:2007:BPI

[And07] Tomohiro Ando. Bayesian predictive information criterion for the evaluation of hierarchical Bayesian and empirical Bayes models. Biometrika, 94(2):443–458, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/2/443;   http://
www.jstor.org/stable/20441383.

Anonymous:2000:BMa

[Ano00a] Anonymous. Back matter. Biometrika, 87(1):??, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673579.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. Biometrika, 87(2):??, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673479.

Anonymous:2000:BMc

[Ano00c] Anonymous. Back matter. Biometrika, 87(3):??, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673647.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. Biometrika, 87(4):??, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673624.

Anonymous:2000:FMa

[Ano00e] Anonymous. Front matter. Biometrika, 87(1):??, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673556.

Anonymous:2000:FMb

[Ano00f] Anonymous. Front matter. Biometrika, 87(2):??, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673460.

Anonymous:2000:FMc

[Ano00g] Anonymous. Front matter. Biometrika, 87(3):??, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673625.

Anonymous:2000:FMd

[Ano00h] Anonymous. Front matter. Biometrika, 87(4):??, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673606.

Anonymous:2000:VI

[Ano00i] Anonymous. Volume information. Biometrika, 87(4):i–viii, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673605.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Biometrika, 88(1):??, March 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673688.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Biometrika, 88(2):??, June 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673506.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Biometrika, 88(3):??, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673459.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Biometrika, 88(4):??, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673714.

Anonymous:2001:FMa

[Ano01e] Anonymous. Front matter. Biometrika, 88(1):??, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673671.

Anonymous:2001:FMb

[Ano01f] Anonymous. Front matter. Biometrika, 88(2):??, June 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673480.

Anonymous:2001:FMc

[Ano01g] Anonymous. Front matter. Biometrika, 88(3):??, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673433.

Anonymous:2001:FMd

[Ano01h] Anonymous. Front matter. Biometrika, 88(4):??, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673690.

Anonymous:2001:VI

[Ano01i] Anonymous. Volume information. Biometrika, 88(4):i–viii, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673689.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Biometrika, 89(1):??, March 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140573.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Biometrika, 89(2):??, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140595.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Biometrika, 89(3):??, September 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140616.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Biometrika, 89(4):??, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140553.

Anonymous:2002:FMa

[Ano02e] Anonymous. Front matter. Biometrika, 89(1):??, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140554.

Anonymous:2002:FMb

[Ano02f] Anonymous. Front matter. Biometrika, 89(2):??, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140574.

Anonymous:2002:FMc

[Ano02g] Anonymous. Front matter. Biometrika, 89(3):??, September 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140596.

Anonymous:2002:FMd

[Ano02h] Anonymous. Front matter. Biometrika, 89(4):??, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/4140532.

Anonymous:2002:VI

[Ano02i] Anonymous. Volume information. Biometrika, 89(4):i–viii, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/4140531.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. Biometrika, 90(1):??, March 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042035.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. Biometrika, 90(2):??, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042059.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. Biometrika, 90(3):??, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042082.

Anonymous:2003:BMd

[Ano03d] Anonymous. Back matter. Biometrika, 90(4):??, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042106.

Anonymous:2003:FMa

[Ano03e] Anonymous. Front matter. Biometrika, 90(1):??, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042014.

Anonymous:2003:FMb

[Ano03f] Anonymous. Front matter. Biometrika, 90(2):??, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042037.

Anonymous:2003:FMc

[Ano03g] Anonymous. Front matter. Biometrika, 90(3):??, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042061.

Anonymous:2003:FMd

[Ano03h] Anonymous. Front matter. Biometrika, 90(4):??, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/30042085.

Anonymous:2003:VI

[Ano03i] Anonymous. Volume information. Biometrika, 90(4):??, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/30042084.

Anonymous:2004:AC

[Ano04a] Anonymous. Amendments and corrections: “Is the Simes Improved Bonferroni Procedure Conservative?”. Biometrika, 91(1):249–250, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/1/249;   http://www.jstor.org/stable/20441095.

See [SC96].

Anonymous:2004:BMa

[Ano04b] Anonymous. Back matter. Biometrika, 91(1):??, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441096.

Anonymous:2004:BMb

[Ano04c] Anonymous. Back matter. Biometrika, 91(2):??, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441117.

Anonymous:2004:BMc

[Ano04d] Anonymous. Back matter. Biometrika, 91(3):??, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441139.

Anonymous:2004:BMd

[Ano04e] Anonymous. Back matter. Biometrika, 91(4):??, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441158.

Anonymous:2004:FMa

[Ano04f] Anonymous. Front matter. Biometrika, 91(1):??, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441074.

Anonymous:2004:FMb

[Ano04g] Anonymous. Front matter. Biometrika, 91(2):??, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441098.

Anonymous:2004:FMc

[Ano04h] Anonymous. Front matter. Biometrika, 91(3):??, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441119.

Anonymous:2004:FMd

[Ano04i] Anonymous. Front matter. Biometrika, 91(4):??, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441142.

Anonymous:2004:VIa

[Ano04j] Anonymous. Volume information. Biometrika, 91(1):??, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441073.

Anonymous:2004:VIb

[Ano04k] Anonymous. Volume information. Biometrika, 91(4):??, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441141.

Anonymous:2005:BMa

[Ano05a] Anonymous. Back matter. Biometrika, 92(1):??, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441184.

Anonymous:2005:BMb

[Ano05b] Anonymous. Back matter. Biometrika, 92(2):??, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441207.

Anonymous:2005:BMc

[Ano05c] Anonymous. Back matter. Biometrika, 92(3):??, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441229.

Anonymous:2005:BMd

[Ano05d] Anonymous. Back matter. Biometrika, 92(4):??, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441252.

Anonymous:2005:FMa

[Ano05e] Anonymous. Front matter. Biometrika, 92(1):??, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441161.

Anonymous:2005:FMb

[Ano05f] Anonymous. Front matter. Biometrika, 92(2):??, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441186.

Anonymous:2005:FMc

[Ano05g] Anonymous. Front matter. Biometrika, 92(3):??, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441209.

Anonymous:2005:FMd

[Ano05h] Anonymous. Front matter. Biometrika, 92(4):??, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441232.

Anonymous:2005:VIa

[Ano05i] Anonymous. Volume information. Biometrika, 92(1):??, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441160.

Anonymous:2005:VIb

[Ano05j] Anonymous. Volume information. Biometrika, 92(4):??, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441231.

Anonymous:2006:AC

[Ano06a] Anonymous. Amendments and corrections: “Nonparametric Maximum Likelihood Estimation of the Structural Mean of a Sample of Curves”. Biometrika, 93(4):1025, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/20441345. See [GG05].

Anonymous:2006:BMa

[Ano06b] Anonymous. Back matter. Biometrika, 93(1):??, March 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441275.

Anonymous:2006:BMb

[Ano06c] Anonymous. Back matter. Biometrika, 93(2):??, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441300.

Anonymous:2006:BMc

[Ano06d] Anonymous. Back matter. Biometrika, 93(3):??, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441321.

Anonymous:2006:BMd

[Ano06e] Anonymous. Back matter. Biometrika, 93(4):??, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441346.

Anonymous:2006:FMa

[Ano06f] Anonymous. Front matter. Biometrika, 93(1):??, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441255.

Anonymous:2006:FMb

[Ano06g] Anonymous. Front matter. Biometrika, 93(2):??, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441277.

Anonymous:2006:FMc

[Ano06h] Anonymous. Front matter. Biometrika, 93(3):??, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441302.

Anonymous:2006:FMd

[Ano06i] Anonymous. Front matter. Biometrika, 93(4):??, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441324.

Anonymous:2006:VIa

[Ano06j] Anonymous. Volume information. Biometrika, 93(1):??, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441254.

Anonymous:2006:VIb

[Ano06k] Anonymous. Volume information. Biometrika, 93(4):??, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441323.

Anonymous:2007:AC

[Ano07a] Anonymous. Amendments and corrections: “Nonparametric Inference in Multivariate Mixtures”. Biometrika, 94(3):767, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441412. See cor-

rection [HNPE05].

Anonymous:2007:BMa

[Ano07b] Anonymous. Back matter. Biometrika, 94(1):??, March 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441367.

Anonymous:2007:BMb

[Ano07c] Anonymous. Back matter. Biometrika, 94(2):??, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441390.

Anonymous:2007:BMc

[Ano07d] Anonymous. Back matter. Biometrika, 94(3):??, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441413.

Anonymous:2007:BMd

[Ano07e] Anonymous. Back matter. Biometrika, 94(4):??, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441435.

Anonymous:2007:FMa

[Ano07f] Anonymous. Front matter. Biometrika, 94(1):??, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441349.

Anonymous:2007:FMb

[Ano07g] Anonymous. Front matter. Biometrika, 94(2):??, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441369.

Anonymous:2007:FMc

[Ano07h] Anonymous. Front matter. Biometrika, 94(3):??, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441392.

Anonymous:2007:FMd

[Ano07i] Anonymous. Front matter. Biometrika, 94(4):??, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441416.

Anonymous:2007:VIa

[Ano07j] Anonymous. Volume information. Biometrika, 94(1):??, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441348.

Anonymous:2007:VIb

[Ano07k] Anonymous. Volume information. Biometrika, 94(4):??, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441415.

Anonymous:2008:BMa

[Ano08a] Anonymous. Back matter. Biometrika, 95(1):??, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441459.

Anonymous:2008:BMb

[Ano08b] Anonymous. Back matter. Biometrika, 95(2):??, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441481.

Anonymous:2008:BMc

[Ano08c] Anonymous. Back matter. Biometrika, 95(3):??, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441502.

Anonymous:2008:BMd

[Ano08d] Anonymous. Back matter. Biometrika, 95(4):??, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441525.

Anonymous:2008:FMa

[Ano08e] Anonymous. Front matter. Biometrika, 95(1):??, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441438.

Anonymous:2008:FMb

[Ano08f] Anonymous. Front matter. Biometrika, 95(2):??, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441462.

Anonymous:2008:FMc

[Ano08g] Anonymous. Front matter. Biometrika, 95(3):??, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441483.

Anonymous:2008:FMd

[Ano08h] Anonymous. Front matter. Biometrika, 95(4):??, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441505.

Anonymous:2008:VIa

[Ano08i] Anonymous. Volume information. Biometrika, 95(1):??, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441437.

Anonymous:2008:VIb

[Ano08j] Anonymous. Volume information. Biometrika, 95(2):??, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441461.

Anonymous:2008:VIc

[Ano08k] Anonymous. Volume information. Biometrika, 95(4):??, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441504.

Atkinson:2002:FSA

[AR02] Anthony C. Atkinson and Marco Riani. Forward search addedvariable t-tests and the effect of masked outliers on model selection. Biometrika, 89(4):939–946, December 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/4/939;
 http://www.jstor.org/stable/4140549.

Antoniadis:2001:WSN

[AS01] Anestis Antoniadis and Theofanis Sapatinas. Wavelet shrinkage for natural exponential families with quadratic variance functions. Biometrika, 88(3):805–820, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673448.

Auh:2006:ILD

[AS06] Sungyoung Auh and Allan R. Sampson. Isotonic logistic discrimination. Biometrika, 93(4):961–972, December 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/4/961;
 http://www.jstor.org/stable/20441338.

Allen:2005:LER

[AST05] Andrew S. Allen, Glen A. Satten, and Anastasios A. Tsiatis. Locally-efficient robust estimation of haplotype-disease association in family-based studies. Biometrika, 92(3):559–571, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/3/559;   http://www.jstor.org/stable/20441214.

Abbring:2007:UHD

[AV07] Jaap H. Abbring and Gerard J. Van Den Berg. The unobserved heterogeneity distribution in duration analysis. Biometrika, 94(1):87– 99, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/1/87;   http://www.jstor.org/stable/20441356.

Bandyopadhyay:2001:ADN

[BB01] Uttam Bandyopadhyay and Atanu Biswas. Adaptive designs for normal responses with prognostic factors. Biometrika, 88(2):409– 419, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2673489.

Bartolucci:2002:RAM

[BB02] Francesco Bartolucci and Julian Besag. A recursive algorithm for Markov random fields. Biometrika, 89(3):724–730, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/3/724;
 http://www.jstor.org/stable/4140615.

Broemeling:2003:SHP

[BB03] Lyle Broemeling and Ana Broemeling. Studies in the history of probability and statistics. XLVIII. The Bayesian contributions of Ernest Lhoste. Biometrika, 90(3):728–731, September 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/728;
 http://www.jstor.org/stable/30042079.

Becker:2004:EVE

[BB04] Niels G. Becker and Tom Britton. Estimating vaccine efficacy from small outbreaks. Biometrika, 91(2):363–382, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/363;
 http://www.jstor.org/stable/20441107.

Ball:2004:SME

[BBL04] Frank G. Ball, Tom Britton, and Owen D. Lyne. Stochastic multitype epidemics in a community of households: Estimation of threshold parameter R_{∗} and secure vaccination coverage. Biometrika, 91(2):345–362, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/345;   http://
www.jstor.org/stable/20441106.

Braun:2000:MCF

[BBM00] J. V. Braun, R. K. Braun, and H.-G. Müller. Multiple changepoint fitting via quasilikelihood, with application to DNA sequence segmentation. Biometrika, 87(2):301–314, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/301;
 http://www.jstor.org/stable/2673465.

Ball:2002:ECS

[BBO02] Frank G. Ball, Tom Britton, and Philip D. O’Neill. Empty confidence sets for epidemics, branching processes and Brownian motion. Biometrika, 89(1):211–224, March 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/1/211;
 http://www.jstor.org/stable/4140568.

Biswas:2002:EDM

[BC02] Atanu Biswas and Probal Chaudhuri. An efficient design for model discrimination and parameter estimation in linear models. Biometrika, 89(3):709–718, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/709;   http://
www.jstor.org/stable/4140613.

Beaumont:2009:AAB

[BCMR09] Mark A. Beaumont, Jean-Marie Cornuet, Jean-Michel Marin, and Christian P. Robert. Adaptive approximate Bayesian computation. Biometrika, 96(4):983–990, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/983;
 http://www.jstor.org/stable/27798882.

Breidt:2005:MAE

[BCO05] F. J. Breidt, G. Claeskens, and J. D. Opsomer. Modelassisted estimation for complex surveys using penalised splines. Biometrika, 92(4):831–846, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/4/831;   http://
www.jstor.org/stable/20441239.

Botts:2006:SES

[BD06] Carsten H. Botts and Michael J. Daniels. A shrinkage estimator for spectral densities. Biometrika, 93(1):179–195, March 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/1/179;
 http://www.jstor.org/stable/20441269.

Bortnick:2006:EDO

[BDS06] Steven M. Bortnick, Angela M. Dean, and Thomas J. Santner. Efficient designs for one-sided comparisons of two or three treatments with a control in a one-way layout. Biometrika, 93(1):127–135, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/1/127;   http://www.jstor.org/stable/20441265.

Basu:2001:BCR

[BE01] Sanjib Basu and Nader Ebrahimi. Bayesian capture-recapture methods for error detection and estimation of population size: Heterogeneity and dependence. Biometrika, 88(1):269–279, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673684.

Beaumont:2008:NAW

[Bea08] Jean-François Beaumont. A new approach to weighting and inference in sample surveys. Biometrika, 95(3):539–553, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/3/539;   http://www.jstor.org/stable/20441485.

Berger:2007:JVE

[Ber07] Yves G. Berger. A jackknife variance estimator for unistage stratified samples with unequal probabilities. Biometrika, 94(4):953–964, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/4/953;   http://www.jstor.org/stable/20441428.

Betensky:2000:NNC

[Bet00] Rebecca A. Betensky. On nonidentifiability and noninformative censoring for current status data. Biometrika, 87(1):218–221, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/1/218;   http://www.jstor.org/stable/2673576.

Bayarri:2007:ECP

[BGD07] M. J. Bayarri and Gonzalo García-Donato. Extending conventional priors for testing general hypotheses in linear models. Biometrika, 94(1):135–152, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/1/135;   http://
www.jstor.org/stable/20441359.

Brown:2001:SMB

[BHY01] B. M. Brown, Peter Hall, and G. A. Young. The smoothed median and the bootstrap. Biometrika, 88(2):519–534, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673497.

Bingham:2000:SHP

[Bin00] N. H. Bingham. Studies in the history of probability and statistics. XLVI. Measure into probability: from Lebesgue to Kolmogorov. Biometrika, 87(1):145–156, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/145;
 http://www.jstor.org/stable/2673568.

Barber:2002:OAO

[BJ02] Stuart Barber and Christopher Jennison. Optimal asymmetric one-sided group sequential tests. Biometrika, 89(1):49–60, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/1/49;   http://www.jstor.org/stable/4140558.

Bailey:2006:OCD

[BK06] R. A. Bailey and J. Kunert. On optimal crossover designs when carryover effects are proportional to direct effects. Biometrika, 93(3):613–625, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/3/613;   http://
www.jstor.org/stable/20441311.

Behseta:2005:HMA

[BKW05] Sam Behseta, Robert E. Kass, and Garrick L. Wallstrom. Hierarchical models for assessing variability among functions. Biometrika, 92(2):419–434, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/419;   http://
www.jstor.org/stable/20441197.

Benjamini:2006:ALS

[BKY06] Yoav Benjamini, Abba M. Krieger, and Daniel Yekutieli. Adaptive linear step-up procedures that control the false discovery rate. Biometrika, 93(3):491–507, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/491;
 http://www.jstor.org/stable/20441303.

Balasubramanian:2003:EFT

[BL03] R. Balasubramanian and S. W. Lagakos. Estimation of a failure time distribution based on imperfect diagnostic tests. Biometrika, 90(1):171–182, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/1/171;   http://
www.jstor.org/stable/30042027.

Blackwell:2003:BIM

[Bla03] P. G. Blackwell. Bayesian inference for Markov processes with diffusion and discrete components. Biometrika, 90(3):613–627, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/613;   http://www.jstor.org/stable/30042070.

Barbieri:2000:BFF

[BLP00] Maria Maddalena Barbieri, Brunero Liseo, and Lea Petrella. Bayes factors for Fieller’s problem. Biometrika, 87(3):717–723, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/3/717;   http://www.jstor.org/stable/2673644.

Besag:2005:FOI

[BM05a] Julian Besag and Debashis Mondal. First-order intrinsic autoregressions and the De Wijs process. Biometrika, 92(4):909–920, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/4/909;   http://www.jstor.org/stable/20441244.

Bischoff:2005:ATS

[BM05b] Wolfgang Bischoff and Frank Miller. Adaptive two-stage test procedures to find the best treatment in clinical trials. Biometrika, 92(1):197–212, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/1/197;   http://
www.jstor.org/stable/20441176.

Boik:2002:SMC

[Boi02] Robert J. Boik. Spectral models for covariance matrices. Biometrika, 89(1):159–182, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/1/159;   http://
www.jstor.org/stable/4140565.

Boik:2003:PCM

[Boi03] Robert J. Boik. Principal component models for correlation matrices. Biometrika, 90(3):679–701, September 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/679;
 http://www.jstor.org/stable/30042075.

Bondell:2005:MDE

[Bon05] Howard D. Bondell. Minimum distance estimation for the logistic regression model. Biometrika, 92(3):724–731, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/724;
 http://www.jstor.org/stable/20441226.

Bondell:2007:TGF

[Bon07] Howard D. Bondell. Testing goodness-of-fit in logistic casecontrol studies. Biometrika, 94(2):487–495, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/487;
 http://www.jstor.org/stable/20441386.

Bottai:2003:CRW

[Bot03] Matteo Bottai. Confidence regions when the Fisher information is zero. Biometrika, 90(1):73–84, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/90/1/73;   http://
www.jstor.org/stable/30042020.

Bartolucci:2007:AQE

[BP07] Francesco Bartolucci and Fulvia Pennoni. On the approximation of the quadratic exponential distribution in a latent variable context. Biometrika, 94(3):745–754, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/745;
 http://www.jstor.org/stable/20441409.

Boente:2002:IFO

[BPR02] Graciela Boente, Ana M. Pires, and Isabel M. Rodrigues. Influence functions and outlier detection under the common principal components model: A robust approach. Biometrika, 89(4):861–875, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/861;   http://www.jstor.org/stable/4140543.

Bravo:2003:SOP

[Bra03] Francesco Bravo. Second-order power comparisons for a class of nonparametric likelihood-based tests. Biometrika, 90(4):881–890, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/4/881;   http://www.jstor.org/stable/30042094.

Bandeen-Roche:2002:MMF

[BRL02] Karen Bandeen-Roche and Kung-Yee Liang. Modelling multivariate failure time associations in the presence of a competing risk. Biometrika, 89(2):299–314, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/299;
 http://www.jstor.org/stable/4140578.

Bandeen-Roche:2008:NEB

[BRN08] Karen Bandeen-Roche and Jing Ning. Nonparametric estimation of bivariate failure time associations in the presence of a competing risk. Biometrika, 95(1):221–232, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/221;
 http://www.jstor.org/stable/20441453.

Benhin:2005:MEE

[BRS05] E. Benhin, J. N. K. Rao, and A. J. Scott. Mean estimating equation approach to analysing cluster-correlated data with nonignorable cluster sizes. Biometrika, 92(2):435–450, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/2/435;
 http://www.jstor.org/stable/20441198.

Betensky:2001:CSA

[BRT01] Rebecca A. Betensky, Daniel Rabinowitz, and Anastasios A. Tsiatis. Computationally simple accelerated failure time regression for interval censored data. Biometrika, 88(3):703–711, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673440.

Bittman:2009:OTM

[BRVW09] Richard M. Bittman, Joseph P. Romano, Carlos Vallarino, and Michael Wolf. Optimal testing of multiple hypotheses with common effect direction. Biometrika, 96(2):399–410, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/399;
 http://www.jstor.org/stable/27798832.

Beyersmann:2008:NNQ

[BS08] Jan Beyersmann and Martin Schumacher. A note on nonparametric quantile inference for competing risks and more complex multistate models. Biometrika, 95(4):1006–1008, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/4/1006;   http://www.jstor.org/stable/20441524.

Beerenwinkel:2009:MMA

[BS09] N. Beerenwinkel and S. Sullivant. Markov models for accumulating mutations. Biometrika, 96(3):645–661, September 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/645;
 http://www.jstor.org/stable/27798854.

Brix:2001:AGF

[BSCC01] A. Brix, R. Senoussi, P. Couteron, and J. Chadoeuf. Assessing goodness of fit of spatially inhomogeneous Poisson processes. Biometrika, 88(2):487–497, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673495.

Bartolucci:2006:EBF

[BSM06] Francesco Bartolucci, Luisa Scaccia, and Antonietta Mira. Efficient Bayes factor estimation from the reversible jump output. Biometrika, 93(1):41–52, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/93/1/41;   http://
www.jstor.org/stable/20441258.

Bingham:2009:ONO

[BST09] Derek Bingham, Randy R. Sitter, and Boxin Tang. Orthogonal and nearly orthogonal designs for computer experiments. Biometrika, 96(1):51–65, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/96/1/51;   http://
www.jstor.org/stable/27798801.

Bang:2000:EMC

[BT00] Heejung Bang and Anastasios A. Tsiatis. Estimating medical costs with censored data. Biometrika, 87(2):329–343, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/329;
 http://www.jstor.org/stable/2673467.

Butler:2001:OOL

[But01] Neil A. Butler. Optimal and orthogonal Latin hypercube designs for computer experiments. Biometrika, 88(3):847–857, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673451.

Butler:2003:MAC

[But03a] Neil A. Butler. Minimum aberration construction results for nonregular two-level fractional factorial designs. Biometrika, 90 (4):891–898, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/4/891;   http://www.jstor.
org/stable/30042095.

Butler:2003:STC

[But03b] Neil A. Butler. Some theory for constructing minimum aberration fractional factorial designs. Biometrika, 90(1):233–238, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/1/233;   http://www.jstor.org/stable/30042032.

Butler:2005:GMA

[But05] Neil A. Butler. Generalised minimum aberration construction results for symmetrical orthogonal arrays. Biometrika, 92(2):485– 491, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/2/485;   http://www.jstor.org/stable/20441202.

Butler:2006:OBT

[But06] Neil A. Butler. Optimal blocking of two-level factorial designs. Biometrika, 93(2):289–302, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/2/289;   http://
www.jstor.org/stable/20441282.

Brown:2005:SEC

[BW05] B. M. Brown and You-Gan Wang. Standard errors and covariance matrices for smoothed rank estimators. Biometrika, 92(1):149–158, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/1/149;   http://www.jstor.org/stable/20441172.

Bailey:2007:ONR

[BW07] R. A. Bailey and E. R. Williams. Optimal nested row-column designs with specified components. Biometrika, 94(2):459–468, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/2/459;   http://www.jstor.org/stable/20441384.

Bjornstad:2008:TSS

[BY08] Jan F. Bjørnstad and Elinor Ytterstad. Two-stage sampling from a prediction point of view when the cluster sizes are unknown.

Biometrika, 95(1):187–204, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/1/187;   http://
www.jstor.org/stable/20441451.

Carota:2005:SDA

[Car05] Cinzia Carota. Symmetric diagnostics for the analysis of the residuals in regression models. Biometrika, 92(4):787–799, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/4/787;   http://www.jstor.org/stable/20441236.

Chandler:2007:ICD

[CB07a] Richard E. Chandler and Steven Bate. Inference for clustered data using the independence loglikelihood. Biometrika, 94(1):167–183, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/1/167;   http://www.jstor.org/stable/20441361.

Cheng:2007:OAR

[CB07b] Yi Cheng and Donald A. Berry. Optimal adaptive randomized designs for clinical trials. Biometrika, 94(3):673–689, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/673;
 http://www.jstor.org/stable/20441404.

Caffo:2002:ESR

[CBD02] Brian S. Caffo, James G. Booth, and A. C. Davison. Empirical supremum rejection sampling. Biometrika, 89(4):745–754, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/745;   http://www.jstor.org/stable/4140534.

Copas:2002:ORO

[CC02] J. B. Copas and P. Corbett. Overestimation of the receiver operating characteristic curve for logistic regression. Biometrika, 89(2):315–331, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/315;   http://
www.jstor.org/stable/4140579.

Cai:2004:SRA

[CC04] T. Cai and S. Cheng. Semiparametric regression analysis for doubly censored data. Biometrika, 91(2):277–290, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/277;
 http://www.jstor.org/stable/20441101.

Chatterjee:2005:SML

[CC05a] Nilanjan Chatterjee and Raymond J. Carroll. Semiparametric maximum likelihood estimation exploiting gene-environment independence in case-control studies. Biometrika, 92(2):399–418, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/2/399;   http://www.jstor.org/stable/20441196.

Chen:2005:SRA

[CC05b] Y. Q. Chen and S. Cheng. Semiparametric regression analysis of mean residual life with censored survival data. Biometrika, 92(1): 19–29, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/1/19;   http://www.jstor.org/stable/20441163.

Cavallini:2006:FIR

[CC06a] David Cavallini and Fabio Corradi. Forensic identification of relatives of individuals included in a database of DNA profiles. Biometrika, 93(3):525–536, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/525;
 http://www.jstor.org/stable/20441305.

Chen:2006:BCE

[CC06b] Song Xi Chen and Hengjian Cui. On Bartlett correction of empirical likelihood in the presence of nuisance parameters. Biometrika, 93(1):215–220, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/1/215;   http://
www.jstor.org/stable/20441272.

Chen:2006:LLE

[CC06c] Y. Q. Chen and S. Cheng. Linear life expectancy regression with censored data. Biometrika, 93(2):303–313, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/303;
 http://www.jstor.org/stable/20441283.

Claeskens:2007:ATM

[CC07] Gerda Claeskens and Raymond J. Carroll. An asymptotic theory for model selection inference in general semiparametric problems. Biometrika, 94(2):249–265, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/249;
 http://www.jstor.org/stable/20441370.

Chen:2008:EBI

[CC08] Jiahua Chen and Zehua Chen. Extended Bayesian information criteria for model selection with large model spaces. Biometrika, 95(3):759–771, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/3/759;   http://
www.jstor.org/stable/20441500.

Craiu:2004:IBE

[CD04] Radu V. Craiu and Thierry Duchesne. Inference based on the EM algorithm for the competing risks model with masked causes of failure. Biometrika, 91(3):543–558, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/20441122.

Chen:2004:BCB

[CDI04] Ming-Hui Chen, Dipak K. Dey, and Joseph G. Ibrahim. Bayesian criterion based model assessment for categorical data. Biometrika, 91(1):45–63, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/91/1/45;   http://
www.jstor.org/stable/20441078.

Crainiceanu:2008:AUE

[CDP08] Ciprian M. Crainiceanu, Francesca Dominici, and Giovanni Parmigiani. Adjustment uncertainty in effect estimation. Biometrika, 95(3):635–651, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/3/635;   http://
www.jstor.org/stable/20441491.

Chaudhuri:2007:ECM

[CDR07] Sanjay Chaudhuri, Mathias Drton, and Thomas S. Richardson. Estimation of a covariance matrix with zeros. Biometrika, 94(1): 199–216, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/1/199;   http://www.jstor.org/stable/20441363.

Colombi:2001:MRM

[CF01] R. Colombi and A. Forcina. Marginal regression models for the analysis of positive association of ordinal response variables. Biometrika, 88(4):1007–1019, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/88/4/1007;   http://
www.jstor.org/stable/2673698.

Cheng:2008:NEC

[CF08a] Yu Cheng and Jason P. Fine. Nonparametric estimation of cause-specific cross hazard ratio with bivariate competing risks data. Biometrika, 95(1):233–240, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/1/233;   http://
www.jstor.org/stable/20441454.

Cook:2008:CRM

[CF08b] R. Dennis Cook and Liliana Forzani. Covariance reducing models: An alternative to spectral modelling of covariance matrices. Biometrika, 95(4):799–812, December 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/799;
 http://www.jstor.org/stable/20441507.

Cai:2005:VSM

[CFLZ05] Jianwen Cai, Jianqing Fan, Runze Li, and Haibo Zhou. Variable selection for multivariate failure time data. Biometrika, 92(2):303– 316, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/2/303;   http://www.jstor.org/stable/20441190.

Chernozhukov:2009:IPI

[CFVG09] V. Chernozhukov, I. Fernández-Val, and A. Galichon. Improving point and interval estimators of monotone functions by rearrangement. Biometrika, 96(3):559–575, September 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/559;
 http://www.jstor.org/stable/27798848.

Conti:2009:GPE

[CGOO09] S. Conti, J. P. Gosling, J. E. Oakley, and A. O’Hagan. Gaussian process emulation of dynamic computer codes. Biometrika, 96(3):663–676, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/96/3/663;   http://www.jstor.
org/stable/27798855.

Choudhuri:2004:CWM

[CGR04] Nidhan Choudhuri, Subhashis Ghosal, and Anindya Roy. Contiguity of the Whittle measure for a Gaussian time series. Biometrika, 91(1):211–218, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/1/211;   http://
www.jstor.org/stable/20441089.

Choi:2000:EPI

[CH00a] Edwin Choi and Peter Hall. On the estimation of poles in intensity functions. Biometrika, 87(2):251–263, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/251;
 http://www.jstor.org/stable/2673462.

Croux:2000:PCA

[CH00b] Christophe Croux and Gentiane Haesbroeck. Principal component analysis based on robust estimators of the covariance or correlation matrix: Influence functions and efficiencies. Biometrika, 87(3):603–618, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/3/603;   http://
www.jstor.org/stable/2673633.

Cantoni:2002:DFT

[CH02a] Eva Cantoni and Trevor Hastie. Degrees-of-freedom tests for smoothing splines. Biometrika, 89(2):251–263, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/251;
 http://www.jstor.org/stable/4140575.

Cox:2002:ESR

[CH02b] D. R. Cox and Peter Hall. Estimation in a simple random effects model with nonnormal distributions. Biometrika, 89(4):831–840, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/831;   http://www.jstor.org/stable/4140540.

Chan:2009:SAC

[CH09] Yao-Ban Chan and Peter Hall. Scale adjustments for classifiers in high-dimensional, low sample size settings. Biometrika, 96(2):469– 478, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/2/469;   http://www.jstor.org/stable/27798837.

Coull:2006:CTM

[CHB06] Brent A. Coull, E. Andres Houseman, and Rebecca A. Betensky. A computationally tractable multivariate random effects model for clustered binary data. Biometrika, 93(3):587–599, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/587;
 http://www.jstor.org/stable/20441309.

Chen:2000:RIM

[Che00] Yi-Hau Chen. A robust imputation method for surrogate outcome data. Biometrika, 87(3):711–716, September 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/3/711;
 http://www.jstor.org/stable/2673643.

Chen:2001:FST

[Che01] Hua Yun Chen. Fitting semiparametric transformation regression models to data from a modified case-cohort design. Biometrika, 88 (1):255–268, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673683.

Cheung:2005:CPD

[Che05] Ying Kuen Cheung. Coherence principles in dose-finding studies. Biometrika, 92(4):863–873, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/4/863;   http://
www.jstor.org/stable/20441241.

Chen:2009:WBT

[Che09] Yi-Hau Chen. Weighted Breslow-type and maximum likelihood estimation in semiparametric transformation models. Biometrika, 96(3):591–600, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/3/591;   http://
www.jstor.org/stable/27798850.

Crump:2009:DLO

[CHIM09] Richard K. Crump, V. Joseph Hotz, Guido W. Imbens, and Oscar A. Mitnik. Dealing with limited overlap in estimation of average treatment effects. Biometrika, 96(1):187–199, March 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/1/187;
 http://www.jstor.org/stable/27798811.

Cui:2003:ADP

[CHN03] Hengjian Cui, Xuming He, and Kai W. Ng. Asymptotic distributions of principal components based on robust dispersions. Biometrika, 90(4):953–966, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/953;   http://
www.jstor.org/stable/30042100.

Chopin:2002:SPF

[Cho02] Nicolas Chopin. A sequential particle filter method for static models. Biometrika, 89(3):539–551, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/539;   http://
www.jstor.org/stable/4140600.

Choi:2000:RPP

[CHP00] Edwin Choi, Peter Hall, and Brett Presnell. Rendering parametric procedures more robust by empirically tilting the model. Biometrika, 87(2):453–465, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/453;   http://
www.jstor.org/stable/2673475.

Chan:2006:NTR

[CHT06a] Kung-Sik Chan, Lop-Hing Ho, and Howell Tong. A note on timereversibility of multivariate linear processes. Biometrika, 93(1):

221–227, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/93/1/221;   http://www.jstor.org/stable/20441273.

Cheng:2006:CBH

[CHT06b] Ming-Yen Cheng, Peter Hall, and Dongsheng Tu. Confidence bands for hazard rates under random censorship. Biometrika, 93(2):357– 366, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/2/357;   http://www.jstor.org/stable/20441287.

Chervoneva:2005:OBA

[CI05] Inna Chervoneva and Boris Iglewicz. Orthogonal bases approach for comparing nonnormal continuous distributions. Biometrika, 92(3):679–690, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/3/679;   http://www.jstor.
org/stable/20441222.

Chen:2006:PPC

[CIS06] Ming-Hui Chen, Joseph G. Ibrahim, and Qi-Man Shao. Posterior propriety and computation for the Cox regression model with applications to missing covariates. Biometrika, 93(4):791–807, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/4/791;   http://www.jstor.org/stable/20441328.

Chen:2001:GCS

[CJ01] Y. Q. Chen and N. P. Jewell. On a general class of semiparametric hazards regression models. Biometrika, 88(3):687–702, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673439.

Chen:2005:LPR

[CJ05] Kani Chen and Zhezhen Jin. Local polynomial regression analysis of clustered data. Biometrika, 92(1):59–74, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/1/59;   http://www.jstor.org/stable/20441166.

Chaganty:2006:RCM

[CJ06] N. Rao Chaganty and Harry Joe. Range of correlation matrices for dependent Bernoulli random variables. Biometrika, 93(1):197–206, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/1/197;   http://www.jstor.org/stable/20441270.

Chen:2002:SAT

[CJY02] Kani Chen, Zhezhen Jin, and Zhiliang Ying. Semiparametric analysis of transformation models with censored data. Biometrika, 89(3):659–668, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/659;   http://
www.jstor.org/stable/4140609.

Chang:2008:UCW

[CK08] Ted Chang and Phillip S. Kott. Using calibration weighting to adjust for nonresponse under a plausible model. Biometrika, 95(3):555–571, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/95/3/555;   http://www.jstor.
org/stable/20441486.

Claeskens:2009:APP

[CKO09] Gerda Claeskens, Tatyana Krivobokova, and Jean D. Opsomer. Asymptotic properties of penalized spline estimators. Biometrika, 96(3):529–544, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/3/529;   http://
www.jstor.org/stable/27798846.

Chan:2003:BET

[CKS03] K. S. Chan, A. B. Kristoffersen, and N. Chr. Stenseth. Bürmann expansion and test for additivity. Biometrika, 90(1):209–222, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/1/209;   http://www.jstor.org/stable/30042030.

Chen:2000:NAA

[CL00] Song Xi Chen and Chris J. Lloyd. A nonparametric approach to the analysis of two-stage mark-recapture experiments. Biometrika, 87(3):633–649, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/3/633;   http://
www.jstor.org/stable/2673635.

Chen:2006:BMT

[CL06a] Shu-Chuan Chen and Bruce G. Lindsay. Building mixture trees from binary sequence data. Biometrika, 93(4):843–860, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/843;   http://www.jstor.org/stable/20441331.

Consonni:2006:RPD

[CL06b] Guido Consonni and Valentina Leucari. Reference priors for discrete graphical models. Biometrika, 93(1):23–40, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/1/23;   http://www.jstor.org/stable/20441257.

Cox:2008:PTF

[CL08] Dennis D. Cox and Jong Soo Lee. Pointwise testing with functional data using the Westfall-Young randomization method. Biometrika, 95(3):621–634, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/3/621;   http://
www.jstor.org/stable/20441490.

Cui:2009:NLT

[CL09] Yunwei Cui and Robert Lund. A new look at time series of counts. Biometrika, 96(4):781–792, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/781;
 http://www.jstor.org/stable/27798866.

Clayton:2003:CLI

[Cla03] David Clayton. Conditional likelihood inference under complex ascertainment using data augmentation. Biometrika, 90(4):976–981, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/4/976;   http://www.jstor.org/stable/30042102.

Cook:2007:DRR

[CLC07] R. Dennis Cook, Bing Li, and Francesca Chiaromonte. Dimension reduction in regression without matrix inversion. Biometrika, 94(3):569–584, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/94/3/569;   http://www.jstor.
org/stable/20441397.

Catchpole:2001:DPR

[CM01] E. A. Catchpole and B. J. T. Morgan. Deficiency of parameterredundant models. Biometrika, 88(2):593–598, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673504.

Critchley:2004:DII

[CM04] Frank Critchley and Paul Marriott. Data-informed influence analysis. Biometrika, 91(1):125–140, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/1/125;   http://
www.jstor.org/stable/20441083.

Chang:2008:BFC

[CM08] In Hong Chang and Rahul Mukerjee. Bayesian and frequentist confidence intervals arising from empirical-type likelihoods. Biometrika, 95(1):139–147, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/1/139;   http://
www.jstor.org/stable/20441448.

Carroll:2009:NAR

[CMMY09] Raymond J. Carroll, Arnab Maity, Enno Mammen, and Kyusang Yu. Nonparametric additive regression for repeatedly measured data. Biometrika, 96(2):383–398, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/383;
 http://www.jstor.org/stable/27798831.

Carvalho:2007:SHI

[CMW07] Carlos M. Carvalho, Hélène Massam, and Mike West. Simulation of hyper-inverse Wishart distributions in graphical models. Biometrika, 94(3):647–659, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/3/647;   http://
www.jstor.org/stable/20441402.

Cook:2006:UIC

[CN06] R. Dennis Cook and Liqiang Ni. Using intraslice covariances for improved estimation of the central subspace in regression. Biometrika, 93(1):65–74, March 2006. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/93/1/65;   http://
www.jstor.org/stable/20441260.

Cribari-Neto:2000:IHC

[CNFC00] Francisco Cribari-Neto, Silvia L. P. Ferrari, and Gauss M. Cordeiro. Improved heteroscedasticity-consistent covariance matrix estimators. Biometrika, 87(4):907–918, December 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/907;
 http://www.jstor.org/stable/2673618.

Cox:2001:BFY

[Cox01] D. R. Cox. Biometrika: The first 100 years. Biometrika, 88(1):3– 11, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2673673.

Cox:2003:CMA

[Cox03] D. R. Cox. Conditional and marginal association for binary random variables. Biometrika, 90(4):982–984, December 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/982;
 http://www.jstor.org/stable/30042103.

Cox:2007:GRW

[Cox07] D. R. Cox. On a generalization of a result of W. G. Cochran. Biometrika, 94(3):755–759, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/3/755;   http://
www.jstor.org/stable/20441410.

Cox:2008:IMD

[Cox08] D. R. Cox. On an internal method for deriving a summary measure. Biometrika, 95(4):1002–1005, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/4/1002;   http://www.jstor.org/stable/20441523.

Carota:2002:SRC

[CP02a] Cinzia Carota and Giovanni Parmigiani. Semiparametric regression for count data. Biometrika, 89(2):265–281, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/265;
 http://www.jstor.org/stable/4140576.

Coles:2002:MIU

[CP02b] Stuart Coles and Francesco Pauli. Models and inference for uncertainty in extremal dependence. Biometrika, 89(1):183–196, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/1/183;   http://www.jstor.org/stable/4140566.

Chan:2005:WLA

[CP05] Ngai Hang Chan and Liang Peng. Weighted least absolute deviations estimation for an AR(1) process with ARCH(1) errors. Biometrika, 92(2):477–484, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/477;   http://
www.jstor.org/stable/20441201.

Chen:2009:EDD

[CPQ09] Song Xi Chen, Liang Peng, and Ying-Li Qin. Effects of data dimension on empirical likelihood. Biometrika, 96(3):711–722, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/3/711;   http://www.jstor.org/stable/27798858.

Chen:2000:ELC

[CQ00] Song Xi Chen and Yong Song Qin. Empirical likelihood confidence intervals for local linear smoothers. Biometrika, 87(4):946–953, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/946;   http://www.jstor.org/stable/2673622.

Cox:2004:NPC

[CR04] D. R. Cox and N. Reid. A note on pseudolikelihood constructed from marginal densities. Biometrika, 91(3):729–737, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441134.

Chaieb:2006:ESU

[CRA06] Lajmi Lakhal Chaieb, Louis-Paul Rivest, and Belkacem Abdous. Estimating survival under a dependent truncation. Biometrika, 93(3):655–669, September 2006. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/3/655;   http://
www.jstor.org/stable/20441314.

Crainiceanu:2005:ELR

[CRCW05] Ciprian Crainiceanu, David Ruppert, Gerda Claeskens, and M. P. Wand. Exact likelihood ratio tests for penalised splines. Biometrika, 92(1):91–103, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/92/1/91;   http://
www.jstor.org/stable/20441168.

Chen:2002:AHM

[CRW02] Y. Q. Chen, C. A. Rohde, and M.-C.-C. Wang. Additive hazards models with latent treatment effectiveness lag time. Biometrika, 89(4):917–931, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/917;   http://
www.jstor.org/stable/4140547.

Cheng:2005:BAD

[CS05] Yi Cheng and Yu Shen. Bayesian adaptive designs for clinical trials. Biometrika, 92(3):633–646, September 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/633;
 http://www.jstor.org/stable/20441219.

Carvalho:2009:OBM

[CS09] C. M. Carvalho and J. G. Scott. Objective Bayesian model selection in Gaussian graphical models. Biometrika, 96(3):497–512, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/3/497;   http://www.jstor.org/stable/27798844.

Cheng:2003:CSS

[CSB03]  Yi Cheng, Fusheng Su, and Donald A. Berry.  Choos-

ing sample size for a clinical trial using decision analysis. Biometrika, 90(4):923–936, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/923;   http://
www.jstor.org/stable/30042098.

Cheng:2009:ENE

[CSTT09] Jing Cheng, Dylan S. Small, Zhiqiang Tan, and Thomas R. Ten Have. Efficient nonparametric estimation of causal effects in randomized trials with noncompliance. Biometrika, 96(1):19–36, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/1/19;   http://www.jstor.org/stable/27798799.

Chen:2002:UEL

[CSW02] J. Chen, R. R. Sitter, and C. Wu. Using empirical likelihood methods to obtain range restricted weights in regression estimators for surveys. Biometrika, 89(1):230–237, March 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/1/230;
 http://www.jstor.org/stable/4140570.

Cohen:2008:FBM

[CSXB08] Arthur Cohen, H. B. Sackrowitz, Minya Xu, and Steven Buyske. A family of Bayes multiple testing procedures. Biometrika, 95(2):295–305, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/295;   http://
www.jstor.org/stable/20441465.

Cheng:2001:UBN

[CT01] Ching-Shui Cheng and Boxin Tang. Upper bounds on the number of columns in supersaturated designs. Biometrika, 88(4):1169– 1174, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/88/4/1169;   http://www.jstor.org/stable/2673708.

Chan:2004:TMD

[CT04] K. S. Chan and H. Tong. Testing for multimodality with dependent data. Biometrika, 91(1):113–123, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/113;
 http://www.jstor.org/stable/20441082.

Carolan:2005:NTA

[CT05] Christopher A. Carolan and Joshua M. Tebbs. Nonparametric tests for and against likelihood ratio ordering in the twosample problem. Biometrika, 92(1):159–171, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/159;
 http://www.jstor.org/stable/20441173.

Chambers:2006:QMS

[CT06] Ray Chambers and Nikos Tzavidis. M-quantile models for small area estimation. Biometrika, 93(2):255–268, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/255;
 http://www.jstor.org/stable/20441279.

Cheng:2009:OTL

[CT09] Ching-Shui Cheng and Pi-Wen Tsai. Optimal two-level regular fractional factorial block and split-plot designs. Biometrika, 96(1): 83–93, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/96/1/83;   http://www.jstor.org/stable/27798803.

Cao:2009:IER

[CTD09] Weihua Cao, Anastasios A. Tsiatis, and Marie Davidian. Improving efficiency and robustness of the doubly robust estimator for a population mean with incomplete data. Biometrika, 96(3):723–734, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/3/723;   http://www.jstor.org/stable/27798859.

Cai:2008:PFR

[CTSW08] Tianxi Cai, Lu Tian, Scott D. Solomon, and L. J. Wei. Predicting future responses based on possibly mis-specified working models. Biometrika, 95(1):75–92, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/95/1/75;   http://
www.jstor.org/stable/20441444.

Cai:2005:SBC

[CTW05] Tianxi Cai, Lu Tian, and L. J. Wei. Semiparametric Box: Cox power transformation models for censored survival observations. Biometrika, 92(3):619–632, September 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/619;
 http://www.jstor.org/stable/20441218.

Cuzick:2001:EBA

[Cuz01] Jack Cuzick. Event-based analysis times for randomised clinical trials. Biometrika, 88(1):245–253, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673682.

Cheng:2000:ISM

[CW00a] S. C. Cheng and L. J. Wei. Inferences for a semiparametric model with panel data. Biometrika, 87(1):89–97, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/1/89;   http://www.jstor.org/stable/2673563.

Cox:2000:GCF

[CW00b] D. R. Cox and Nanny Wermuth. On the generation of the chordless four-cycle. Biometrika, 87(1):206–212, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/206;
 http://www.jstor.org/stable/2673574.

Cox:2002:SMM

[CW02] D. R. Cox and Nanny Wermuth. On some models for multivariate binary variables parallel in complexity with the multivariate Gaussian distribution. Biometrika, 89(2):462–469, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/462;
 http://www.jstor.org/stable/4140591.

Carroll:2008:NVE

[CW08] Raymond J. Carroll and Yuedong Wang. Nonparametric variance estimation in the analysis of microarray data: A measurement error approach. Biometrika, 95(2):437–449, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/437;
 http://www.jstor.org/stable/20441474.

Cai:2000:SRA

[CWW00] T. Cai, L. J. Wei, and M. Wilcox. Semiparametric regression analysis for clustered failure time data. Biometrika, 87(4):867–878, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/867;   http://www.jstor.org/stable/2673615.

Chen:2007:ETE

[CYCJ07] Y. Q. Chen, J. Yang, S. Cheng, and J. B. Jackson. Estimating a treatment effect with repeated measurements accounting for varying effectiveness duration. Biometrika, 94(2):387–402, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/387;
 http://www.jstor.org/stable/20441379.

Chen:2008:ALA

[CYZZ08] Kani Chen, Zhiliang Ying, Hong Zhang, and Lincheng Zhao. Analysis of least absolute deviation. Biometrika, 95(1):107–122, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/1/107;   http://www.jstor.org/stable/20441446.

Datta:2005:ADD

[Dat05] Gauri Sankar Datta. An alternative derivation of the distributions of the maximum likelihood estimators of the parameters in an inverse Gaussian distribution. Biometrika, 92(4):975–977, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/4/975;   http://www.jstor.org/stable/20441251.

Davison:2001:BCT

[Dav01] A. C. Davison. Biometrika Centenary: Theory and general methodology. Biometrika, 88(1):13–52, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673674.

Davies:2002:HTW

[Dav02] Robert B. Davies. Hypothesis testing when a nuisance parameter is present only under the alternative: Linear model case. Biometrika, 89(2):484–489, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/484;   http://
www.jstor.org/stable/4140594.

Du:2003:NAC

[DAV03] Yunling Du, Michael G. Akritas, and Ingrid Van Keilegom. Nonparametric analysis of covariance for censored data. Biometrika, 90(2):269–287, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/2/269;   http://www.jstor.
org/stable/30042039.

DeStavola:2008:CO

[DC08] B. L. De Stavola and D. R. Cox. On the consequences of overstratification. Biometrika, 95(4):992–996, December 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/992;
 http://www.jstor.org/stable/20441521.

Dominici:2005:SQR

[DCNZ05] Francesca Dominici, Leslie Cope, Daniel Q. Naiman, and Scott L. Zeger. Smooth quantile ratio estimation. Biometrika, 92(3):543– 557, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/3/543;   http://www.jstor.org/stable/20441213.

Davis:2003:ODM

[DDS03] Richard A. Davis, William T. M. Dunsmuir, and Sarah B. Streett. Observation-driven models for Poisson counts. Biometrika, 90 (4):777–790, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/4/777;   http://www.jstor.
org/stable/30042088.

Davis:2000:APR

[DDW00] Richard A. Davis, William T. M. Dunsmuir, and Ying Wang. On autocorrelation in a Poisson regression model. Biometrika, 87(3):491–505, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/3/491;   http://www.jstor.
org/stable/2673626.

Delicado:2006:LLD

[Del06] Pedro Delicado. Local likelihood density estimation based on smooth truncation. Biometrika, 93(2):472–480, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/472;
 http://www.jstor.org/stable/20441297.

Demidenko:2000:LSE

[Dem00] Eugene Demidenko. Is this the least squares estimate? Biometrika, 87(2):437–452, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/2/437;   http://www.jstor.
org/stable/2673474.

Dellaportas:2006:BMS

[DFR06] Petros Dellaportas, Nial Friel, and Gareth O. Roberts. Bayesian model selection for partially observed diffusion models. Biometrika, 93(4):809–825, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/4/809;   http://www.jstor.
org/stable/20441329.

Dai:2004:MSA

[DG04] Ming Dai and Wensheng Guo. Multivariate spectral analysis using Cholesky decomposition. Biometrika, 91(3):629–643, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/20441128.

Duan:2007:GSD

[DGG07] Jason A. Duan, Michele Guindani, and Alan E. Gelfand. Generalized spatial Dirichlet process models. Biometrika, 94(4):809–825, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/4/809;   http://www.jstor.org/stable/20441419.

DiMatteo:2001:BCF

[DGK01] Ilaria DiMatteo, Christopher R. Genovese, and Robert E. Kass. Bayesian curve-fitting with free-knot splines. Biometrika, 88(4): 1055–1071, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/88/4/1055;   http://www.jstor.
org/stable/2673701.

Davis:2002:EIM

[DHO02] B. A. Davis, C. R. Heathcote, and T. J. O’Neill. Estimating and interpolating a Markov chain from aggregate data. Biometrika, 89(1):95–110, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/89/1/95;   http://
www.jstor.org/stable/4140561.

Denne:2000:GST

[DJ00] Jonathan S. Denne and Christopher Jennison. A group sequential t-test with updating of sample size. Biometrika, 87(1):125–134, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/1/125;   http://www.jstor.org/stable/2673566.

Durbin:2002:SES

[DK02] J. Durbin and S. J. Koopman. A simple and efficient simulation smoother for state space time series analysis. Biometrika, 89(3): 603–615, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/3/603;   http://www.jstor.org/stable/4140605.

Dauxois:2003:TPO

[DK03]  Jean-Yves Dauxois and Syed N. U. A. Kirmani.  Test-

ing the proportional odds model under random censoring. Biometrika, 90(4):913–922, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/913;   http://
www.jstor.org/stable/30042097.

Dette:2004:CSN

[DK04] Holger Dette and Robert Kwiecien. A comparison of sequential and non-sequential designs for discrimination between nested regression models. Biometrika, 91(1):165–176, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/165;
 http://www.jstor.org/stable/20441086.

Dryden:2008:MDS

[DKLW08] Ian L. Dryden, Alfred Kume, Huiling Le, and Andrew T. A. Wood. A multi-dimensional scaling approach to shape analysis. Biometrika, 95(4):779–798, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/779;   http://
www.jstor.org/stable/20441506.

DiRienzo:2001:BCS

[DL01] A. G. DiRienzo and S. W. Lagakos. Bias correction for score tests arising from misspecified proportional hazards regression models. Biometrika, 88(2):421–434, June 2001. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673490.

DoHa:2005:CHL

[DL05] Il Do Ha and Youngjo Lee. Comparison of hierarchical likelihood versus orthodox best linear unbiased predictor approaches for frailty models. Biometrika, 92(3):717–723, September 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/717;
 http://www.jstor.org/stable/20441225.

DoHa:2001:HLA

[DLS01] Il Do Ha, Youngjo Lee, and Jae-Kee Song. Hierarchical likelihood approach for frailty models. Biometrika, 88(1):233–243, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673681.

Delgado:2000:NTS

[DM00] Miguel A. Delgado and Juan Mora. A nonparametric test for serial independence of regression errors. Biometrika, 87(1):228–234, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/1/228;   http://www.jstor.org/stable/2673578.

DiCiccio:2001:APE

[DM01] Thomas J. DiCiccio and Anna Clara Monti. Approximations to the profile empirical likelihood function for a scalar parameter in the context of M-estimation. Biometrika, 88(2):337–351, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673483.

DiCiccio:2002:ACL

[DM02a] Thomas J. DiCiccio and Anna Clara Monti. Accurate confidence limits for scalar functions of vector M-estimands. Biometrika, 89(2):437–450, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/437;   http://
www.jstor.org/stable/4140588.

Downs:2002:CR

[DM02b] T. D. Downs and K. V. Mardia. Circular regression. Biometrika, 89(3):683–697, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/89/3/683;   http://www.jstor.
org/stable/4140611.

Davison:2009:SAM

[DM09] A. C. Davison and D. Mastropietro. Saddlepoint approximation for mixture models. Biometrika, 96(2):479–486, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/479;
 http://www.jstor.org/stable/27798838.

Downs:2003:SR

[Dow03] T. D. Downs. Spherical regression. Biometrika, 90(3):655–668, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/655;   http://www.jstor.org/stable/30042073.

Daniels:2002:BAC

[DP02] Michael J. Daniels and Mohsen Pourahmadi. Bayesian analysis of covariance matrices and dynamic models for longitudinal data. Biometrika, 89(3):553–566, August 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/3/553;
 http://www.jstor.org/stable/4140601.

Drton:2004:MSG

[DP04] Mathias Drton and Michael D. Perlman. Model selection for Gaussian concentration graphs. Biometrika, 91(3):591–602, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441125.

Dunson:2008:KSB

[DP08a] David B. Dunson and Ju-Hyun Park. Kernel stick-breaking processes. Biometrika, 95(2):307–323, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/307;   http://
www.jstor.org/stable/20441466.

Dunson:2008:BNI

[DP08b] David B. Dunson and Shyamal D. Peddada. Bayesian nonparametric inference on stochastic ordering. Biometrika, 95(4):859–874, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/4/859;   http://www.jstor.org/stable/20441511.

Drton:2004:MLB

[DR04] Mathias Drton and Thomas S. Richardson. Multimodality of the likelihood in the bivariate seemingly unrelated regressions model. Biometrika, 91(2):383–392, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/383;
 http://www.jstor.org/stable/20441108.

deRossi:2001:HOA

[dRG01] François-Xavier de Rossi and Riccardo Gatto. High-order asymptotic expansions for robust tests. Biometrika, 88(4):1153–1168, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/1153;   http://www.jstor.org/stable/2673707.

Datta:2005:MVS

[DRS05] Gauri Sankar Datta, J. N. K. Rao, and David Daniel Smith. On measuring the variability of small area estimators under a basic area level model. Biometrika, 92(1):183–196, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/1/183;   http://www.jstor.org/stable/20441175. See

correction [DRS12].

Datta:2012:ACM

[DRS12] Gauri Sankar Datta, J. N. K. Rao, and David Daniel Smith. Amendments and corrections: “On Measuring the Variability of Small Area Estimators under a Basic Area Level Model” [MR2158619]. Biometrika, 99(2):509, June 2012. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/41720709. See [DRS05].

Deville:2004:EBS

[DT04a] Jean-Claude Deville and Yves Tillé. Efficient balanced sampling: The cube method. Biometrika, 91(4):893–912, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/4/893;
 http://www.jstor.org/stable/20441151.

DiMarzio:2004:BKD

[DT04b] Marco Di Marzio and Charles C. Taylor. Boosting kernel density estimates: A bias reduction technique? Biometrika, 91(1):226–233, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/1/226;   http://www.jstor.org/stable/20441091.

Druilhet:2009:ORM

[DT09] P. Druilhet and W. Tinsson. Optimal repeated measurement designs for a model with partial interactions. Biometrika, 96(3):677– 690, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/96/3/677;   http://www.jstor.org/stable/27798856.

Dunson:2009:NBL

[Dun09] David B. Dunson. Nonparametric Bayes local partition models for random effects. Biometrika, 96(2):249–262, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/249;
 http://www.jstor.org/stable/27798822.

Davison:2002:SAS

[DW02] A. C. Davison and Suojin Wang. Saddlepoint approximations as smoothers. Biometrika, 89(4):933–938, December 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/4/933;
 http://www.jstor.org/stable/4140548.

Davis:2009:NBM

[DW09] Richard A. Davis and Rongning Wu. A negative binomial model for time series of counts. Biometrika, 96(3):735–749, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/735;
 http://www.jstor.org/stable/27798860.

Diciccio:2008:CPU

[DY08] Thomas J. Diciccio and G. Alastair Young. Conditional properties of unconditional parametric bootstrap procedures for inference in exponential families. Biometrika, 95(3):747–758, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/747;
 http://www.jstor.org/stable/20441499.

Ebrahimi:2006:MRE

[Ebr06] Nader Ebrahimi. Models for recurring events with marginal proportional hazards. Biometrika, 93(2):481–485, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/481;
 http://www.jstor.org/stable/20441298.

Eguchi:2002:CLT

[EC02] Shinto Eguchi and John Copas. A class of logistic-type discriminant functions. Biometrika, 89(1):1–22, March 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/1/1;
http://www.jstor.org/stable/4140555.

Edmondson:2002:GIT

[Edm02] R. N. Edmondson. Generalised incomplete Trojan designs. Biometrika, 89(4):877–891, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/877;   http://
www.jstor.org/stable/4140544.

Edwards:2001:TAM

[EL01] David Edwards and Steffen L. Lauritzen. The TM algorithm for maximising a conditional likelihood function. Biometrika, 88 (4):961–972, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/88/4/961;   http://www.jstor.
org/stable/2673695.

Epifani:2003:EFM

[ELP03] Ilenia Epifani, Antonio Lijoi, and Igor Prünster. Exponential functionals and means of neutral-to-the-right priors. Biometrika, 90(4):791–808, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/791;   http://
www.jstor.org/stable/30042089.

Ebrahimi:2003:ICD

[EMY03] Nader Ebrahimi, Daniel Molefe, and Zhiliang Ying. Identifiability and censored data. Biometrika, 90(3):724–727, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/724;
 http://www.jstor.org/stable/30042078.

Evans:2001:CDE

[ES01] Richard Evans and J. Sedransk. Combining data from experiments that May be similar. Biometrika, 88(3):643–656, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673436.

Fattorini:2006:AHT

[Fat06] Lorenzo Fattorini. Applying the Horvitz-Thompson criterion in complex designs: a computer-intensive perspective for estimating inclusion probabilities. Biometrika, 93(2):269–278, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/269;
 http://www.jstor.org/stable/20441280.

Fried:2003:DSG

[FD03] Roland Fried and Vanessa Didelez. Decomposability and selection of graphical models for multivariate time series. Biometrika, 90(2):251–267, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/251;   http://
www.jstor.org/stable/30042038.

Firth:2004:QV

[FD04] David Firth and Renée X. De Menezes. Quasi-variances. Biometrika, 91(1):65–80, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/91/1/65;   http://
www.jstor.org/stable/20441079.

Falconer:2007:IFP

[FF07] Kenneth Falconer and Carmen Fernández. Inference on fractal processes using multiresolution approximation. Biometrika, 94(2):313– 334, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/2/313;   http://www.jstor.org/stable/20441374.

Fonseca:2008:OBA

[FFM08] Thaís C. O. Fonseca, Marco A. R. Ferreira, and Helio S. Migon. Objective Bayesian analysis for the Student-t regression model. Biometrika, 95(2):325–333, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/325;
 http://www.jstor.org/stable/20441467.

Fuh:2004:EIS

[FH04] Cheng-Der Fuh and Inchi Hu. Efficient importance sampling for events of moderate deviations with applications. Biometrika, 91(2):471–490, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/471;   http://
www.jstor.org/stable/20441114.

Fine:2002:CNC

[Fin02] J. P. Fine. Comparing nonnested Cox models. Biometrika, 89(3): 635–647, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/3/635;   http://www.jstor.org/stable/4140607.

Fine:2000:ACT

[FJ00] J. P. Fine and H. Jiang. On association in a copula with time transformations. Biometrika, 87(3):559–571, September 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/3/559;
 http://www.jstor.org/stable/2673630.

Fewster:2009:IPS

[FJ09] R. M. Fewster and P. E. Jupp. Inference on population size in binomial detectability models. Biometrika, 96(4):805–820, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/4/805;   http://www.jstor.org/stable/27798868.

Fine:2001:SCR

[FJC01] J. P. Fine, H. Jiang, and R. Chappell. On semi-competing risks data. Biometrika, 88(4):907–919, December 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/88/4/907;
 http://www.jstor.org/stable/2673691.

Fienberg:2007:PAA

[FK07] Stephen E. Fienberg and Sung-Ho Kim. Positive association among three binary variables and cross-product ratios. Biometrika, 94(4):999–1005, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/4/999;   http://
www.jstor.org/stable/20441433.

Fu:2002:TMA

[FLF02] Bo Fu, Wai-Keung Li, and Wing-Kam Fung. Testing model adequacy for dynamic panel data with intercorrelation. Biometrika, 89(3):591–601, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/89/3/591;   http://www.jstor.
org/stable/4140604.

Fang:2000:CBU

[FM00] Kai-Tai Fang and Rahul Mukerjee. A connection between uniformity and aberration in regular fractions of two-level factorials. Biometrika, 87(1):193–198, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/1/193;   http://
www.jstor.org/stable/2673572.

Fisher:2001:MTE

[FM01] N. I. Fisher and J. S. Marron. Mode testing via the excess mass estimate. Biometrika, 88(2):499–517, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673496.

Fang:2005:ELC

[FM05] Kai-Tai Fang and Rahul Mukerjee. Expected lengths of confidence intervals based on empirical discrepancy statistics. Biometrika, 92(2):499–503, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/499;   http://
www.jstor.org/stable/20441204.

Fang:2006:ETL

[FM06] Kai-Tai Fang and Rahul Mukerjee. Empirical-type likelihoods allowing posterior credible sets with frequentist validity: Higherorder asymptotics. Biometrika, 93(3):723–733, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/723;
 http://www.jstor.org/stable/20441318.

Felix-Medina:2004:ACD

[FMT04] Martín H. Félix-Medina and Steven K. Thompson. Adaptive cluster double sampling. Biometrika, 91(4):877–891, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/4/877;
 http://www.jstor.org/stable/20441150.

Friel:2007:RCS

[FR07] Nial Friel and Håvard Rue. Recursive computing and simulationfree inference for general factorizable models. Biometrika, 94(3): 661–672, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/3/661;   http://www.jstor.org/stable/20441403.

Fraser:2008:SDA

[FR08] D. A. S. Fraser and Judith Rousseau. Studentization and deriving accurate P -values. Biometrika, 95(1):1–16, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/1;
http://www.jstor.org/stable/20441439.

Fraser:2003:LCP

[Fra03]  D. A. S. Fraser.  Likelihood for component parameters.

Biometrika, 90(2):327–339, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/327;   http://
www.jstor.org/stable/30042043.

Fryzlewicz:2006:HFT

[FSR06] Piotr Fryzlewicz, Theofanis Sapatinas, and Suhasini Subba Rao. A Haar-Fisz technique for locally stationary volatility estimation. Biometrika, 93(3):687–704, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/3/687;   http://
www.jstor.org/stable/20441316.

Fruhwirth-Schnatter:2006:AMS

[FSW06] Sylvia Frühwirth-Schnatter and Helga Wagner. Auxiliary mixture sampling for parameter-driven models of time series of counts with applications to state space modelling. Biometrika, 93(4):827–841, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/4/827;   http://
www.jstor.org/stable/20441330.

Fuentes:2002:SMN

[Fue02] Montserrat Fuentes. Spectral methods for nonstationary spatial processes. Biometrika, 89(1):197–210, March 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/1/197;
 http://www.jstor.org/stable/4140567.

Fuller:2009:SDP

[Ful09] Wayne A. Fuller. Some design properties of a rejective sampling procedure. Biometrika, 96(4):933–944, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/933;
 http://www.jstor.org/stable/27798877.

Feng:2008:SWL

[FW08] Wentao Feng and Abdus S. Wahed. Supremum weighted logrank test and sample size for comparing two-stage adaptive treatment strategies. Biometrika, 95(3):695–707, September 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/695;
 http://www.jstor.org/stable/20441495.

Fan:2004:CME

[FY04] Jianqing Fan and Tsz Ho Yim. A crossvalidation method for estimating conditional densities. Biometrika, 91(4):819–834, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/4/819;   http://www.jstor.org/stable/20441146.

Fine:2004:TPR

[FYK04] J. P. Fine, J. Yan, and M. R. Kosorok. Temporal process regression. Biometrika, 91(3):683–703, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/20441131.

Fan:2004:GLR

[FZ04] Jianqing Fan and Wenyang Zhang. Generalised likelihood ratio tests for spectral density. Biometrika, 91(1):195–209, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/195;
 http://www.jstor.org/stable/20441088.

Gabriel:2002:GFB

[Gab02] K. Ruben Gabriel. Goodness of fit of biplots and correspondence analysis. Biometrika, 89(2):423–436, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/423;
 http://www.jstor.org/stable/4140587.

Grabarnik:2002:GFT

[GC02] P. Grabarnik and S. N. Chiu. Goodness-of-fit test for complete spatial randomness against mixtures of regular and clustered spatial point processes. Biometrika, 89(2):411–421, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/411;
 http://www.jstor.org/stable/4140586.

Garthwaite:2004:DDB

[GC04] Paul H. Garthwaite and John R. Crawford. The distribution of the difference between two t-variates. Biometrika, 91(4):987–994, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/4/987;   http://www.jstor.org/stable/20441157.

Goga:2009:UFL

[GDRG09] C. Goga, J.-C. Deville, and A. Ruiz-Gazen. Use of functionals in linearization and composite estimation with application to twosample survey data. Biometrika, 96(3):691–709, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/691;
 http://www.jstor.org/stable/27798857.

Gervini:2008:RFE

[Ger08] Daniel Gervini. Robust functional estimation using the median and spherical principal components. Biometrika, 95(3):587–600, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/3/587;   http://www.jstor.org/stable/20441488.

George:2000:CEB

[GF00a] Edward I. George and Dean P. Foster. Calibration and empirical Bayes variable selection. Biometrika, 87(4):731–747, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/4/731;   http://www.jstor.org/stable/2673607.

Goldenshluger:2000:NEW

[GF00b] Alexander Goldenshluger and David Faraggi. Nonparametric estimation of weighted proportions with application to particle size measurement. Biometrika, 87(1):157–171, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/157;
 http://www.jstor.org/stable/2673569.

Gervini:2005:NML

[GG05] Daniel Gervini and Theo Gasser. Nonparametric maximum likelihood estimation of the structural mean of a sample of curves. Biometrika, 92(4):801–820, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/92/4/801;   http:/
/www.jstor.org/stable/20441237. See corrections [GG06a,

Ano06a].

Gervini:2006:ACN

[GG06a] D. Gervini and T. Gasser. Amendments and corrections: “Nonparametric maximum likelihood estimation of the structural mean of a sample of curves” [Biometrika 92 (2005), no. 4, 801– 820; MR2234187]. Biometrika, 93(4):1025, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/1025.1. See [GG05].

Guttorp:2006:SHP

[GG06b] Peter Guttorp and Tilmann Gneiting. Studies in the history of probability and statistics. XLIX. On the Matérn correlation family. Biometrika, 93(4):989–995, December 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/4/989;
 http://www.jstor.org/stable/20441340.

Gonen:2005:CPD

[GH05] Mithat Gönen and Glenn Heller. Concordance probability and discriminatory power in proportional hazards regression. Biometrika, 92(4):965–970, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/4/965;   http://
www.jstor.org/stable/20441249.

Gijbels:2000:TMR

[GHJK00] Irene Gijbels, Peter Hall, M. C. Jones, and Inge Koch. Tests for monotonicity of a regression mean with guaranteed level. Biometrika, 87(3):663–673, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/3/663;   http://
www.jstor.org/stable/2673637.

Gibbs:2000:BCT

[Gib00] Alison L. Gibbs. Bounding the convergence time of the Gibbs sampler in Bayesian image restoration. Biometrika, 87(4):749–766, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/749;   http://www.jstor.org/stable/2673608.

Goudie:2007:PCE

[GJA07] I. B. J. Goudie, P. E. Jupp, and J. Ashbridge. Plantcapture estimation of the size of a homogeneous population. Biometrika, 94(1):243–248, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/1/243;   http://
www.jstor.org/stable/20441366.

Gangnon:2004:SSF

[GK04] Ronald E. Gangnon and Michael R. Kosorok. Sample-size formula for clustered survival data using weighted log-rank statistics. Biometrika, 91(2):263–275, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/263;   http://
www.jstor.org/stable/20441100.

Ganesh:2008:NCA

[GL08] N. Ganesh and P. Lahiri. A new class of average moment matching priors. Biometrika, 95(2):514–520, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/514;
 http://www.jstor.org/stable/20441480.

Glidden:2007:PDD

[Gli07] David V. Glidden. Pairwise dependence diagnostics for clustered failure-time data. Biometrika, 94(2):371–385, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/371;
 http://www.jstor.org/stable/20441378.

Green:2001:DRR

[GM01] Peter J. Green and Antonietta Mira. Delayed rejection in reversible jump Metropolis-Hastings. Biometrika, 88(4):1035–1053, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/1035;   http://www.jstor.org/stable/2673700.

Ghosh:2004:SAE

[GM04] Malay Ghosh and Tapabrata Maiti. Small-area estimation based on natural exponential family quadratic variance function models and survey weights. Biometrika, 91(1):95–112, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/1/95;   http://www.jstor.org/stable/20441081.

Green:2006:BAU

[GM06] Peter J. Green and Kanti V. Mardia. Bayesian alignment using hierarchical models, with applications in protein bioinformatics. Biometrika, 93(2):235–254, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/2/235;   http://
www.jstor.org/stable/20441278.

Ghosh:2008:IFR

[GMR08] Malay Ghosh, Tapabrata Maiti, and Ananya Roy. Influence functions and robust Bayes and empirical Bayes small area estimation. Biometrika, 95(3):573–585, September 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/573;
 http://www.jstor.org/stable/20441487.

Gower:2004:GBS

[Gow04] J. C. Gower. The geometry of biplot scaling. Biometrika, 91 (3):705–714, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/20441132.

Genovese:2006:FDC

[GRW06] Christopher R. Genovese, Kathryn Roeder, and Larry Wasserman. False discovery control with p-value weighting. Biometrika, 93(3):509–524, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/3/509;   http://www.jstor.
org/stable/20441304.

Gerds:2001:FMC

[GS01] T. A. Gerds and M. Schumacher. On functional misspecification of covariates in the Cox regression model. Biometrika, 88(2):572–580, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673501.

Goldstein:2004:BLK

[GS04] Michael Goldstein and Simon C. Shaw. Bayes linear kinematics and Bayes linear Bayes graphical models. Biometrika, 91(2):425–446, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/2/425;   http://www.jstor.org/stable/20441111.

Gander:2007:SIS

[GS07] Matthew P. S. Gander and David A. Stephens. Simulation and inference for stochastic volatility models driven by Lévy processes. Biometrika, 94(3):627–646, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/627;
 http://www.jstor.org/stable/20441401.

Gao:2005:SER

[GT05] Guozhi Gao and Anastasios A. Tsiatis. Semiparametric estimators for the regression coefficients in the linear transformation competing risks model with missing cause of failure. Biometrika, 92(4):875–891, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/4/875;   http://
www.jstor.org/stable/20441242.

Goldwasser:2004:SII

[GTW04] M. A. Goldwasser, L. Tian, and L. J. Wei. Statistical inference for infinite-dimensional parameters via asymptotically pivotal estimating functions. Biometrika, 91(1):81–94, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/1/81;   http://www.jstor.org/stable/20441080.

Guan:2004:SCM

[Gua04] Zhong Guan. A semiparametric changepoint model. Biometrika, 91(4):849–862, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/91/4/849;   http://www.jstor.
org/stable/20441148.

Guan:2008:GFT

[Gua08] Yongtao Guan. A goodness-of-fit test for inhomogeneous spatial Poisson processes. Biometrika, 95(4):831–845, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/831;
 http://www.jstor.org/stable/20441509.

Guan:2009:FBV

[Gua09] Yongtao Guan. Fast block variance estimation procedures for inhomogeneous spatial point processes. Biometrika, 96(1):213–220, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/1/213;   http://www.jstor.org/stable/27798813.

Guo:2009:NAB

[Guo09] Wenge Guo. A note on adaptive Bonferroni and Holm procedures under dependence. Biometrika, 96(4):1012–1018, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/4/1012;   http://www.jstor.org/stable/27798886.

Grzebyk:2004:IMF

[GWC04] Michel Grzebyk, Pascal Wild, and Dominique Chouanière. On identification of multi-factor models with correlated residuals. Biometrika, 91(1):141–151, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/1/141;   http://
www.jstor.org/stable/20441084.

Geyer:2007:AML

[GWS07] Charles J. Geyer, Stuart Wagenius, and Ruth G. Shaw. Aster models for life history analysis. Biometrika, 94(2):415–426, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/415;
 http://www.jstor.org/stable/20441381.

Gaetan:2003:MIM

[GY03] Carlo Gaetan and Jian-Feng Yao. A multiple-imputation Metropolis version of the EM algorithm. Biometrika, 90(3):643–654, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/643;   http://www.jstor.org/stable/30042072.

Gorfine:2006:PSA

[GZH06] Malka Gorfine, David M. Zucker, and Li Hsu. Prospective survival analysis with a general semiparametric shared frailty model: A pseudo full likelihood approach. Biometrika, 93(3):735–741, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/3/735;   http://www.jstor.org/stable/20441319.

Holmes:2003:LIN

[HA03] Christopher C. Holmes and Niall M. Adams. Likelihood inference in nearest-neighbour classification models. Biometrika, 90(1): 99–112, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/90/1/99;   http://www.jstor.org/stable/30042022.

Hald:2000:SHP

[Hal00] Anders Hald. Studies in the history of probability and statistics. XLVII. Pizzetti’s contributions to the statistical analysis of normally distributed observations, 1891. Biometrika, 87(1):213–217, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/1/213;   http://www.jstor.org/stable/2673575.

Hall:2001:BCN

[Hal01] Peter Hall. Biometrika Centenary: Nonparametrics. Biometrika, 88(1):143–165, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673677.

Hanfelt:2003:CRS

[Han03] John J. Hanfelt. Conditioning to reduce the sensitivity of general estimating functions to nuisance parameters. Biometrika, 90(3):517–531, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/3/517;   http://www.jstor.
org/stable/30042063.

Hansen:2008:PAP

[Han08] Ben B. Hansen. The prognostic analogue of the propensity score. Biometrika, 95(2):481–488, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/481;   http://
www.jstor.org/stable/20441477.

Hans:2009:BLR

[Han09] Chris Hans. Bayesian lasso regression. Biometrika, 96(4):835–845, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/835;   http://www.jstor.org/stable/27798870.

Huang:2006:SEM

[HB06] Yijian Huang and Kristin Berry. Semiparametric estimation of marginal mark distribution. Biometrika, 93(4):895–910, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/895;   http://www.jstor.org/stable/20441334.

Houseman:2006:FBD

[HCR06] E. Andrés Houseman, Brent A. Coull, and Louise M. Ryan. A functional-based distribution diagnostic for a linear model with correlated outcomes. Biometrika, 93(4):911–926, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/4/911;
 http://www.jstor.org/stable/20441335.

Henmi:2004:PCN

[HE04] Masayuki Henmi and Shinto Eguchi. A paradox concerning nuisance parameters and projected estimating functions. Biometrika, 91(4):929–941, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/4/929;   http://
www.jstor.org/stable/20441153.

He:2009:NTO

[He09] H. He. A note on time-ordered classification [mr0368317]. Biometrika, 96(1):248, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/96/1/248;   http://www.jstor.
org/stable/27798818.

Helmers:2000:IRE

[Hel00] Roelof Helmers. Inference on rare errors using asymptotic expansions and bootstrap calibration. Biometrika, 87(3):689–694, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/3/689;   http://www.jstor.org/stable/2673639.

Hall:2002:EDS

[HH02] Peter Hall and Nancy E. Heckman. Estimating and depicting the structure of a distribution of random functions. Biometrika, 89(1):145–158, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/1/145;   http://
www.jstor.org/stable/4140564.

Hwang:2005:EEH

[HH05] Wen-Han Hwang and Richard Huggins. An examination of the effect of heterogeneity on the estimation of population size using capture-recapture data. Biometrika, 92(1):229–233, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/229;
 http://www.jstor.org/stable/20441178.

Huang:2007:HSM

[HH07] Yifan Huang and Jason C. Hsu. Hochberg’s step-up method: Cutting corners off Holm’s step-down method. Biometrika, 94(4):965– 975, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/4/965;   http://www.jstor.org/stable/20441429.

Hall:2004:NCI

[HHF04] Peter Hall, Rob J. Hyndman, and Yanan Fan. Nonparametric confidence intervals for receiver operating characteristic curves. Biometrika, 91(3):743–750, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/20441136.

Hobert:2002:ARS

[HJPR02] James P. Hobert, Galin L. Jones, Brett Presnell, and Jeffrey S. Rosenthal. On the applicability of regenerative simulation in Markov chain Monte Carlo. Biometrika, 89(4):731–743, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/4/731;   http://www.jstor.org/stable/4140533.

Heagerty:2001:MML

[HK01] Patrick J. Heagerty and Brenda F. Kurland. Misspecified maximum likelihood estimates and generalised linear mixed models. Biometrika, 88(4):973–985, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/88/4/973;   http://
www.jstor.org/stable/2673696.

Hall:2002:STE

[HL02a] W. J. Hall and Aiyi Liu. Sequential tests and estimators after overrunning based on maximum-likelihood ordering. Biometrika, 89(3):699–707, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/699;   http://
www.jstor.org/stable/4140612.

Hickernell:2002:UDL

[HL02b] Fred J. Hickernell and Min-Qian Liu. Uniform designs limit aliasing. Biometrika, 89(4):893–904, December 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/4/893;
 http://www.jstor.org/stable/4140545.

Ho:2005:CIC

[HL05] Yvonne H. S. Ho and Stephen M. S. Lee. Calibrated interpolated confidence intervals for population quantiles. Biometrika, 92(1): 234–241, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/1/234;   http://www.jstor.org/stable/20441179.

Hall:2006:AC

[HL06a] Peter Hall and Ming Li. Amendments and corrections: “Using the Periodogram to Estimate Period in Nonparametric Regression”. Biometrika, 93(4):1026, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/4/1025.3. See [HL06b].

Hall:2006:UPE

[HL06b] Peter Hall and Ming Li. Using the periodogram to estimate period in nonparametric regression. Biometrika, 93(2):411–424, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/2/411;   http://www.jstor.org/stable/20441291. See

correction [HL06a, HNPE07].

Hannig:2009:GFI

[HL09] Jan Hannig and Thomas C. M. Lee. Generalized fiducial inference for wavelet regression. Biometrika, 96(4):847–860, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/847;
 http://www.jstor.org/stable/27798871.

Hu:2009:MAP

[HLL09] X. Joan Hu, Stephen W. Lagakos, and Richard A. Lockhart. Marginal analysis of panel counts through estimating functions. Biometrika, 96(2):445–456, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/445;
 http://www.jstor.org/stable/27798835.

Huang:2006:CMS

[HLPL06] Jianhua Z. Huang, Naiping Liu, Mohsen Pourahmadi, and Linxu Liu. Covariance matrix selection and estimation via penalised normal likelihood. Biometrika, 93(1):85–98, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/1/85;   http://www.jstor.org/stable/20441262.

Huang:2009:GBA

[HMXZ09a] Jian Huang, Shuange Ma, Huiliang Xie, and Cun-Hui Zhang. A group bridge approach for variable selection. Biometrika, 96(2): 339–355, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/96/2/339;   http://www.jstor.org/stable/27798828.

See amendments and corrections [HMXZ09b].

Huang:2009:ACG

[HMXZ09b] Jian Huang, Shuangge Ma, Huiliang Xie, and Cun-Hui Zhang. Amendments and corrections: “A group bridge approach for variable selection” [MR2507147]. Biometrika, 96(4):1024, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/27798888;
http://www.jstor.org/stable/27798889. See [HMXZ09a].

Hrafnkelsson:2000:ASC

[HN00] Birgir Hrafnkelsson and H. Joseph Newton. Asymptotic simultaneous confidence bands for vector autoregressive spectra. Biometrika, 87(1):173–182, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/1/173;   http://
www.jstor.org/stable/2673570.

Heimann:2001:ADP

[HN01] Günter Heimann and Georg Neuhaus. On the asymptotic distribution for Peto’s combined test for carcinogenicity assays under equal and unequal censoring. Biometrika, 88(2):435–445, June 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673491.

Hall:2006:EBD

[HN06] Peter Hall and Natalie Neumeyer. Estimating a bivariate density when there are extra data on one or both components. Biometrika, 93(2):439–450, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/2/439;   http://
www.jstor.org/stable/20441293.

Hall:2005:NIM

[HNPE05] Peter Hall, Amnon Neeman, Reza Pakyari, and Ryan Elmore. Nonparametric inference in multivariate mixtures. Biometrika, 92(3):667–678, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/3/667;   http://www.jstor.
org/stable/20441221. See [Ano07a].

Hall:2007:ACN

[HNPE07] Peter Hall, Amnon Neeman, Reza Pakyari, and Ryan Elmore. Amendments and corrections: “Nonparametric inference in multivariate mixtures” Biometrika (2005), 92, pp. 667–678 [MR2202653]. Biometrika, 94(3):767, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://biomet.oxfordjournals.org/content/94/3/767. See

[HL06b].

Hall:2005:TPS

[HO05] Peter Hall and J. D. Opsomer. Theory for penalised spline regression. Biometrika, 92(1):105–118, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/105;
 http://www.jstor.org/stable/20441169.

Hoff:2003:BMP

[Hof03] Peter D. Hoff. Bayesian methods for partial stochastic orderings. Biometrika, 90(2):303–317, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/303;   http://
www.jstor.org/stable/30042041.

Hughes-Oliver:2000:EEP

[HOR00] Jacqueline M. Hughes-Oliver and William F. Rosenberger. Efficient estimation of the prevalence of multiple rare traits. Biometrika, 87(2):315–327, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/315;   http://
www.jstor.org/stable/2673466.

Hall:2001:ORS

[HP01] Daniel B. Hall and Jens T. Præstgaard. Order-restricted score tests for homogeneity in generalised linear and nonlinear mixed models. Biometrika, 88(3):739–751, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673443.

Huang:2009:SME

[HP09] Ying Huang and Margaret Sullivan Pepe. Semiparametric methods for evaluating risk prediction markers in case-control studies. Biometrika, 96(4):991–997, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/991;
 http://www.jstor.org/stable/27798883.

Hall:2005:DTA

[HQ05] Peter Hall and Peihua Qiu. Discrete-transform approach to deconvolution problems. Biometrika, 92(1):135–148, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/135;
 http://www.jstor.org/stable/20441171.

Hettmansperger:2002:PAE

[HR02] Thomas P. Hettmansperger and Ronald H. Randles. A practical affine equivariant multivariate median. Biometrika, 89(4):851–860, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/851;   http://www.jstor.org/stable/4140542.

Hall:2009:RVC

[HR09] Peter Hall and Andrew P. Robinson. Reducing variability of crossvalidation for smoothing-parameter choice. Biometrika, 96(1):175– 186, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/96/1/175;   http://www.jstor.org/stable/27798810.

Hall:2000:NEP

[HRR00] Peter Hall, James Reimann, and John Rice. Nonparametric estimation of a periodic function. Biometrika, 87(3):545–557, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/3/545;   http://www.jstor.org/stable/2673629.

Hodges:2001:CDF

[HS01] James S. Hodges and Daniel J. Sargent. Counting degrees of freedom in hierarchical and other richly-parameterised models. Biometrika, 88(2):367–379, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673485.

Henderson:2003:SCG

[HS03] Robin Henderson and Silvia Shimakura. A serially correlated gamma frailty model for longitudinal count data. Biometrika, 90(2):355–366, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/2/355;   http://www.jstor.
org/stable/30042045.

Harel:2009:PLI

[HS09] Ofer Harel and Joseph L. Schafer. Partial and latent ignorability in missing-data problems. Biometrika, 96(1):37–50, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/1/37;   http://www.jstor.org/stable/27798800.

Huang:2006:LMR

[HSD06] Xianzheng Huang, Leonard A. Stefanski, and Marie Davidian. Latent-model robustness in structural measurement error models. Biometrika, 93(1):53–64, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/93/1/53;   http://
www.jstor.org/stable/20441259.

He:2000:BPT

[HSW00] Xuming He, Douglas G. Simpson, and Guangyu Wang. Breakdown points of t-type regression estimators. Biometrika, 87(3):675–687, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/3/675;   http://www.jstor.org/stable/2673638.

Hoffman:2001:WCR

[HSW01] Elaine B. Hoffman, Pranab K. Sen, and Clarice R. Weinberg. Within-cluster resampling. Biometrika, 88(4):1121–1134, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/1121;   http://www.jstor.org/stable/2673705.

Hall:2002:PTE

[HT02] Peter Hall and Nader Tajvidi. Permutation tests for equality of distributions in high-dimensional settings. Biometrika, 89(2):359– 374, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/2/359;   http://www.jstor.org/stable/4140582.

Huggins:2006:UNE

[Hug06] Richard Huggins. Understanding nonparametric estimation for clustered data. Biometrika, 93(2):486–489, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/486;
 http://www.jstor.org/stable/20441299.

Hubbard:2008:PIM

[HV08] Alan E. Hubbard and Mark J. Van Der Laan. Population intervention models in causal inference. Biometrika, 95(1):35–47, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/1/35;   http://www.jstor.org/stable/20441441.

Haywood:2000:ISS

[HW00] John Haywood and Granville Tunnicliffe Wilson. An improved state space representation for cyclical time series. Biometrika, 87(3):724–726, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/3/724;   http://www.jstor.
org/stable/2673645.

Hu:2004:PKV

[HWC04] Zonghui Hu, Naisyin Wang, and Raymond J. Carroll. Profilekernel versus backfitting in the partially linear models for longitudinal/clustered data. Biometrika, 91(2):251–262, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/251;
 http://www.jstor.org/stable/20441099.

Huang:2002:VCM

[HWZ02] Jianhua Z. Huang, Colin O. Wu, and Lan Zhou. Varying-coefficient models and basis function approximations for the analysis of repeated measurements. Biometrika, 89(1):111–128, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/1/111;
 http://www.jstor.org/stable/4140562.

Huang:2006:APC

[HWZ06] Chiung-Yu Huang, Mei-Cheng Wang, and Ying Zhang. Analysing panel count data with informative observation times. Biometrika, 93(4):763–775, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/4/763;   http://www.jstor.
org/stable/20441326.

Hall:2003:IAS

[HY03] W. J. Hall and Benjamin Yakir. Inference about a secondary process following a sequential trial. Biometrika, 90(3):597–611, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/597;   http://www.jstor.org/stable/30042069.

Hardouin:2008:MPA

[HY08] Cécile Hardouin and Jian-Feng Yao. Multi-parameter automodels and their applications. Biometrika, 95(2):335–349, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/335;
 http://www.jstor.org/stable/20441468.

Henmi:2007:ISE

[HYE07] Masayuki Henmi, Ryo Yoshida, and Shinto Eguchi. Importance sampling via the estimated sampler. Biometrika, 94(4):985–991, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/4/985;   http://www.jstor.org/stable/20441431.

Heckman:2000:CSR

[HZ00] Nancy E. Heckman and Ruben H. Zamar. Comparing the shapes of regression functions. Biometrika, 87(1):135–144, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/135;
 http://www.jstor.org/stable/2673567.

He:2002:ESM

[HZF02] Xuming He, Zhong-Yi Zhu, and Wing-Kam Fung. Estimation in a semiparametric model for longitudinal data with unspecified dependence structure. Biometrika, 89(3):579–590, August 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/3/579;
 http://www.jstor.org/stable/4140603.

Ibrahim:2001:MRG

[ICL01] Joseph G. Ibrahim, Ming-Hui Chen, and Stuart R. Lipsitz. Missing responses in generalised linear mixed models when the missing data mechanism is nonignorable. Biometrika, 88(2):551–564, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673499.

Imbens:2000:RPS

[Imb00] Guido W. Imbens. The role of the propensity score in estimating dose-response functions. Biometrika, 87(3):706–710, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/3/706;   http://www.jstor.org/stable/2673642.

Ip:2003:TET

[IWLA03] Wai-Cheung Ip, Heung Wong, Yuan Li, and Hongzhi An. Testing and estimation of thresholds based on wavelets in heteroscedastic threshold autoregressive models. Biometrika, 90(3):703–716, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/703;   http://www.jstor.org/stable/30042076.

Ishwaran:2000:MCM

[IZ00] Hemant Ishwaran and Mahmoud Zarepour. Markov chain Monte Carlo in approximate Dirichlet and beta two-parameter process hierarchical models. Biometrika, 87(2):371–390, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/371;
 http://www.jstor.org/stable/2673470.

Joseph:2009:BIM

[JAW09] V. Roshan Joseph, Mingyao Ai, and C. F. Jeff Wu. Bayesianinspired minimum aberration two- and four-level designs. Biometrika, 96(1):95–106, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/96/1/95;   http://www.jstor.
org/stable/27798804.

Jeffries:2003:NTN

[Jef03] Neal O. Jeffries. A note on: “Testing the number of components in a normal mixture” [Biometrika 88 (2001), no. 3, 767–778; MR1859408] by Y. Lo, N. R. Mendell, and D. B. Rubin. Biometrika, 90(4):991–994, December 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/991;
 http://www.jstor.org/stable/30042105.

Jeong:2009:NCS

[JF09] J.-H. Jeong and J. P. Fine. A note on cause-specific residual life. Biometrika, 96(1):237–242, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/1/237;   http://
www.jstor.org/stable/27798816.

Jones:2009:ODS

[JG09] Bradley Jones and Peter Goos. D-optimal design of splitsplit-plot experiments. Biometrika, 96(1):67–82, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/1/67;   http://www.jstor.org/stable/27798802.

Jones:2007:KTD

[JH07] M. C. Jones and D. A. Henderson. Kernel-type density estimation on the unit interval. Biometrika, 94(4):977–984, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/4/977;
 http://www.jstor.org/stable/20441430.

Jones:2001:CRD

[JHHB01] M. C. Jones, Nils Lid Hjort, Ian R. Harris, and Ayanendranath Basu. A comparison of related density-based minimum divergence estimators. Biometrika, 88(3):865–873, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673453.

James:2000:PCM

[JHS00] Gareth M. James, Trevor J. Hastie, and Catherine A. Sugar. Principal component models for sparse functional data. Biometrika, 87(3):587–602, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/3/587;   http://www.jstor.
org/stable/2673632.

Jungbacker:2007:MCE

[JK07] Borus Jungbacker and Siem Jan Koopman. Monte Carlo estimation for nonlinear non-Gaussian state space models. Biometrika, 94(4):827–839, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/4/827;   http://
www.jstor.org/stable/20441420.

Jones:2004:MDS

[JL04] M. C. Jones and P. V. Larsen. Multivariate distributions with support above the diagonal. Biometrika, 91(4):975–986, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/4/975;   http://www.jstor.org/stable/20441156.

Jin:2003:RBI

[JLWY03] Zhezhen Jin, D. Y. Lin, L. J. Wei, and Zhiliang Ying. Rank-based inference for the accelerated failure time model. Biometrika, 90(2):341–353, June 2003. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/341;   http://
www.jstor.org/stable/30042044.

Jin:2006:LSR

[JLY06] Zhezhen Jin, D. Y. Lin, and Zhiliang Ying. On least-squares regression with censored data. Biometrika, 93(1):147–161, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/1/147;
 http://www.jstor.org/stable/20441267.

John:2001:UFA

[Joh01] J. A. John. Updating formula in an analysis of variance model. Biometrika, 88(4):1175–1178, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/88/4/1175;   http://
www.jstor.org/stable/2673709.

Joseph:2004:ERM

[Jos04] V. Roshan Joseph. Efficient Robbins-Monro procedure for binary data. Biometrika, 91(2):461–470, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/461;
 http://www.jstor.org/stable/20441113.

Janes:2009:ACE

[JP09a] Holly Janes and Margaret S. Pepe. Adjusting for covariate effects on classification accuracy using the covariate-adjusted receiver operating characteristic curve. Biometrika, 96(2):371–382, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/371;
 http://www.jstor.org/stable/27798830.

Jones:2009:SAD

[JP09b] M. C. Jones and Arthur Pewsey. Sinh-arcsinh distributions. Biometrika, 96(4):761–780, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/4/761;   http://
www.jstor.org/stable/27798865.

James:2009:GDS

[JR09] Gareth M. James and Peter Radchenko. A generalized Dantzig selector with shrinkage tuning. Biometrika, 96(2):323–337, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/2/323;   http://www.jstor.org/stable/27798827.

Jowaheer:2002:ALC

[JS02] Vandna Jowaheer and Brajendra C. Sutradhar. Analysing longitudinal count data with overdispersion. Biometrika, 89(2):389–399, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/2/389;   http://www.jstor.org/stable/4140584.

Johnson:2009:ISS

[JS09] Lynn M. Johnson and Robert L. Strawderman. Induced smoothing for the semiparametric accelerated failure time model: asymptotics and extensions to clustered data. Biometrika, 96(3):577–590, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/3/577;   http://www.jstor.org/stable/27798849.

Jasra:2007:PBR

[JSH07] Ajay Jasra, David A. Stephens, and Christopher C. Holmes. Population-based reversible jump Markov chain Monte Carlo. Biometrika, 94(4):787–807, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/4/787;   http://
www.jstor.org/stable/20441418.

Johnson:2005:SIO

[JT05] Brent A. Johnson and Anastasios A. Tsiatis. Semiparametric inference in observational duration-response studies, with duration possibly right-censored. Biometrika, 92(3):605–618, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/605;
 http://www.jstor.org/stable/20441217.

Jennison:2006:ANG

[JT06] Christopher Jennison and Bruce W. Turnbull. Adaptive and nonadaptive group sequential tests. Biometrika, 93(1):1–21, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/1/1;   http://www.jstor.org/stable/20441256.

Jewell:2004:CCC

[JV04] Nicholas P. Jewell and Mark Van Der Laan. Case-control current status data. Biometrika, 91(3):529–541, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/20441121.

Jewell:2003:NEC

[JvdLH03] Nicholas P. Jewell, Mark van der Laan, and Tanya Henneman. Nonparametric estimation from current status data with competing risks. Biometrika, 90(1):183–197, March 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/1/183;
 http://www.jstor.org/stable/30042028.

Jewell:2005:BCS

[JVL05] Nicholas P. Jewell, Mark Van Der Laan, and Xiudong Lei. Bivariate current status data with univariate monitoring times. Biometrika, 92(4):847–862, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/4/847;   http://
www.jstor.org/stable/20441240.

Jones:2005:CDU

[JW05] Beatrix Jones and Mike West. Covariance decomposition in undirected Gaussian graphical models. Biometrika, 92(4):779–786, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/4/779;   http://www.jstor.org/stable/20441235.

Jung:2003:RBR

[JY03] Sin-Ho Jung and Zhiliang Ying. Rank-based regression with repeated measurements data. Biometrika, 90(3):732–740, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/3/732;   http://www.jstor.org/stable/30042080.

Jin:2001:SRM

[JYW01] Zhezhen Jin, Zhiliang Ying, and L. J. Wei. A simple resampling method by perturbing the minimand. Biometrika, 88(2):381–390, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673486.

Jiang:2001:REG

[JZ01] Jiming Jiang and Weihong Zhang. Robust estimation in generalised linear mixed models. Biometrika, 88(3):753–765, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673444.

Jiang:2007:AHR

[JZ07] Jiancheng Jiang and Haibo Zhou. Additive hazard regression with auxiliary covariates. Biometrika, 94(2):359–369, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/359;
 http://www.jstor.org/stable/20441377.

Kagan:2001:NLL

[Kag01] Abram Kagan. A note on the logistic link function. Biometrika, 88(2):599–601, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673505.

Konishi:2004:BIC

[KAI04] Sadanori Konishi, Tomohiro Ando, and Seiya Imoto. Bayesian information criteria and smoothing parameter selection in radial basis function networks. Biometrika, 91(1):27–43, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/1/27;   http://www.jstor.org/stable/20441077.

King:2001:BAP

[KB01] R. King and S. P. Brooks. On the Bayesian analysis of population size. Biometrika, 88(2):317–336, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673482.

King:2002:BMD

[KB02] R. King and S. P. Brooks. Bayesian model discrimination for multiple strata capture-recapture data. Biometrika, 89(4):785–806, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/785;   http://www.jstor.org/stable/4140537.

King:2003:CFL

[KB03] R. King and S. P. Brooks. Closed-form likelihoods for Arnason-Schwarz models. Biometrika, 90(2):435–444, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/435;
 http://www.jstor.org/stable/30042051.

Kuroki:2006:RPC

[KC06] Manabu Kuroki and Zhihong Cai. On recovering a population covariance matrix in the presence of selection bias. Biometrika, 93(3):601–611, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/3/601;   http://
www.jstor.org/stable/20441310.

Kang:2009:MHM

[KC09] S. Kang and J. Cai. Marginal hazards model for case-cohort studies with multiple disease outcomes. Biometrika, 96(4):887–901, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/887;   http://www.jstor.org/stable/27798874.

Kim:2008:FGL

[KCD08] Sungduk Kim, Ming-Hui Chen, and Dipak K. Dey. Flexible generalized t-link models for binary response data. Biometrika, 95(1): 93–106, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/1/93;   http://www.jstor.org/stable/20441445.

Kong:2004:WEE

[KCS04] Lan Kong, Jianwen Cai, and Pranab K. Sen. Weighted estimating equations for semiparametric transformation models with censored data from a case-cohort design. Biometrika, 91(2):305–319, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/2/305;   http://www.jstor.org/stable/20441103.

Kent:2000:UCS

[KDA00] John T. Kent, Ian L. Dryden, and Catherine R. Anderson. Using circulant symmetry to model featureless objects. Biometrika, 87(3):527–544, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/3/527;   http://
www.jstor.org/stable/2673628.

Kume:2007:SSS

[KDL07] Alfred Kume, Ian L. Dryden, and Huiling Le. Shape-space smoothing splines for planar landmark data. Biometrika, 94(3):513–528, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/3/513;   http://www.jstor.org/stable/20441393.

Kelderman:2004:MEN

[Kel04] Henk Kelderman. Measurement exchangeability and normal onefactor models. Biometrika, 91(3):738–742, September 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441135.

Kim:2004:FHD

[KF04] Jae Kwang Kim and Wayne Fuller. Fractional hot deck imputation. Biometrika, 91(3):559–578, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/20441123.

Kosmidis:2009:BRE

[KF09] Ioannis Kosmidis and David Firth. Bias reduction in exponential family nonlinear models. Biometrika, 96(4):793–804, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/4/793;   http://www.jstor.org/stable/27798867.

Kempton:2001:OCD

[KFD01] R. A. Kempton, S. J. Ferris, and O. David. Optimal change-over designs when carry-over effects are proportional to direct effects of treatments. Biometrika, 88(2):391–399, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673487.

Kim:2002:NAB

[Kim02] J. K. Kim. A note on approximate Bayesian bootstrap imputation. Biometrika, 89(2):470–477, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/470;   http://
www.jstor.org/stable/4140592.

Kim:2004:NDC

[KKPS04] Tae Yoon Kim, Donghoh Kim, Byeong U. Park, and Douglas G. Simpson. Nonparametric detection of correlated errors. Biometrika, 91(2):491–496, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/491;   http://
www.jstor.org/stable/20441115.

Kulich:2000:AHR

[KL00] Michal Kulich and D. Y. Lin. Additive hazards regression for case-cohort studies. Biometrika, 87(1):73–87, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/1/73;   http://www.jstor.org/stable/2673562.

Kulinskaya:2009:FFE

[KL09] Elena Kulinskaya and Alex Lewin. On fuzzy familywise error rate and false discovery rate procedures for discrete distributions. Biometrika, 96(1):201–211, March 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/1/201;
 http://www.jstor.org/stable/27798812.

Kent:2001:SPT

[KM01] J. T. Kent and K. V. Mardia. Shape, Procrustes tangent projections and bilateral symmetry. Biometrika, 88(2):469–485, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673494. See cor-

rection [KM05].

Kent:2005:ACS

[KM05] J. T. Kent and K. V. Mardia. Amendments and corrections: “Shape, Procrustes tangent projections and bilateral symmetry” [Biometrika 88 (2001), no. 2, 469–485 [MR1844846]]. Biometrika, 92(1):249, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/1/249.1;   http://www.jstor.
org/stable/20441181. See [KM01].

Kou:2009:AP

[KM09] S. C. Kou and P. McCullagh. Approximating the α-permanent. Biometrika, 96(3):635–644, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/3/635;   http://
www.jstor.org/stable/27798853.

Kenward:2003:PMM

[KMT03] M. G. Kenward, G. Molenberghs, and H. Thijs. Pattern-mixture models with proper time dependence. Biometrika, 90(1):53–71, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/1/53;   http://www.jstor.org/stable/30042019.

Kolaczyk:2005:MGL

[KN05] Eric D. Kolaczyk and Robert D. Nowak. Multiscale generalised linear models for nonparametric function estimation. Biometrika, 92(1):119–133, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/1/119;   http://
www.jstor.org/stable/20441170.

Kuang:2008:FAP

[KNN08a] D. Kuang, B. Nielsen, and J. P. Nielsen. Forecasting with the age-period-cohort model and the extended chain-ladder model. Biometrika, 95(4):987–991, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/987;   http://
www.jstor.org/stable/20441520.

Kuang:2008:IAP

[KNN08b] D. Kuang, B. Nielsen, and J. P. Nielsen. Identification of the age-period-cohort model and the extended chain-ladder model. Biometrika, 95(4):979–986, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/979;   http://
www.jstor.org/stable/20441519.

Kennedy:2000:POC

[KO00] M. C. Kennedy and A. O’Hagan. Predicting the output from a complex computer code when fast approximations are available. Biometrika, 87(1):1–13, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/1/1;   http://www.jstor.
org/stable/2673557.

Komaki:2001:SPD

[Kom01] Fumiyasu Komaki. A shrinkage predictive distribution for multivariate normal observables. Biometrika, 88(3):859–864, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673452.

Kosorok:1999:TSQ

[Kos99] Michael R. Kosorok. Two-sample quantile tests under general conditions. Biometrika, 86(4):909–921, December 1999. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673594. See correction

[Kos01a].

Kosorok:2001:ACT

[Kos01a] Michael R. Kosorok. Amendments and corrections: “Two-sample quantile tests under general conditions” [Biometrika 86 (1999), no. 4, 909–921; 1741986]. Biometrika, 88(4):1201, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/88/4/1201.

See [Kos99].

Kosorok:2001:TSQ

[Kos01b] Michael R. Kosorok. Two-sample quantile tests under general conditions. Biometrika, 88(4):1201, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673713.

Kowalski:2004:NIS

[KP04] Jeanne Kowalski and Jonathan Powell. Nonparametric inference for stochastic linear hypotheses: Application to highdimensional data. Biometrika, 91(2):393–408, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/393;
 http://www.jstor.org/stable/20441109.

Kim:2009:UAL

[KR09] Jae Kwang Kim and J. N. K. Rao. A unified approach to linearization variance estimation from survey data after imputation for item nonresponse. Biometrika, 96(4):917–932, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/917;
 http://www.jstor.org/stable/27798876.

Kim:2006:VSC

[KTV06] Sinae Kim, Mahlet G. Tadesse, and Marina Vannucci. Variable selection in clustering via Dirichlet process mixture models.

Biometrika, 93(4):877–893, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/4/877;   http://
www.jstor.org/stable/20441333.

Kuk:2004:PIA

[Kuk04] Anthony Y. C. Kuk. Permutation invariance of alternating logistic regression for multivariate binary data. Biometrika, 91(3): 758–761, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/20441138.

Kuk:2007:HPL

[Kuk07] Anthony Y. C. Kuk. A hybrid pairwise likelihood method. Biometrika, 94(4):939–952, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/4/939;   http://
www.jstor.org/stable/20441427.

Kuroki:2007:GIC

[Kur07] Manabu Kuroki. Graphical identifiability criteria for causal effects in studies with an unobserved treatment/response variable. Biometrika, 94(1):37–47, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/94/1/37;   http://
www.jstor.org/stable/20441352.

Kume:2005:SAB

[KW05] A. Kume and Andrew T. A. Wood. Saddlepoint approximations for the Bingham and Fisher-Bingham normalising constants. Biometrika, 92(2):465–476, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/465;   http://
www.jstor.org/stable/20441200.

Kong:2007:VSS

[KX07] Efang Kong and Yingcun Xia. Variable selection for the singleindex model. Biometrika, 94(1):217–229, March 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/1/217;
 http://www.jstor.org/stable/20441364.

Lazar:2003:BEL

[Laz03] Nicole A. Lazar. Bayesian empirical likelihood. Biometrika, 90(2): 319–326, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/90/2/319;   http://www.jstor.org/stable/30042042.

Lawless:2006:MIC

[LB06] J. F. Lawless and Denise Babineau. Models for interval censoring and simulation-based inference for lifetime distributions. Biometrika, 93(3):671–686, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/3/671;   http://
www.jstor.org/stable/20441315.

Larsen:2002:ILR

[LBS02] P. V. Larsen, P. Blæsild, and M. K. Sørensen. Improved likelihood ratio tests on the von Mises–Fisher distribution. Biometrika, 89(4):947–951, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/947;   http://
www.jstor.org/stable/4140550.

Lin:2001:SRC

[LC01] Xihong Lin and Raymond J. Carroll. Semiparametric regression for clustered data. Biometrika, 88(4):1179–1185, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
88/4/1179;   http://www.jstor.org/stable/2673710.

Levine:2003:IMP

[LC03] Richard A. Levine and George Casella. Implementing matching priors for frequentist inference. Biometrika, 90(1):127–137, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/1/127;   http://www.jstor.org/stable/30042024.

Li:2009:NFF

[LCM09] P. Li, J. Chen, and P. Marriott. Non-finite Fisher information and homogeneity: an EM approach. Biometrika, 96(2):411–426, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/2/411;   http://www.jstor.org/stable/27798833.

Li:2007:PIR

[LCT07] Lexin Li, R. Dennis Cook, and Chih-Ling Tsai. Partial inverse regression. Biometrika, 94(3):615–625, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/615;
 http://www.jstor.org/stable/20441400.

Lee:2006:ICR

[Lee06] Sokbae Lee. Identification of a competing risks model with unknown transformations of latent failure times. Biometrika, 93(4):996–1002, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/4/996;   http://www.jstor.
org/stable/20441341.

Lawless:2005:FPI

[LF05] J. F. Lawless and Marc Fredette. Frequentist prediction intervals and predictive distributions. Biometrika, 92(3):529–542, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/3/529;   http://www.jstor.org/stable/20441212.

Li:2008:TCS

[LGS08] Bo Li, Marc G. Genton, and Michael Sherman. Testing the covariance structure of multivariate random fields. Biometrika, 95 (4):813–829, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/95/4/813;   http://www.jstor.
org/stable/20441508.

Liu:2001:UES

[LH01] Aiyi Liu and W. J. Hall. Unbiased estimation of secondary parameters following a sequential test. Biometrika, 88(3):895–900, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2673457.

Lin:2006:REE

[LH06] Nan Lin and Xuming He. Robust and efficient estimation under data grouping. Biometrika, 93(1):99–112, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/1/99;   http://www.jstor.org/stable/20441263.

Li:2007:SSD

[Li07] Lexin Li. Sparse sufficient dimension reduction. Biometrika, 94(3): 603–613, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/3/603;   http://www.jstor.org/stable/20441399.

Lieberman:2001:PML

[Lie01] Offer Lieberman. Penalised maximum likelihood estimation for fractional Gaussian processes. Biometrika, 88(3):888–894, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673456.

Lin:2000:FCP

[Lin00] D. Y. Lin. On fitting Cox’s proportional hazards models to survey data. Biometrika, 87(1):37–47, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/87/1/37;   http://
www.jstor.org/stable/2673559.

Lin:2008:EES

[Lin08] Pei-Sheng Lin. Estimating equations for spatially correlated data in multi-dimensional space. Biometrika, 95(4):847–858, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/4/847;   http://www.jstor.org/stable/20441510.

Lin:2003:DAT

[LJ03] Y. Lin and Y. Jeon. Discriminant analysis through a semiparametric model. Biometrika, 90(2):379–392, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/379;
 http://www.jstor.org/stable/30042047.

Lin:2007:ISI

[LK07] Wei Lin and K. B. Kulasekera. Identifiability of single-index models and additive-index models. Biometrika, 94(2):496–501, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/496;
 http://www.jstor.org/stable/20441387.

Lewin-Koh:2003:HFA

[LKA03] Sock-Cheng Lewin-Koh and Yasuo Amemiya. Heteroscedastic factor analysis. Biometrika, 90(1):85–97, March 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/1/85;
 http://www.jstor.org/stable/30042021.

Lee:2006:SMS

[LKLK06] Yoonkyung Lee, Yuwon Kim, Sangjun Lee, and Ja-Yong Koo. Structured multicategory support vector machines with analysis of variance decomposition. Biometrika, 93(3):555–571, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/3/555;   http://www.jstor.org/stable/20441307.

Li:2000:CME

[LL00] Yi Li and Xihong Lin. Covariate measurement errors in frailty models for clustered survival data. Biometrika, 87(4):849–866, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/849;   http://www.jstor.org/stable/2673614.

Liao:2002:TRM

[LL02] J. G. Liao and Stuart R. Lipsitz. A type of restricted maximum likelihood estimator of variance components in generalised linear mixed models. Biometrika, 89(2):401–409, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/401;
 http://www.jstor.org/stable/4140585.

Lee:2004:RSL

[LL04] Jaechoul Lee and Robert Lund. Revisiting simple linear regression with autocorrelated errors. Biometrika, 91(1):240–245, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/1/240;   http://www.jstor.org/stable/20441093.

Li:2005:DCT

[LL05] Guodong Li and Wai Keung Li. Diagnostic checking for time series models with conditional heteroscedasticity estimated by the least absolute deviation approach. Biometrika, 92(3):691–701, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/3/691;   http://www.jstor.org/stable/20441223.

Lai:2006:CIG

[LL06] Tze Leung Lai and Wenzhi Li. Confidence intervals in group sequential trials with random group sizes and applications to survival analysis. Biometrika, 93(3):641–654, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/641;
 http://www.jstor.org/stable/20441313.

Li:2008:LAD

[LL08a] Guodong Li and Wai Keung Li. Least absolute deviation estimation for fractionally integrated autoregressive moving average time series models with conditional heteroscedasticity. Biometrika, 95(2):399–414, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/399;   http://
www.jstor.org/stable/20441472.

Ling:2008:AIN

[LL08b] Shiqing Ling and Dong Li. Asymptotic inference for a nonstationary double AR(1) model. Biometrika, 95(1):257–263, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/257;
 http://www.jstor.org/stable/20441458.

Lee:2009:DBB

[LL09] Stephen M. S. Lee and P. Y. Lai. Double block bootstrap confidence intervals for dependent data. Biometrika, 96(2):427–443, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/2/427;   http://www.jstor.org/stable/27798834.

Liechty:2004:BCE

[LLM04] John C. Liechty, Merrill W. Liechty, and Peter Müller. Bayesian correlation estimation. Biometrika, 91(1):1–14, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/1;
http://www.jstor.org/stable/20441075.

Lloyd:2000:MLE

[Llo00] Chris J. Lloyd. Maximum likelihood estimation of misclassification rates of a binomial regression. Biometrika, 87(3):700–705, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/3/700;   http://www.jstor.org/stable/2673641.

Li:2001:EPN

[LLW01] W. K. Li, Shiqing Ling, and H. Wong. Estimation for partially nonstationary multivariate autoregressive models with conditional heteroscedasticity. Biometrika, 88(4):1135–1152, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
88/4/1135;   http://www.jstor.org/stable/2673706.

Li:2009:SRO

[LM09] Gang Li and Dibyen Majumdar. Some results on D-optimal designs for nonlinear models with applications. Biometrika, 96(2):487–493, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/2/487;   http://www.jstor.org/stable/27798839.

Lahiri:2007:RBE

[LMKP07] Soumendra N. Lahiri, Tapabrata Maiti, Myron Katzoff, and Van Parsons. Resampling-based empirical prediction: An application to small area estimation. Biometrika, 94(2):469–485, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/469;
 http://www.jstor.org/stable/20441385.

Liseo:2001:BSI

[LMP01] Brunero Liseo, Domenico Marinucci, and Lea Petrella. Bayesian semiparametric inference on long-range dependence. Biometrika, 88(4):1089–1104, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/88/4/1089;   http://www.jstor.
org/stable/2673703.

Lijoi:2007:BNE

[LMP07] Antonio Lijoi, Ramsés H. Mena, and Igor Prünster. Bayesian nonparametric estimation of the probability of discovering new species. Biometrika, 94(4):769–786, December 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/4/769;
 http://www.jstor.org/stable/20441417.

Lo:2001:TNC

[LMR01] Yungtai Lo, Nancy R. Mendell, and Donald B. Rubin. Testing the number of components in a normal mixture. Biometrika, 88 (3):767–778, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673445.

Lin:2009:CON

[LMT09] C. Devon Lin, Rahul Mukerjee, and Boxin Tang. Construction of orthogonal and nearly orthogonal Latin hypercubes. Biometrika, 96(1):243–247, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/1/243;   http://
www.jstor.org/stable/27798817.

Lee:2001:HGL

[LN01] Youngjo Lee and John A. Nelder. Hierarchical generalised linear models: A synthesis of generalised linear models, random-effect models and structured dispersions. Biometrika, 88(4):987–1006, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/987;   http://www.jstor.org/stable/2673697.

Larocque:2007:WMS

[LNO07] Denis Larocque, Jaakko Nevalainen, and Hannu Oja. A weighted multivariate sign test for cluster-correlated data. Biometrika, 94(2):267–283, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/2/267;   http://
www.jstor.org/stable/20441371.

Lockhart:2007:UGS

[LOS07] Richard A. Lockhart, Federico J. O’Reilly, and Michael A. Stephens. Use of the Gibbs sampler to obtain conditional tests, with applications. Biometrika, 94(4):992–998, December 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/4/992;
 http://www.jstor.org/stable/20441432.

Li:2008:SML

[LPL08] Yi Li, Ross L. Prentice, and Xihong Lin. Semiparametric maximum likelihood estimation in normal transformation models for bivariate survival data. Biometrika, 95(4):947–960, December 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/947;
 http://www.jstor.org/stable/20441517.

Lund:2000:MPP

[LR00] Jens Lund and Mats Rudemo. Models for point processes observed with noise. Biometrika, 87(2):235–249, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/235;
 http://www.jstor.org/stable/2673461.

Liechty:2001:MCM

[LR01] John C. Liechty and Gareth O. Roberts. Markov chain Monte Carlo methods for switching diffusion models. Biometrika, 88(2):299–315, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673481.

Li:2008:APS

[LR08] Yingxing Li and David Ruppert. On the asymptotics of penalized splines. Biometrika, 95(2):415–436, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/415;
 http://www.jstor.org/stable/20441473.

Lohr:2009:JEM

[LR09] Sharon L. Lohr and J. N. K. Rao. Jackknife estimation of mean squared error of small area predictors in nonlinear mixed models. Biometrika, 96(2):457–468, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/457;
 http://www.jstor.org/stable/27798836.

Liu:2000:GGS

[LS00] Jun S. Liu and Chiara Sabatti. Generalised Gibbs sampler and multigrid Monte Carlo for Bayesian computation. Biometrika, 87(2):353–369, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/353;   http://
www.jstor.org/stable/2673469.

Lai:2003:HEN

[LS03a] Tze Leung Lai and Mei-Chiung Shih. A hybrid estimator in nonlinear and generalised linear mixed effects models. Biometrika, 90(4):859–879, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/859;   http://
www.jstor.org/stable/30042093.

Lai:2003:NEN

[LS03b] Tze Leung Lai and Mei-Chiung Shih. Nonparametric estimation in nonlinear mixed effects models. Biometrika, 90(1):1–13, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/1/1;   http://www.jstor.org/stable/30042015.

Lai:2004:PSS

[LS04] Tze Leung Lai and Mei-Chiung Shih. Power, sample size and adaptation considerations in the design of group sequential clinical trials. Biometrika, 91(3):507–528, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/20441120.

Lai:2009:TCI

[LSS09] Tze Leung Lai, Mei-Chiung Shih, and Zheng Su. Tests and confidence intervals for secondary endpoints in sequential clinical trials. Biometrika, 96(4):903–915, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/903;
 http://www.jstor.org/stable/27798875.

Lee:2006:FBR

[LSW06] A. J. Lee, A. J. Scott, and C. J. Wild. Fitting binary regression models with case-augmented samples. Biometrika, 93(2):385–397, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/2/385;   http://www.jstor.org/stable/20441289.

Lindqvist:2005:MCC

[LT05] Bo Henry Lindqvist and Gunnar Taraldsen. Monte Carlo conditioning on a sufficient statistic. Biometrika, 92(2):451–464, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/2/451;   http://www.jstor.org/stable/20441199.

Lu:2006:STM

[LT06] Wenbin Lu and Anastasios A. Tsiatis. Semiparametric transformation models for the case-cohort study. Biometrika, 93(1):207–214, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/1/207;   http://www.jstor.org/stable/20441271.

Lu:2008:IEL

[LT08] Xiaomin Lu and Anastasios A. Tsiatis. Improving the efficiency of the log-rank test using auxiliary covariates. Biometrika, 95(3):679– 694, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/3/679;   http://www.jstor.org/stable/20441494.

Luo:2009:NER

[LT09] Xiaodong Luo and Wei Yann Tsai. Nonparametric estimation for right-censored length-biased data: a pseudo-partial likelihood approach. Biometrika, 96(4):873–886, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/873;
 http://www.jstor.org/stable/27798873.

Lindqvist:2003:CCA

[LTLE03] Bo Henry Lindqvist, Gunnar Taraldsen, Magnar Lillegård, and Steinar Engen. A counterexample to a claim about stochastic simulations. Biometrika, 90(2):489–490, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/489;
 http://www.jstor.org/stable/30042058.

Liang:2008:APL

[LTR^{+}08] Hua Liang, Sally W. Thurston, David Ruppert, Tatiyana Apanasovich, and Russ Hauser. Additive partial linear models with measurement errors. Biometrika, 95(3):667–678, September 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/667;
 http://www.jstor.org/stable/20441493.

Luo:2009:PPL

[LTX09] Xiaodong Luo, Wei Yann Tsai, and Qiang Xu. Pseudo-partial likelihood estimators for the Cox regression model with missing covariates. Biometrika, 96(3):617–633, September 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/617;
 http://www.jstor.org/stable/27798852.

Lue:2004:PHD

[Lue04] Heng-Hui Lue. Principal Hessian directions for regression with measurement error. Biometrika, 91(2):409–423, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/409;
 http://www.jstor.org/stable/20441110.

Liang:2007:PLM

[LWC07] Hua Liang, Suojin Wang, and Raymond J. Carroll. Partially linear models with missing response variables and errorprone covariates. Biometrika, 94(1):185–198, March 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/1/185;
 http://www.jstor.org/stable/20441362.

Lin:2004:EKS

[LWWC04] Xihong Lin, Naisyin Wang, Alan H. Welsh, and Raymond J. Carroll. Equivalent kernels of smoothing splines in nonparametric regression for clustered/longitudinal data. Biometrika, 91(1):177– 193, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/91/1/177;   http://www.jstor.org/stable/20441087.

Lin:2006:NKP

[LWY06] Chien-Tai Lin, Jyh-Shyang Wu, and Chia-Hung Yen. A note on kernel polygons. Biometrika, 93(1):228–234, March 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/1/228;
 http://www.jstor.org/stable/20441274.

Liang:2008:NCA

[LWZ08] Hua Liang, Hulin Wu, and Guohua Zou. A note on conditional AIC for linear mixed-effects models. Biometrika, 95(3):773–778, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/3/773;   http://www.jstor.org/stable/20441501.

Lam:2005:SRC

[LX05] K. F. Lam and Hongqi Xue. A semiparametric regression cure model with current status data. Biometrika, 92(3):573–586, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/3/573;   http://www.jstor.org/stable/20441215.

Lee:2003:PWB

[LY03] Stephen M. S. Lee and G. Alastair Young. Prepivoting by weighted bootstrap iteration. Biometrika, 90(2):393–410, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/393;
 http://www.jstor.org/stable/30042048.

Lu:2004:STC

[LY04] Wenbin Lu and Zhiliang Ying. On semiparametric transformation cure models. Biometrika, 91(2):331–343, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/331;
 http://www.jstor.org/stable/20441105.

Li:2009:ACG

[LY09a] Hui Li and Guosheng Yin. Amendments and corrections: “Generalized method of moments estimation for linear regression with clustered failure time data” [MR2507144]. Biometrika, 96(4):1024, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/1024. See [LY09b].

Li:2009:GMM

[LY09b] Hui Li and Guosheng Yin. Generalized method of moments estimation for linear regression with clustered failure time data. Biometrika, 96(2):293–306, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/2/293;   http://
www.jstor.org/stable/27798825. See amendments and correc-

tions [LY09a].

Liang:2008:EFD

[LZ08] Faming Liang and Jian Zhang. Estimating the false discovery rate using the stochastic approximation algorithm. Biometrika, 95(4):961–977, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/95/4/961;   http://www.jstor.
org/stable/20441518.

Lu:2007:EMF

[LZH07] Minggen Lu, Ying Zhang, and Jian Huang. Estimation of the mean function with panel count data using monotone polynomial splines. Biometrika, 94(3):705–718, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/705;
 http://www.jstor.org/stable/20441406.

Martin:2000:EGM

[Mar00] R. J. Martin. Exact Gaussian maximum likelihood and simulation for regularly-spaced observations with Gaussian correlations. Biometrika, 87(3):727–730, September 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/3/727;
 http://www.jstor.org/stable/2673646.

Marschner:2001:SVE

[Mar01] Ian C. Marschner. On stochastic versions of the EM algorithm. Biometrika, 88(1):281–286, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673685.

Marriott:2002:LGM

[Mar02] Paul Marriott. On the local geometry of mixture models. Biometrika, 89(1):77–93, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/89/1/77;   http://
www.jstor.org/stable/4140560.

Marriott:2003:GME

[Mar03] Paul Marriott. On the geometry of measurement error models. Biometrika, 90(3):567–576, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/3/567;   http://
www.jstor.org/stable/30042066.

Matsuda:2006:TSG

[Mat06] Yasumasa Matsuda. A test statistic for graphical modelling of multivariate time series. Biometrika, 93(2):399–409, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/399;
 http://www.jstor.org/stable/20441290.

Meinshausen:2005:LBN

[MB05] Nicolai Meinshausen and Peter Bühlmann. Lower bounds for the number of false null hypotheses for multiple testing of associations under general dependence structures. Biometrika, 92(4):893–907, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/4/893;   http://www.jstor.org/stable/20441243.

Malzahn:2000:NEH

[MBH00] Uwe Malzahn, Dankmar Böhning, and Heinz Holling. Nonparametric estimation of heterogeneity variance for the standardised difference used in meta-analysis. Biometrika, 87(3):619–632, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/3/619;   http://www.jstor.org/stable/2673634.

Mardia:2000:SAB

[MBM00] Kanti V. Mardia, Fred L. Bookstein, and Ian J. Moreton. Statistical assessment of bilateral symmetry of shapes. Biometrika, 87(2):285–300, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/285;   http://
www.jstor.org/stable/2673464. See amendments and correction

[MBM05].

Mardia:2005:ACS

[MBM05] K. V. Mardia, F. L. Bookstein, and I. J. Moreton. Amendments and corrections: “Statistical assessment of bilateral symmetry of shapes” [Biometrika 87 (2000), no. 2, 285–300; MR1782479]. Biometrika, 92(1):249–250, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/92/1/249.2;   http:/
/www.jstor.org/stable/20441182. See [MBM00].

Molchanov:2000:STE

[MC00] I. S. Molchanov and S. N. Chiu. Smoothing techniques and estimation methods for nonstationary Boolean models with applications to coverage processes. Biometrika, 87(2):265–283, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/265;
 http://www.jstor.org/stable/2673463.

Ma:2006:ESE

[MCW06] Yanyuan Ma, Jeng-Min Chiou, and Naisyin Wang. Efficient semiparametric estimator for heteroscedastic partially linear models. Biometrika, 93(1):75–84, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/93/1/75;   http://
www.jstor.org/stable/20441261.

Mukerjee:2002:OME

[MDC02] Rahul Mukerjee, Aloke Dey, and Kashinath Chatterjee. Optimal main effect plans with non-orthogonal blocking. Biometrika, 89(1): 225–229, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/1/225;   http://www.jstor.org/stable/4140569.

Meinshausen:2008:HTV

[Mei08] Nicolai Meinshausen. Hierarchical testing of variable importance. Biometrika, 95(2):265–278, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/265;   http://
www.jstor.org/stable/20441463.

Meyer:2003:TLV

[Mey03] Mary C. Meyer. A test for linear versus convex regression function using shape-restricted regression. Biometrika, 90(1):223–232, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/1/223;   http://www.jstor.org/stable/30042031.

Mandel:2009:NEP

[MF09] M. Mandel and R. Fluss. Nonparametric estimation of the probability of illness in the illness-death model under cross-sectional sampling. Biometrika, 96(4):861–872, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/861;
 http://www.jstor.org/stable/27798872.

Ma:2007:CLL

[MH07] Yanyuan Ma and Jeffrey D. Hart. Constrained local likelihood estimators for semiparametric skew-normal distributions. Biometrika, 94(1):119–134, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/1/119;   http://
www.jstor.org/stable/20441358.

Ma:2003:REC

[MKB03] Renjun Ma, Daniel Krewski, and Richard T. Burnett. Random effects Cox models: A Poisson modelling approach. Biometrika, 90(1):157–169, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/1/157;   http://
www.jstor.org/stable/30042026.

Mardia:2009:MLE

[MKHT09] Kanti V. Mardia, John T. Kent, Gareth Hughes, and Charles C. Taylor. Maximum likelihood estimation using composite likelihoods for closed exponential families. Biometrika, 96(4):975–982, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/975;   http://www.jstor.org/stable/27798881.

Mao:2002:PMC

[ML02] Chang Xuan Mao and Bruce G. Lindsay. A Poisson model for the coverage problem with a genomic application. Biometrika, 89(3): 669–681, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/3/669;   http://www.jstor.org/stable/4140610.

Mammen:2003:GSM

[MN03] Enno Mammen and Jens Perch Nielsen. Generalised structured models. Biometrika, 90(3):551–566, September 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/551;
 http://www.jstor.org/stable/30042065.

Mammen:2007:GAP

[MN07] Enno Mammen and Jens Perch Nielsen. A general approach to the predictability issue in survival analysis with applications. Biometrika, 94(4):873–892, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/4/873;   http://
www.jstor.org/stable/20441423.

Marshall:1997:NMA

[MO97] Albert W. Marshall and Ingram Olkin. A new method for adding a parameter to a family of distributions with application to the exponential and Weibull families. Biometrika, 84(3):641–652, September 1997. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2337585. See

correction [MO05].

Marshall:2005:ACN

[MO05] Albert W. Marshall and Ingram Olkin. Amendments and corrections: “A new method for adding a parameter to a family of distributions with application to the exponential and Weibull families” [Biometrika 84 (1997), no. 3, 641–652; MR1603936]. Biometrika, 92(2):505, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/2/505.1;   http://www.jstor.org/stable/

20441205. See [MO97].

Moodie:2009:NVD

[Moo09] E. E. M. Moodie. A note on the variance of doubly-robust Gestimators. Biometrika, 96(4):998–1004, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/998;
 http://www.jstor.org/stable/27798884.

Muller:2003:MDE

[MP03] Werner G. Müller and Andrej Pázman. Measures for designs in experiments with correlated errors. Biometrika, 90(2):423–434, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/2/423;   http://www.jstor.org/stable/30042050.

Moller:2006:EMC

[MPRB06] J. Møller, A. N. Pettitt, R. Reeves, and K. K. Berthelsen. An efficient Markov chain Monte Carlo method for distributions with intractable normalising constants. Biometrika, 93(2):451–458, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/2/451;   http://www.jstor.org/stable/20441294.

Mukerjee:2001:SOP

[MR01] Rahul Mukerjee and N. Reid. Second-order probability matching priors for a parametric function with application to Bayesian tolerance limits. Biometrika, 88(2):587–592, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673503.

Martinussen:2002:EEA

[MS02a] Torben Martinussen and Thomas H. Scheike. Efficient estimation in additive hazards regression with current status data. Biometrika, 89(3):649–658, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/649;   http://
www.jstor.org/stable/4140608.

Martinussen:2002:FAM

[MS02b] Torben Martinussen and Thomas H. Scheike. A flexible additive multiplicative hazard model. Biometrika, 89(2):283–298, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/2/283;   http://www.jstor.org/stable/4140577.

Martinussen:2007:AAH

[MS07] Torben Martinussen and Thomas H. Scheike. Aalen additive hazards change-point model. Biometrika, 94(4):861–872, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/4/861;   http://www.jstor.org/stable/20441422.

Mukherjee:2007:AMT

[MSV07] K. Mukherjee, R. H. Shumway, and K. L. Verosub. On the alignment of multiple time series fragments. Biometrika, 94(2):347–358, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/2/347;   http://www.jstor.org/stable/20441376.

Muller:2001:EVC

[MU01] Christine H. Müller and Steffen Uhlig. Estimation of variance components with high breakdown point and high efficiency. Biometrika, 88(2):353–366, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673484.

Nguyen:2003:PCP

[NCGD03] Truc T. Nguyen, John T. Chen, Arjun K. Gupta, and Khoan T. Dinh. A proof of the conjecture on positive skewness of generalised inverse Gaussian distributions. Biometrika, 90(1):245–250, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/1/245;   http://www.jstor.org/stable/30042034.

Ni:2005:NSS

[NCT05] Liqiang Ni, R. Dennis Cook, and Chih-Ling Tsai. A note on shrinkage sliced inverse regression. Biometrika, 92(1):242–247, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/1/242;   http://www.jstor.org/stable/20441180.

Nott:2002:ENS

[ND02] David J. Nott and William T. M. Dunsmuir. Estimation of nonstationary spatial covariance structure. Biometrika, 89(4):819–829, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/4/819;   http://www.jstor.org/stable/4140539.

Nobile:2000:BAF

[NG00] A. Nobile and P. J. Green. Bayesian analysis of factorial experiments by mixture modelling. Biometrika, 87(1):15–35, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/1/15;   http://www.jstor.org/stable/2673558.

Naveau:2009:MPD

[NGCD09] Philippe Naveau, Armelle Guillou, Daniel Cooley, and Jean Diebolt. Modelling pairwise dependence of maxima in space. Biometrika, 96(1):1–17, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/96/1/1;   http://www.jstor.
org/stable/27798798.

Nielsen:2001:CTR

[Nie01] B. Nielsen. Conditional test for rank in bivariate canonical correlation analysis. Biometrika, 88(3):874–880, September 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673454.

Nott:2005:ASB

[NK05] David J. Nott and Robert Kohn. Adaptive sampling for Bayesian variable selection. Biometrika, 92(4):747–763, December 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/4/747;
 http://www.jstor.org/stable/20441233.

Nordman:2009:TEL

[Nor09] Daniel J. Nordman. Tapered empirical likelihood for time series data in time and frequency domains. Biometrika, 96(1):119–132, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/1/119;   http://www.jstor.org/stable/27798806.

Neumann:2008:SCB

[NP08] Michael H. Neumann and Efstathios Paparoditis. Simultaneous confidence bands in spectral density estimation. Biometrika, 95(2): 381–397, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/2/381;   http://www.jstor.org/stable/20441471.

Neuhaus:2002:ARF

[NSW02] J. Neuhaus, A. J. Scott, and C. J. Wild. The analysis of retrospective family studies. Biometrika, 89(1):23–37, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/1/23;   http://www.jstor.org/stable/4140556.

Naik:2001:SIM

[NT01] Prasad A. Naik and Chih-Ling Tsai. Single-index model selections. Biometrika, 88(3):821–832, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673449.

Nedyalkova:2008:OSE

[NT08] Desislava Nedyalkova and Yves Tillé. Optimal sampling and estimation strategies under the linear model. Biometrika, 95(3):521– 537, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/3/521;   http://www.jstor.org/stable/20441484.

Nan:2006:CLR

[NYK06] Bin Nan, Menggang Yu, and John D. Kalbfleisch. Censored linear regression for case-cohort studies. Biometrika, 93(4):747–762, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/4/747;   http://www.jstor.org/stable/20441325.

Oakes:2001:BCS

[Oak01] David Oakes. Biometrika Centenary: Survival analysis. Biometrika, 88(1):99–142, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673676.

Oakes:2008:CKT

[Oak08] David Oakes. On consistency of Kendall’s Tau under censoring. Biometrika, 95(4):997–1001, December 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/997;
 http://www.jstor.org/stable/20441522.

Oller:2007:ICI

[OGC07] Ramon Oller, Guadalupe Gómez, and M. Luz Calle. Interval censoring: Identifiability and the constant-sum property. Biometrika, 94(1):61–70, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/94/1/61;   http://
www.jstor.org/stable/20441354.

Olhede:2004:LSP

[OMS04] Sofia C. Olhede, Emma J. McCoy, and David A. Stephens. Largesample properties of the periodogram estimator of seasonally persistent processes. Biometrika, 91(3):613–628, September 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441127.

Oh:2001:PBT

[ONL01] Hee-Seok Oh, Philippe Naveau, and Geunghee Lee. Polynomial boundary treatment for wavelet regression. Biometrika, 88(1):291– 298, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2673687.

Oh:2007:RPD

[ONL07] Hee-Seok Oh, Douglas W. Nychka, and Thomas C. M. Lee. The role of pseudo data for robust smoothing with application to wavelet regression. Biometrika, 94(4):893–904, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/4/893;
 http://www.jstor.org/stable/20441424.

Oakley:2002:BIU

[OO02] Jeremy Oakley and Anthony O’Hagan. Bayesian inference for the uncertainty distribution of computer model outputs. Biometrika, 89(4):769–784, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/769;   http://
www.jstor.org/stable/4140536.

Oakley:2007:UPE

[OO07] Jeremy E. Oakley and Anthony O’Hagan. Uncertainty in prior elicitations: A nonparametric approach. Biometrika, 94(2):427–441, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/2/427;   http://www.jstor.org/stable/20441382.

OQuigley:2003:KTG

[O’Q03] John O’Quigley. Khmaladze-type graphical evaluation of the proportional hazards assumption. Biometrika, 90(3):577–584, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/3/577;   http://www.jstor.org/stable/30042067.

Oakes:2000:RBC

[OR00] David Oakes and John Ritz. Regression in a bivariate copula model. Biometrika, 87(2):345–352, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/345;   http://
www.jstor.org/stable/2673468.

Ombao:2001:SGC

[ORSV01] Hernando C. Ombao, Jonathan A. Raz, Robert L. Strawderman, and Rainer Von Sachs. A simple generalised crossvalidation method of span selection for periodogram smoothing. Biometrika, 88(4):1186–1192, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/88/4/1186;   http://
www.jstor.org/stable/2673711.

Oakes:2003:CMG

[OW03] David Oakes and Antai Wang. Copula model generated by Dabrowska’s association measure. Biometrika, 90(2):478–481, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/2/478;   http://www.jstor.org/stable/30042056.

Olhede:2004:AWT

[OW04] Sofia C. Olhede and Andrew T. Walden. ‘Analytic’ wavelet thresholding. Biometrika, 91(4):955–973, December 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/4/955;
 http://www.jstor.org/stable/20441155.

Oman:2001:MGC

[OZ01] Samuel D. Oman and David M. Zucker. Modelling and generating correlated binary variables. Biometrika, 88(1):287–290, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2673686.

Paddock:2002:BNM

[Pad02] Susan M. Paddock. Bayesian nonparametric multiple imputation of partially observed data with ignorable nonresponse. Biometrika, 89(3):529–538, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/529;   http://
www.jstor.org/stable/4140599.

Pal:2009:NCS

[Pal09] Soumik Pal. A note on a conjectured sharpness principle for probabilistic forecasting with calibration. Biometrika, 96(4):1019–1023, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/1019;   http://www.jstor.org/stable/27798887.

See correction [Pal10].

Pal:2010:ACC

[Pal10] Soumik Pal. Amendments and corrections: ’On a conjectured sharpness principle for probabilistic forecasting with calibration’ [errata to MR2767288]. Biometrika, 97(4):1013, December 2010. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/29777157. See [Pal09].

Pan:2001:RVE

[Pan01] Wei Pan. On the robust variance estimator in generalised estimating equations. Biometrika, 88(3):901–906, September 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673458.

Paige:2001:BIN

[PB01] R. L. Paige and R. W. Butler. Bayesian inference in neural networks. Biometrika, 88(3):623–641, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673435.

Perez:2002:EPP

[PB02] José M. Pérez and James O. Berger. Expected-posterior prior distributions for model selection. Biometrika, 89(3):491–511, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/3/491;   http://www.jstor.org/stable/4140597.

Pierce:2006:ERS

[PB06] Donald A. Pierce and Ruggero Bellio. Effects of the reference set on frequentist inferences. Biometrika, 93(2):425–438, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/425;
 http://www.jstor.org/stable/20441292.

Pitt:2006:EBI

[PCK06] Michael Pitt, David Chan, and Robert Kohn. Efficient Bayesian inference for Gaussian copula regression models. Biometrika, 93(3):537–554, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/3/537;   http://www.jstor.
org/stable/20441306.

Pfeffermann:2006:MLM

[PDD06] Danny Pfeffermann, Fernando Antonio Da Silva Moura, and Pedro Luis Do Nascimento Silva. Multi-level modelling under informative sampling. Biometrika, 93(4):943–959, December 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/4/943;
 http://www.jstor.org/stable/20441337.

Peddada:2005:EOR

[PDT05] Shyamal D. Peddada, David B. Dunson, and Xiaofeng Tan. Estimation of order-restricted means from correlated data. Biometrika, 92(3):703–715, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/3/703;   http://
www.jstor.org/stable/20441224.

Pena:2009:DRT

[Peñ09] Daniel Peña. Dimension reduction in time series and the dynamic factor model. Biometrika, 96(2):494–496, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/494;
 http://www.jstor.org/stable/27798840.

Peng:2006:NEL

[PF06] L. Peng and J. P. Fine. Nonparametric estimation with lefttruncated semicompeting risks data. Biometrika, 93(2):367–383, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/2/367;   http://www.jstor.org/stable/20441288.

Peng:2007:NQI

[PF07] L. Peng and J. P. Fine. Nonparametric quantile inference with competing–risks data. Biometrika, 94(3):735–744, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/735;
 http://www.jstor.org/stable/20441408.

Pfeiffer:2001:ICA

[PGP01] Ruth M. Pfeiffer, Mitchell H. Gail, and David Pee. Inference for covariates that accounts for ascertainment and random genetic effects in family studies. Biometrika, 88(4):933–948, December 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/88/4/933;
 http://www.jstor.org/stable/2673693.

Peng:2007:SAT

[PH07] Limin Peng and Yijian Huang. Survival analysis with temporal covariate effects. Biometrika, 94(3):719–733, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/719;
 http://www.jstor.org/stable/20441407.

Parner:2001:MPH

[PK01] Erik T. Parner and Niels Keiding. Misspecified proportional hazard models and confirmatory analysis of survival data. Biometrika, 88(2):459–468, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673493.

Park:2004:ALD

[PK04a] Eunsik Park and Yongdai Kim. Analysis of longitudinal data in case-control studies. Biometrika, 91(2):321–330, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/321;
 http://www.jstor.org/stable/20441104.

Pierce:2004:ACE

[PK04b] Donald A. Pierce and Albrecht M. Kellerer. Adjusting for covariate errors with nonparametric assessment of the true covariate distribution. Biometrika, 91(4):863–876, December 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/4/863;
 http://www.jstor.org/stable/20441149.

Pilla:2001:AEM

[PL01] Ramani S. Pilla and Bruce G. Lindsay. Alternative EM methods for nonparametric finite mixture models. Biometrika, 88(2):535–550, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673498.

Parzen:2005:NRB

[PLF05] Michael Parzen, Stuart R. Lipsitz, and Garrett M. Fitzmaurice. A note on reducing the bias of the approximate Bayesian bootstrap imputation variance estimator. Biometrika, 92(4):971–974, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/4/971;   http://www.jstor.org/stable/20441250.

Pan:2003:MMC

[PM03] Jianxin Pan and Gilbert Mackenzie. On modelling mean-covariance structures in longitudinal studies. Biometrika, 90(1):239–244, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/1/239;   http://www.jstor.org/stable/30042033.

Pourahmadi:2000:MLE

[Pou00] Mohsen Pourahmadi. Maximum likelihood estimation of generalised linear models for multivariate normal covariance matrix. Biometrika, 87(2):425–435, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/425;   http://
www.jstor.org/stable/2673473.

Pourahmadi:2007:CDE

[Pou07] Mohsen Pourahmadi. Cholesky decompositions and estimation of a covariance matrix: Orthogonality of variance–correlation parameters. Biometrika, 94(4):1006–1013, December 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/4/1006;   http://www.jstor.org/stable/20441434.

Paparoditis:2001:TBB

[PP01] Efstathios Paparoditis and Dimitris N. Politis. Tapered block bootstrap. Biometrika, 88(4):1105–1119, December 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
88/4/1105;   http://www.jstor.org/stable/2673704.

Preisser:2008:NDD

[PQP08] John S. Preisser, Bahjat F. Qaqish, and Jamie Perin. A note on deletion diagnostics for estimating equations. Biometrika, 95(2): 509–513, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/2/509;   http://www.jstor.org/stable/20441479.

Papaspiliopoulos:2008:RMC

[PR08] Omiros Papaspiliopoulos and Gareth O. Roberts. Retrospective Markov chain Monte Carlo methods for Dirichlet process hierarchical models. Biometrika, 95(1):169–186, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/169;
 http://www.jstor.org/stable/20441450.

Prendergast:2007:IIF

[Pre07] Luke A. Prendergast. Implications of influence function analysis for sliced inverse regression and sliced average variance estimation. Biometrika, 94(3):585–601, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/585;
 http://www.jstor.org/stable/20441398.

Pintore:2006:SAS

[PSH06] Alexandre Pintore, Paul Speckman, and Chris C. Holmes. Spatially adaptive smoothing splines. Biometrika, 93(1):113–125, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/1/113;   http://www.jstor.org/stable/20441264.

Park:2003:ESS

[PW03] Yuhyun Park and L. J. Wei. Estimating subject-specific survival functions under the accelerated failure time model. Biometrika, 90(3):717–723, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/3/717;   http://
www.jstor.org/stable/30042077.

Posch:2008:NRV

[PWB08] Martin Posch, Gernot Wassmer, and Werner Brannath. A note on repeated p-values for group sequential designs. Biometrika, 95(1): 253–256, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/1/253;   http://www.jstor.org/stable/20441457.

Peng:2003:LAD

[PY03] Liang Peng and Qiwei Yao. Least absolute deviations estimation for ARCH and GARCH models. Biometrika, 90(4):967–975, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/4/967;   http://www.jstor.org/stable/30042101.

Pan:2008:MMT

[PY08] Jiazhu Pan and Qiwei Yao. Modelling multiple time series via common factors. Biometrika, 95(2):365–379, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/365;
 http://www.jstor.org/stable/20441470.

Qaqish:2003:FMB

[Qaq03] Bahjat F. Qaqish. A family of multivariate binary distributions for simulating correlated binary variables with specified marginal means and correlations. Biometrika, 90(2):455–463, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/455;
 http://www.jstor.org/stable/30042053.

Qaqish:2006:MLM

[QI06] Bahjat F. Qaqish and Anastasia Ivanova. Multivariate logistic models. Biometrika, 93(4):1011–1017, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/1011;   http://www.jstor.org/stable/20441343.

Qian:2009:NLH

[Qia09] Peter Z. G. Qian. Nested Latin hypercube designs. Biometrika, 96(4):957–970, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/96/4/957;   http://www.jstor.
org/stable/27798879.

Qin:2000:CPE

[Qin00] Jing Qin. Combining parametric and empirical likelihoods. Biometrika, 87(2):484–490, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/484;   http://
www.jstor.org/stable/2673478.

Qu:2000:IGE

[QLL00] Annie Qu, Bruce G. Lindsay, and Bing Li. Improving generalised estimating equations using quadratic inference functions.

Biometrika, 87(4):823–836, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/4/823;   http://
www.jstor.org/stable/2673612.

Qu:2008:MDT

[QLL08] Annie Qu, J. Jack Lee, and Bruce G. Lindsay. Model diagnostic tests for selecting informative correlation structure in correlated data. Biometrika, 95(4):891–905, December 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/891;
 http://www.jstor.org/stable/20441513.

Qu:2002:TIM

[QS02] Annie Qu and Peter X.-K. Song. Testing ignorable missingness in estimating equation approaches for longitudinal data. Biometrika, 89(4):841–850, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/841;   http://
www.jstor.org/stable/4140541.

Qu:2004:ARG

[QS04] Annie Qu and Peter X.-K. Song. Assessing robustness of generalised estimating equations and quadratic inference functions. Biometrika, 91(2):447–459, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/447;   http://
www.jstor.org/stable/20441112.

Qian:2009:SSF

[QW09] Peter Z. G. Qian and C. F. Jeff Wu. Sliced space-filling designs. Biometrika, 96(4):945–956, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/945;
 http://www.jstor.org/stable/27798878.

Qin:2003:ULR

[QZ03] Jing Qin and Biao Zhang. Using logistic regression procedures for estimating receiver operating characteristic curves. Biometrika, 90(3):585–596, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/3/585;   http://
www.jstor.org/stable/30042068. See correction [QZ05a].

Qin:2005:ACU

[QZ05a] Jing Qin and Biao Zhang. Amendments and corrections: “Using logistic regression procedures for estimating receiver operating characteristic curves” [Biometrika 90 (2003), no. 3, 585– 596; MR2006837]. Biometrika, 92(1):250, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/250;
 http://www.jstor.org/stable/20441183. See [QZ03].

Qin:2005:MLC

[QZ05b] Jing Qin and Biao Zhang. Marginal likelihood, conditional likelihood and empirical likelihood: Connections and applications. Biometrika, 92(2):251–270, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/251;   http://
www.jstor.org/stable/20441187.

Rao:2004:MRM

[Rao04] Suhasini Subba Rao. On multiple regression models with nonstationary correlated errors. Biometrika, 91(3):645–659, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/20441129.

Rodriguez:2009:BNF

[RDG09] Abel Rodríguez, David B. Dunson, and Alan E. Gelfand. Bayesian nonparametric functional data analysis through density estimation. Biometrika, 96(1):149–162, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/1/149;   http://
www.jstor.org/stable/27798808.

Reiter:2007:SSD

[Rei07] Jerome P. Reiter. Small-sample degrees of freedom for multicomponent significance tests with multiple imputation for missing data. Biometrika, 94(2):502–508, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/2/502;   http://
www.jstor.org/stable/20441388.

Reiter:2008:MIW

[Rei08] Jerome P. Reiter. Multiple imputation when records used for imputation are not used or disseminated for analysis.

Biometrika, 95(4):933–946, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/933;   http://
www.jstor.org/stable/20441516.

Renaud:2002:DPP

[Ren02] Olivier Renaud. The discrimination power of projection pursuit with different density estimators. Biometrika, 89(1):129–143, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/1/129;   http://www.jstor.org/stable/4140563.

Raghavarao:2003:SCB

[RF03] Damaraju Raghavarao and Walter T. Federer. Sufficient conditions for balanced incomplete block designs to be minimal fractional combinatorial treatment designs. Biometrika, 90(2):465–470, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/2/465;   http://www.jstor.org/stable/30042054.

Rivest:2001:DMS

[Riv01] Louis-Paul Rivest. A directional model for the statistical analysis of movement in three dimensions. Biometrika, 88(3):779–791, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2673446.

Rosen:2000:MMM

[RJT00] Ori Rosen, Wenxin Jiang, and Martin A. Tanner. Mixtures of marginal models. Biometrika, 87(2):391–404, June 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/2/391;
 http://www.jstor.org/stable/2673471.

Rodland:2006:SPV

[Rød06] Einar Andreas Rødland. Simes’ procedure is ‘valid on average’. Biometrika, 93(3):742–746, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/742;
 http://www.jstor.org/stable/20441320.

Rosenbaum:2001:EAT

[Ros01] Paul R. Rosenbaum. Effects attributable to treatment: Inference in experiments and observational studies with a discrete pivot. Biometrika, 88(1):219–231, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673680.

Rosenbaum:2004:DSO

[Ros04] Paul R. Rosenbaum. Design sensitivity in observational studies. Biometrika, 91(1):153–164, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/1/153;   http://
www.jstor.org/stable/20441085.

Rosenbaum:2006:DEG

[Ros06] Paul R. Rosenbaum. Differential effects and generic biases in observational studies. Biometrika, 93(3):573–586, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/573;
 http://www.jstor.org/stable/20441308.

Rosenbaum:2008:THO

[Ros08] Paul R. Rosenbaum. Testing hypotheses in order. Biometrika, 95(1):248–252, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/95/1/248;   http://www.jstor.
org/stable/20441456.

Roverato:2000:CDH

[Rov00] Alberto Roverato. Cholesky decomposition of a hyper inverse Wishart matrix. Biometrika, 87(1):99–112, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/1/99;   http://www.jstor.org/stable/2673564.

Reeves:2004:ERG

[RP04] R. Reeves and A. N. Pettitt. Efficient recursions for general factorisable models. Biometrika, 91(3):751–757, September 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441137.

Robins:2004:ETE

[RR04] James Robins and Andrea Rotnitzky. Estimation of treatment effects in randomised trials with non-compliance and a dichotomous outcome using structural mean models. Biometrika, 91(4):763–783, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/4/763;   http://www.jstor.org/stable/20441143.

Roberts:2001:IPO

[RS01] G. O. Roberts and O. Stramer. On inference for partially observed nonlinear diffusion models using the Metropolis-Hastings algorithm. Biometrika, 88(3):603–621, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673434.

Ritz:2005:LRT

[RS05] Christian Ritz and Ib M. Skovgaard. Likelihood ratio tests in curved exponential families with nuisance parameters present only under the alternative. Biometrika, 92(3):507–517, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/3/507;
 http://www.jstor.org/stable/20441210.

Rosen:2007:AEM

[RS07] Ori Rosen and David S. Stoffer. Automatic estimation of multivariate spectra via smoothing splines. Biometrika, 94(2):335–345, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/2/335;   http://www.jstor.org/stable/20441375.

Rathouz:2002:SIM

[RSC02] Paul J. Rathouz, Glen A. Satten, and Raymond J. Carroll. Semiparametric inference in matched case-control studies with missing covariate data. Biometrika, 89(4):905–916, December 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/4/905;
 http://www.jstor.org/stable/4140546.

Rathbun:2007:MEP

[RSG07] Stephen L. Rathbun, Saul Shiffman, and Chad J. Gwaltney. Modelling the effects of partially observed covariates on Poisson process intensity. Biometrika, 94(1):153–165, March 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/1/153;
 http://www.jstor.org/stable/20441360.

Robins:2003:UCC

[RSSW03] James M. Robins, Richard Scheines, Peter Spirtes, and Larry Wasserman. Uniform consistency in causal inference. Biometrika, 90(3):491–515, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/3/491;   http://www.jstor.
org/stable/30042062.

Rizopoulos:2008:SPM

[RVM08] Dimitris Rizopoulos, Geert Verbeke, and Geert Molenberghs. Shared parameter models under random effects misspecification. Biometrika, 95(1):63–74, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/95/1/63;   http://
www.jstor.org/stable/20441443.

Robins:2000:IIE

[RW00] James M. Robins and Naisyin Wang. Inference for imputation estimators. Biometrika, 87(1):113–124, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/113;
 http://www.jstor.org/stable/2673565.

Reale:2002:SPC

[RW02] Marco Reale and Granville Tunnicliffe Wilson. The sampling properties of conditional independence graphs for structural vector autoregressions. Biometrika, 89(2):457–461, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/457;
 http://www.jstor.org/stable/4140590.

Sorensen:2000:CRA

[SA00] Per Sørensen and Per Kragh Andersen. Competing risks analysis of the case-cohort design. Biometrika, 87(1):49–59, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/1/49;   http://www.jstor.org/stable/2673560.

Samworth:2003:NMR

[Sam03] Richard Samworth. A note on methods of restoring consistency to the bootstrap. Biometrika, 90(4):985–990, December 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/985;
 http://www.jstor.org/stable/30042104.

Sartori:2003:MPL

[Sar03] N. Sartori. Modified profile likelihoods in models with stratum nuisance parameters. Biometrika, 90(3):533–549, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/533;
 http://www.jstor.org/stable/30042064.

Sen:2007:PMA

[SB07] Bodhisattva Sen and Moulinath Banerjee. A pseudolikelihood method for analyzing interval censored data. Biometrika, 94(1):71– 86, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/94/1/71;   http://www.jstor.org/stable/20441355.

Samuel-Cahn:1996:SIB

[SC96] Ester Samuel-Cahn. Is the Simes improved Bonferroni procedure conservative? Biometrika, 83(4):928–933, December 1996. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2337297. See correction

[Ano04a].

Schaubel:2004:RMG

[SC04] Douglas E. Schaubel and Jianwen Cai. Regression methods for gap time hazard functions of sequentially ordered multivariate failure time data. Biometrika, 91(2):291–303, June 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/2/291;
 http://www.jstor.org/stable/20441102.

Schott:2003:WCS

[Sch03] James R. Schott. Weighted chi-squared tests for partial common principal component subspaces. Biometrika, 90(2):411–421, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/2/411;   http://www.jstor.org/stable/30042049.

Schennach:2005:BET

[Sch05a] Susanne M. Schennach. Bayesian exponentially tilted empirical likelihood. Biometrika, 92(1):31–46, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/31;
 http://www.jstor.org/stable/20441164.

Schott:2005:TCI

[Sch05b] James R. Schott. Testing for complete independence in high dimensions. Biometrika, 92(4):951–956, December 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/4/951;
 http://www.jstor.org/stable/20441247.

Samia:2007:GTM

[SCS07] Noelle I. Samia, Kung-Sik Chan, and Nils Chr. Stenseth. A generalized threshold mixed model for analyzing nonnormal nonlinear time series, with application to plague in Kazakhstan. Biometrika, 94(1):101–118, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/94/1/101;   http://
www.jstor.org/stable/20441357.

Shao:2007:MTS

[SCW07] Yongwu Shao, R. Dennis Cook, and Sanford Weisberg. Marginal tests with sliced average variance estimation. Biometrika, 94(2): 285–296, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/2/285;   http://www.jstor.org/stable/20441372.

Semadeni:2004:PPI

[SDH04] C. Semadeni, A. C. Davison, and D. V. Hinkley. Posterior probability intervals in Bayesian wavelet estimation. Biometrika, 91(2):497–505, June 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/2/497;   http://
www.jstor.org/stable/20441116.

Severini:2000:LRA

[Sev00] Thomas A. Severini. The likelihood ratio approximation to the conditional distribution of the maximum likelihood estimator in the discrete case. Biometrika, 87(4):939–945, December 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/939;
 http://www.jstor.org/stable/2673621.

Severini:2002:MEF

[Sev02] Thomas A. Severini. Modified estimating functions. Biometrika, 89(2):333–343, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/89/2/333;   http://www.jstor.
org/stable/4140580.

Severini:2004:MLR

[Sev04] Thomas A. Severini. A modified likelihood ratio statistic for some nonregular models. Biometrika, 91(3):603–612, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441126.

Severini:2007:ILF

[Sev07] Thomas A. Severini. Integrated likelihood functions for non-Bayesian inference. Biometrika, 94(3):529–542, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/529;
 http://www.jstor.org/stable/20441394.

Sun:2003:NTP

[SF03] Jianguo Sun and Hong-Bin Fang. A nonparametric test for panel count data. Biometrika, 90(1):199–208, March 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/1/199;
 http://www.jstor.org/stable/30042029.

Shao:2009:CIS

[Sha09] Xiaofeng Shao. Confidence intervals for spectral mean and ratio statistics. Biometrika, 96(1):107–117, March 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/1/107;
 http://www.jstor.org/stable/27798805.

Singh:2002:PMT

[SHD02] Harshinder Singh, Vladimir Hnizdo, and Eugene Demchuk. Probabilistic model for two dependent circular variables. Biometrika, 89(3):719–723, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/89/3/719;   http://www.jstor.
org/stable/4140614.

Shieh:2001:SSC

[Shi01] Gwowen Shieh. Sample size calculations for logistic and Poisson regression models. Biometrika, 88(4):1193–1199, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
88/4/1193;   http://www.jstor.org/stable/2673712.

Shoji:2002:NSE

[Sho02] Isao Shoji. Nonparametric state estimation of diffusion processes. Biometrika, 89(2):451–456, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/451;   http://
www.jstor.org/stable/4140589.

Sinha:2003:BJC

[SIC03] Debajyoti Sinha, Joseph G. Ibrahim, and Ming-Hui Chen. A Bayesian justification of Cox’s partial likelihood. Biometrika, 90(3):629–641, September 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/3/629;   http://www.jstor.
org/stable/30042071.

Siegmund:2004:MSI

[Sie04] David Siegmund. Model selection in irregular problems: Applications to mapping quantitative trait loci. Biometrika, 91(4):785– 800, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/91/4/785;   http://www.jstor.org/stable/20441144.

Sutradhar:2000:AOL

[SK00] Brajendra C. Sutradhar and Milorad Kovacevic. Analysing ordinal longitudinal survey data: Generalised estimating equations approach. Biometrika, 87(4):837–848, December 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/837;
 http://www.jstor.org/stable/2673613.

Shu:2005:AHM

[SK05] Youyi Shu and John P. Klein. Additive hazards Markov regression models illustrated with bone marrow transplant data. Biometrika, 92(2):283–301, June 2005. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/283;   http://
www.jstor.org/stable/20441189.

Shao:2000:CBP

[SKP00] Jun Shao, Jürgen Kübler, and Iris Pigeot. Consistency of the bootstrap procedure in individual bioequivalence. Biometrika, 87(3):573–585, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/87/3/573;   http://www.jstor.
org/stable/2673631.

Stoner:2002:ACD

[SL02] Julie A. Stoner and Brian G. Leroux. Analysis of clustered data: A combined estimating equations approach. Biometrika, 89(3):567– 578, August 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/3/567;   http://www.jstor.org/stable/4140602.

Steinberg:2006:AC

[SL06a] David M. Steinberg and Dennis K. J. Lin. Amendments and corrections: “A Construction Method for Orthogonal Latin Hypercube Designs”. Biometrika, 93(4):1025, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/

93/4/1025.2. See [SL06b].

Steinberg:2006:CMO

[SL06b] David M. Steinberg and Dennis K. J. Lin. A construction method for orthogonal Latin hypercube designs. Biometrika, 93(2):279– 288, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/2/279;   http://www.jstor.org/stable/20441281.

See correction [SL06a].

Sun:2009:COL

[SLL09] Fasheng Sun, Min-Qian Liu, and Dennis K. J. Lin. Construction of orthogonal Latin hypercube designs. Biometrika, 96(4):971–974, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/971;   http://www.jstor.org/stable/27798880.

Senturk:2005:CAR

[ŞM05] Damla Şentürk and Hans-Georg Müller. Covariate-adjusted regression. Biometrika, 92(1):75–89, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/92/1/75;   http://
www.jstor.org/stable/20441167.

Senturk:2008:GVC

[ŞM08] Damla Şentürk and Hans-Georg Müller. Generalized varying coefficient models for longitudinal data. Biometrika, 95(3):653–666, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/3/653;   http://www.jstor.org/stable/20441492.

Senturk:2009:CAG

[ŞM09] Damla Şentürk and Hans-Georg Müller. Covariate-adjusted generalized linear models. Biometrika, 96(2):357–370, June 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/2/357;
 http://www.jstor.org/stable/27798829.

Severini:2002:EPM

[SMG02] Thomas A. Severini, Rahul Mukerjee, and Malay Ghosh. On an exact probability matching property of right-invariant priors. Biometrika, 89(4):952–957, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/4/952;   http://
www.jstor.org/stable/4140551.

Smith:2001:BCS

[Smi01] T. M. F. Smith. Biometrika Centenary: Sample surveys. Biometrika, 88(1):167–194, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673678.

Sinha:2009:EBE

[SMLG09] Debajyoti Sinha, M. Brent Mchenry, Stuart R. Lipsitz, and Malay Ghosh. Empirical Bayes estimation for additive hazards regression models. Biometrika, 96(3):545–558, September 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/545;
 http://www.jstor.org/stable/27798847.

Seber:2000:RMM

[SN00] G. A. F. Seber and S. O. Nyangoma. Residuals for multinomial models. Biometrika, 87(1):183–191, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/183;
 http://www.jstor.org/stable/2673571.

Stefanski:2005:ENF

[SND05] Leonard A. Stefanski, Steven J. Novick, and Viswanath Devanarayan. Estimating a nonlinear function of a normal mean. Biometrika, 92(3):732–736, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/3/732;   http://
www.jstor.org/stable/20441227.

Schisterman:2001:EMS

[SR01a] Enrique Schisterman and Andrea Rotnitzky. Estimation of the mean of a K-sample U-statistic with missing outcomes and auxiliaries. Biometrika, 88(3):713–725, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673441.

Seaman:2001:BAC

[SR01b] Shaun R. Seaman and Sylvia Richardson. Bayesian analysis of casecontrol studies with categorical covariates. Biometrika, 88(4):1073– 1088, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/88/4/1073;   http://www.jstor.org/stable/2673702.

Scharfstein:2002:EFT

[SR02] Daniel O. Scharfstein and James M. Robins. Estimation of the failure time distribution in the presence of informative censoring. Biometrika, 89(3):617–634, August 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/3/617;
 http://www.jstor.org/stable/4140606.

Seaman:2004:EPR

[SR04] Shaun R. Seaman and Sylvia Richardson. Equivalence of prospective and retrospective models in the Bayesian analysis of casecontrol studies. Biometrika, 91(1):15–25, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/15;
 http://www.jstor.org/stable/20441076. See amendments

and correction [SR05].

Seaman:2005:ACE

[SR05] Shaun R. Seaman and Sylvia Richardson. Amendments and corrections: “Equivalence of prospective and retrospective models in the Bayesian analysis of case-control studies” [Biometrika 91 (2004), no. 1, 15–25; MR2050457]. Biometrika, 92(2):505, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/2/505.
2;   http://www.jstor.org/stable/20441206. See [SR04].

Sexton:2000:EAC

[SS00] Joe Sexton and Anders Rygh Swensen. ECM algorithms that converge at the rate of EM. Biometrika, 87(3):651–662, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/3/651;   http://www.jstor.org/stable/2673636.

Speckman:2003:FBS

[SS03] Paul L. Speckman and Dongchu Sun. Fully Bayesian spline smoothing and intrinsic autoregressive priors. Biometrika, 90(2):289–302, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/90/2/289;   http://www.jstor.org/stable/30042040.

Salgueiro:2005:PEE

[SSM05] M. Fátima Salgueiro, Peter W. F. Smith, and John W. Mc-Donald. Power of edge exclusion tests in graphical Gaussian models. Biometrika, 92(1):173–182, March 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/1/173;
 http://www.jstor.org/stable/20441174.

Schlather:2003:DMM

[ST03] Martin Schlather and Jonathan A. Tawn. A dependence measure for multivariate and spatial extreme values: Properties and inference. Biometrika, 90(1):139–156, March 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/1/139;
 http://www.jstor.org/stable/30042025.

Stephenson:2005:EOT

[ST05]  Alec Stephenson and Jonathan Tawn.  Exploiting occur-

rence times in likelihood inference for componentwise maxima. Biometrika, 92(1):213–227, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/1/213;   http://
www.jstor.org/stable/20441177.

Stern:2006:SAO

[Ste06] Steven E. Stern. Simple and accurate one-sided inference based on a class of M-estimators. Biometrika, 93(4):973–987, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/973;   http://www.jstor.org/stable/20441339.

Strawderman:2005:AGT

[Str05] Robert L. Strawderman. The accelerated gap times model. Biometrika, 92(3):647–666, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/3/647;   http://
www.jstor.org/stable/20441220.

Sundberg:2002:CRT

[Sun02] Rolf Sundberg. The convergence rate of the TM algorithm of Edwards & Lauritzen. Biometrika, 89(2):478–483, June 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/2/478;
 http://www.jstor.org/stable/4140593.

Sasieni:2003:MDR

[SW03] Peter D. Sasieni and Angela Winnett. Martingale difference residuals as a diagnostic tool for the Cox model. Biometrika, 90(4):899– 912, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/90/4/899;   http://www.jstor.org/stable/30042096.

Stanghellini:2005:IPA

[SW05] Elena Stanghellini and Nanny Wermuth. On the identification of path analysis models with one hidden variable. Biometrika, 92(2):337–350, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/2/337;   http://
www.jstor.org/stable/20441192.

Sweeting:2001:CPB

[Swe01] Trevor J. Sweeting. Coverage probability bias, objective Bayes and the likelihood principle. Biometrika, 88(3):657–675, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673437.

Sweeting:2005:ILP

[Swe05] Trevor J. Sweeting. On the implementation of local probability matching priors for interest parameters. Biometrika, 92(1):47–57, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/1/47;   http://www.jstor.org/stable/20441165.

Stute:2008:MDP

[SXZ08] W. Stute, W. L. Xu, and L. X. Zhu. Model diagnosis for parametric regression in high-dimensional spaces. Biometrika, 95(2):451–467, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/2/451;   http://www.jstor.org/stable/20441475.

Scheike:2008:PCI

[SZG08] Thomas H. Scheike, Mei-Jie Zhang, and Thomas A. Gerds. Predicting cumulative incidence probability by direct binomial regression. Biometrika, 95(1):205–220, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/205;
 http://www.jstor.org/stable/20441452.

Song:2009:NSE

[SZK09] Rui Song, Haibo Zhou, and Michael R. Kosorok. A note on semiparametric efficient inference for two-stage outcome-dependent sampling with a continuous outcome. Biometrika, 96(1):221–228, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/1/221;   http://www.jstor.org/stable/27798814.

Tang:2001:TCF

[Tan01] Boxin Tang. Theory of J-characteristics for fractional factorial designs and projection justification of minimum G_{2}-aberration. Biometrika, 88(2):401–407, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673488.

Tang:2006:OAR

[Tan06] Boxin Tang. Orthogonal arrays robust to nonnegligible twofactor interactions. Biometrika, 93(1):137–146, March 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/1/137;
 http://www.jstor.org/stable/20441266.

Tan:2009:NPL

[Tan09] Z. Tan. A note on profile likelihood for exponential tilt mixture models. Biometrika, 96(1):229–236, March 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/1/229;
 http://www.jstor.org/stable/27798815.

Tardella:2002:NBM

[Tar02] Luca Tardella. A new Bayesian method for nonparametric capturerecapture models in presence of heterogeneity. Biometrika, 89 (4):807–817, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/89/4/807;   http://www.jstor.
org/stable/4140538.

Tsai:2000:TNP

[TC00] Henghsiu Tsai and K. S. Chan. Testing for nonlinearity with partially observed time series. Biometrika, 87(4):805–821, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
87/4/805;   http://www.jstor.org/stable/2673611.

Titterington:2001:BOH

[TC01] D. M. Titterington and D. R. (David Roxbee) Cox, editors. Biometrika: one hundred years. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850993-6. viii + 383 pp. LCCN QH301 .T55 2001. URL http://www.loc.gov/

catdir/enhancements/fy0612/2001036127-d.html;   http:/
/www.loc.gov/catdir/toc/fy031/2001036127.html.

Tsai:2002:NTN

[TC02] Henghsiu Tsai and K. S. Chan. A note on testing for nonlinearity with partially observed time series. Biometrika, 89(1):245–250, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/1/245;   http://www.jstor.org/stable/4140572.

Tchetgen:2006:DTM

[TC06a] Eric J. Tchetgen and Brent A. Coull. A diagnostic test for the mixing distribution in a generalised linear mixed model. Biometrika, 93(4):1003–1010, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

biomet.oxfordjournals.org/content/93/4/1003;   http://
www.jstor.org/stable/20441342.

Tian:2006:AFT

[TC06b] Lu Tian and Tianxi Cai. On the accelerated failure time model for current status and interval censored data. Biometrika, 93(2):329– 342, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/93/2/329;   http://www.jstor.org/stable/20441285.

Tian:2007:MEB

[TCGW07] Lu Tian, Tianxi Cai, Els Goetghebeur, and L. J. Wei. Model evaluation based on the sampling distribution of estimated absolute prediction error. Biometrika, 94(2):297–311, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/297;
 http://www.jstor.org/stable/20441373.

Tsiatis:2001:SEP

[TD01] Anastasios A. Tsiatis and Marie Davidian. A semiparametric estimator for the proportional hazards model with longitudinal covariates measured with error. Biometrika, 88(2):447–458, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673492.

Tsiatis:2002:MIM

[TDM02] Anastasios A. Tsiatis, Marie Davidian, and Brad McNeney. Multiple imputation methods for testing treatment differences in survival distributions with missing cause of failure. Biometrika, 89(1): 238–244, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/89/1/238;   http://www.jstor.org/stable/4140571.

Terrell:2003:WHT

[Ter03] George R. Terrell. The Wilson-Hilferty transformation is locally saddlepoint. Biometrika, 90(2):445–453, June 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/2/445;
 http://www.jstor.org/stable/30042052.

Tille:2004:CCE

[TF04] Yves Tillé and Anne-Catherine Favre. Coordination, combination and extension of balanced samples. Biometrika, 91(4):913–927, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/4/913;   http://www.jstor.org/stable/20441152.

Thompson:2007:FVL

[TG07] Elizabeth A. Thompson and Charles J. Geyer. Fuzzy p-values in latent variable problems. Biometrika, 94(1):49–60, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/1/49;   http://www.jstor.org/stable/20441353.

Tsai:2000:PTL

[TGM00] Pi-Wen Tsai, Steven G. Gilmour, and Roger Mead. Projective three-level main effects designs robust to model uncertainty. Biometrika, 87(2):467–475, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/2/467;   http://
www.jstor.org/stable/2673476.

Taniguchi:2005:SJE

[TH05] Masanobu Taniguchi and Junichi Hirukawa. The Stein: James estimator for short- and long-memory Gaussian processes. Biometrika, 92(3):737–746, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/3/737;   http://
www.jstor.org/stable/20441228.

Tseng:2005:JMA

[THW05] Yi-Kuan Tseng, Fushing Hsieh, and Jane-Ling Wang. Joint modelling of accelerated failure time and longitudinal data. Biometrika, 92(3):587–603, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/92/3/587;   http://
www.jstor.org/stable/20441216.

Titterington:2001:E

[Tit01] D. M. Titterington. Editorial. Biometrika, 88(1):1, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2673672.

Tamhane:2004:SEA

[TL04] Ajit C. Tamhane and Brent R. Logan. A superiority-equivalence approach to one-sided tests on multiple endpoints in clinical trials. Biometrika, 91(3):715–727, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/20441133.

Tamhane:2008:WHP

[TL08] Ajit C. Tamhane and Lingyun Liu. On weighted Hochberg procedures. Biometrika, 95(2):279–294, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/279;   http://
www.jstor.org/stable/20441464.

Tang:2003:AMM

[TLR03] Gong Tang, Roderick J. A. Little, and Trivellore E. Raghunathan. Analysis of multivariate missing data with nonignorable nonresponse. Biometrika, 90(4):747–764, December 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/747;
 http://www.jstor.org/stable/30042086.

Tian:2004:SIB

[TLZW04] L. Tian, J. Liu, Y. Zhao, and L. J. Wei. Statistical inference based on non-smooth estimating functions. Biometrika, 91(4):943–954, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/91/4/943;   http://www.jstor.org/stable/20441154.

Tsiatis:2003:IAD

[TM03] Anastasios A. Tsiatis and Cyrus Mehta. On the inefficiency of the adaptive design for monitoring clinical trials. Biometrika, 90(2):367–378, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/367;   http://
www.jstor.org/stable/30042046.

Tsiatis:2004:LES

[TM04] Anastasios A. Tsiatis and Yanyuan Ma. Locally efficient semiparametric estimators for functional measurement error models. Biometrika, 91(4):835–848, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/4/835;   http://
www.jstor.org/stable/20441147.

Tang:2008:PCS

[TM08] Rong Tang and Hans-Georg Müller. Pairwise curve synchronization for functional data. Biometrika, 95(4):875–889, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/4/875;
 http://www.jstor.org/stable/20441512.

Tong:2001:PJT

[Ton01] Howell Tong. A personal journey through time series in Biometrika. Biometrika, 88(1):195–218, March 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673679.

Tsay:2000:OMT

[TPP00] Ruey S. Tsay, Daniel Peña, and Alan E. Pankratz. Outliers in multivariate time series. Biometrika, 87(4):789–804, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/789;
 http://www.jstor.org/stable/2673610.

Tebbs:2003:EOB

[TS03] Joshua M. Tebbs and William H. Swallow. Estimating ordered binomial proportions with the use of group testing. Biometrika, 90(2):471–477, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/471;   http://
www.jstor.org/stable/30042055.

Tsai:2009:PPL

[Tsa09] Wei Yann Tsai. Pseudo-partial likelihood for proportional hazards models with biased-sampling data. Biometrika, 96(3):601–615, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/3/601;   http://www.jstor.org/stable/27798851.

Tong:2005:ERV

[TW05] Tiejun Tong and Yuedong Wang. Estimating residual variance in nonparametric regression using least squares. Biometrika, 92 (4):821–830, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/4/821;   http://www.jstor.
org/stable/20441238.

Taylor:2007:MDS

[TWG07] J. E. Taylor, K. J. Worsley, and F. Gosselin. Maxima of discretely sampled random fields, with an application to ‘Bubbles’. Biometrika, 94(1):1–18, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/94/1/1;   http://www.jstor.
org/stable/20441350.

Ueki:2009:NAV

[Uek09] Masao Ueki. A note on automatic variable selection using smooththreshold estimating equations. Biometrika, 96(4):1005–1011, December 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/4/1005;   http://www.jstor.org/stable/27798885.

Ueki:2007:AEP

[UF07] Masao Ueki and Kaoru Fueda. Adjusting estimative prediction limits. Biometrika, 94(2):509–511, June 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/2/509;
 http://www.jstor.org/stable/20441389.

Uno:2005:OCR

[UTW05] Hajime Uno, Lu Tian, and L. J. Wei. The optimal confidence region for a random parameter. Biometrika, 92(4):957–964, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/4/957;   http://www.jstor.org/stable/20441248.

Vaida:2005:CAI

[VB05] Florin Vaida and Suzette Blanchard. Conditional Akaike information for mixed-effects models. Biometrika, 92(2):351–370, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
92/2/351;   http://www.jstor.org/stable/20441193.

vanderLaan:2000:NML

[vdLA00] Mark J. van der Laan and Chris Andrews. The nonparametric maximum likelihood estimator in a class of doubly censored current status data models with application to partner studies. Biometrika, 87(1):61–71, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/87/1/61;   http://
www.jstor.org/stable/2673561.

Verdinelli:2000:NBD

[Ver00] Isabella Verdinelli. A note on Bayesian design for the normal linear model with unknown error variance. Biometrika, 87(1):222–227, March 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/1/222;   http://www.jstor.org/stable/2673577.

Vicard:2000:ISF

[Vic00] Paola Vicard. On the identification of a single-factor model with correlated residuals. Biometrika, 87(1):199–205, March 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/1/199;
 http://www.jstor.org/stable/2673573.

Vidoni:2001:IPL

[Vid01] Paolo Vidoni. Improved prediction limits for continuous and discrete observations in generalised linear models. Biometrika, 88 (3):881–887, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673455.

vanLieshout:2001:CMB

[vLMZ01] M. N. M. van Lieshout, I. S. Molchanov, and S. A. Zuyev. Clustering methods based on variational analysis in the space of measures. Biometrika, 88(4):1021–1033, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
88/4/1021;   http://www.jstor.org/stable/2673699.

Vanderweele:2008:ECC

[VR08]  Tyler J. Vanderweele and James M. Robins.  Empirical

and counterfactual conditions for sufficient cause interactions. Biometrika, 95(1):49–61, March 2008. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/95/1/49;   http://
www.jstor.org/stable/20441442.

Vansteelandt:2007:ERM

[VRR07] Stijn Vansteelandt, Andrea Rotnitzky, and James Robins. Estimation of regression models for the mean of repeated outcomes under nonignorable nonmonotone nonresponse. Biometrika, 94(4):841– 860, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/4/841;   http://www.jstor.org/stable/20441421.

Varin:2005:NCL

[VV05] Cristiano Varin and Paolo Vidoni. A note on composite likelihood inference and model selection. Biometrika, 92(3):519–528, September 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/3/519;   http://www.jstor.org/stable/20441211.

Vardi:2001:TST

[VYZ01] Yehuda Vardi, Zhiliang Ying, and Cun-Hui Zhang. Twosample tests for growth curves under dependent right censoring. Biometrika, 88(4):949–960, December 2001. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/88/4/949;
 http://www.jstor.org/stable/2673694.

Waagepetersen:2008:EFI

[Waa08] Rasmus Waagepetersen. Estimating functions for inhomogeneous spatial point processes with incomplete covariate data. Biometrika, 95(2):351–363, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/2/351;   http://
www.jstor.org/stable/20441469.

Walden:2000:UVM

[Wal00] A. T. Walden. A unified view of multitaper multivariate spectral estimation. Biometrika, 87(4):767–788, December 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/767;
 http://www.jstor.org/stable/2673609.

Walker:2003:SCB

[Wal03] Stephen Walker. On sufficient conditions for Bayesian consistency. Biometrika, 90(2):482–488, June 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/2/482;   http://
www.jstor.org/stable/30042057.

Wang:2003:MNK

[Wan03] Naisyin Wang. Marginal nonparametric kernel regression accounting for within-subject correlation. Biometrika, 90(1):43–52, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
90/1/43;   http://www.jstor.org/stable/30042018.

Wang:2004:CAM

[Wan04] Yuchung J. Wang. Compatibility among marginal densities. Biometrika, 91(1):234–239, March 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/91/1/234;   http://
www.jstor.org/stable/20441092.

Wang:2009:WTG

[Wan09] Lan Wang. Wilcoxon-type generalized Bayesian information criterion. Biometrika, 96(1):163–173, March 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/1/163;   http://
www.jstor.org/stable/27798809.

Wang:2003:WCS

[WC03] You-Gan Wang and Vincent Carey. Working correlation structure misspecification, estimation and covariate design: Implications for generalised estimating equations performance. Biometrika, 90(1):29–41, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//biomet.oxfordjournals.org/content/90/1/29;   http://
www.jstor.org/stable/30042017.

Wermuth:2008:DEC

[WC08] Nanny Wermuth and D. R. Cox. Distortion of effects caused by indirect confounding. Biometrika, 95(1):17–33, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/1/17;   http://www.jstor.org/stable/20441440.

Wong:2003:EEC

[WCK03] Frederick Wong, Christopher K. Carter, and Robert Kohn. Efficient estimation of covariance selection models. Biometrika, 90 (4):809–830, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/90/4/809;   http://www.jstor.
org/stable/30042090.

Wong:2009:SMA

[WCK09] C. S. Wong, W. S. Chan, and P. L. Kam. A Student tmixture autoregressive model with applications to heavy-tailed financial data. Biometrika, 96(3):751–760, September 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/3/751;
 http://www.jstor.org/stable/27798861.

Wang:2007:NEA

[WCMR07] Yuanjia Wang, Lorraine N. Clark, Karen Marder, and Daniel Rabinowitzn. Nonparametric estimation of age-at-onset distributions from censored kin-cohort data. Biometrika, 94(2):403–414, June 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/2/403;   http://www.jstor.org/stable/20441380.

Walker:2000:RLP

[WD00a] Stephen Walker and Paul Damien. Representations of Lévy processes without Gaussian components. Biometrika, 87(2):477–483, June 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/2/477;   http://www.jstor.org/stable/2673477.

Wang:2000:AAB

[WD00b] Weijing Wang and A. Adam Ding. On assessing the association for bivariate current status data. Biometrika, 87(4):879–893, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/879;   http://www.jstor.org/stable/2673616.

Wang:2008:SMB

[WD08] Qihua Wang and Pengjie Dai. Semiparametric model-based inference in the presence of missing responses. Biometrika, 95(3):721– 734, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/95/3/721;   http://www.jstor.org/stable/20441497.

Westfall:2008:BHM

[Wes08] Peter H. Westfall. The Benjamini-Hochberg method with infinitely many contrasts in linear models. Biometrika, 95(3):709–719, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/95/3/709;   http://www.jstor.org/stable/20441496.

Wang:2003:APE

[WH03] Molin Wang and John J. Hanfelt. Adjusted profile estimating function. Biometrika, 90(4):845–858, December 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/845;
 http://www.jstor.org/stable/30042092.

Whittemore:2004:EGA

[Whi04] Alice S. Whittemore. Estimating genetic association parameters from family data. Biometrika, 91(1):219–225, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/219;
 http://www.jstor.org/stable/20441090.

Wang:2008:CSC

[WI08] Yuchung J. Wang and Edward H. Ip. Conditionally specified continuous distributions. Biometrika, 95(3):735–746, September 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/735;
 http://www.jstor.org/stable/20441498.

Williams:2000:UAE

[WJ00] E. R. Williams and J. A. John. Updating the average efficiency factor in α-designs. Biometrika, 87(3):695–699, September 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/3/695;
 http://www.jstor.org/stable/2673640.

Wood:2002:BMS

[WJT02] Sally A. Wood, Wenxin Jiang, and Martin Tanner. Bayesian mixture of splines for spatially adaptive nonparametric regression. Biometrika, 89(3):513–528, August 2002. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/3/513;   http://
www.jstor.org/stable/4140598.

Wong:2001:LMA

[WL01] C. S. Wong and W. K. Li. On a logistic mixture autoregressive model. Biometrika, 88(3):833–846, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2673450.

Wang:2003:MCM

[WL03] Zengri Wang and Thomas A. Louis. Matching conditional and marginal shapes in binary random intercept models using a bridge distribution function. Biometrika, 90(4):765–775, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/4/765;
 http://www.jstor.org/stable/30042087.

Walker:2005:DTU

[WLP05] Stephen G. Walker, Antonio Lijoi, and Igor Prünster. Data tracking and the understanding of Bayesian consistency. Biometrika, 92(4):765–778, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/4/765;   http://www.jstor.
org/stable/20441234.

Wang:2006:FPS

[WLS06] Xinlei Wang, Johan Lim, and S. Lynne Stokes. Forming post-strata via Bayesian treed capture-recapture models. Biometrika, 93(4):861–876, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/4/861;   http://
www.jstor.org/stable/20441332.

Wang:2007:TPS

[WLT07] Hansheng Wang, Runze Li, and Chih-Ling Tsai. Tuning parameter selectors for the smoothly clipped absolute deviation method. Biometrika, 94(3):553–568, August 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/553;
 http://www.jstor.org/stable/20441396.

Wang:2009:HPC

[WNZZ09] S. Wang, B. Nan, N. Zhou, and J. Zhu. Hierarchically penalized Cox regression with grouped variables. Biometrika, 96(2):307–322, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/96/2/307;   http://www.jstor.org/stable/27798826.

Watanabe:2004:MMS

[WO04] Toshiaki Watanabe and Yasuhiro Omori. A multi-move sampler for estimating non-Gaussian time series models: Comments on Shephard & Pitt (1997) [“Likelihood analysis of non-Gaussian measurement time series” [Biometrika 84 (1997), no. 3, 653– 667; mr1603940]]. Biometrika, 91(1):246–248, March 2004. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/91/1/246;
 http://www.jstor.org/stable/20441094.

Wu:2003:NEL

[WP03] Wei Biao Wu and Mohsen Pourahmadi. Nonparametric estimation of large covariance matrices of longitudinal data. Biometrika, 90(4):831–844, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/831;   http://
www.jstor.org/stable/30042091.

Wang:2002:ELB

[WR02] Qihua Wang and J. N. K. Rao. Empirical likelihood-based inference in linear errors-in-covariables models with validation data. Biometrika, 89(2):345–358, June 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/89/2/345;   http://
www.jstor.org/stable/4140581.

Winnett:2001:NSS

[WS01] Angela Winnett and Peter Sasieni. A note on scaled Schoenfeld residuals for the proportional hazards model. Biometrika, 88(2):565–571, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2673500.

Wu:2009:ARU

[WSG09] S. Wu, X. Shen, and C. J. Geyer. Adaptive regularization using the entire solution surface. Biometrika, 96(3):513–527, September 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
96/3/513;   http://www.jstor.org/stable/27798845.

Wang:2008:PEL

[WSL08] Junhui Wang, Xiaotong Shen, and Yufeng Liu. Probability estimation for large-margin classifiers. Biometrika, 95(1):149–167, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/1/149;   http://www.jstor.org/stable/20441449.

Wahed:2006:SEE

[WT06] Abdus S. Wahed and Anastasios A. Tsiatis. Semiparametric efficient estimation of survival distributions in two-stage randomisation designs in clinical trials with censored data. Biometrika, 93(1):163–177, March 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/1/163;   http://
www.jstor.org/stable/20441268.

Wu:2003:OCE

[Wu03] Changbao Wu. Optimal calibration estimators in survey sampling. Biometrika, 90(4):937–951, December 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/90/4/937;   http://
www.jstor.org/stable/30042099.

Wang:2006:EBB

[WW06] Xue Wang and Andrew T. A. Wood. Empirical Bayes block shrinkage of wavelet coefficients via the noncentral χ^{2} distribution. Biometrika, 93(3):705–722, September 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/3/705;
 http://www.jstor.org/stable/20441317.

Wang:2009:BAM

[WW09] Hao Wang and Mike West. Bayesian analysis of matrix normal graphical models. Biometrika, 96(4):821–834, December 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/4/821;
 http://www.jstor.org/stable/27798869.

Wu:2001:IRA

[WWM01] Wei Biao Wu, Michael Woodroofe, and Graciela Mentz. Isotonic regression: Another look at the changepoint problem. Biometrika, 88(3):793–804, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673447.

Wang:2006:RBR

[WZ06] You-Gan Wang and Min Zhu. Rank-based regression for analysis of repeated measures. Biometrika, 93(2):459–464, June 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/459;
 http://www.jstor.org/stable/20441295.

Xia:2009:MCR

[Xia09] Yingcun Xia. Model checking in regression via dimension reduction. Biometrika, 96(1):133–148, March 2009. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/96/1/133;
 http://www.jstor.org/stable/27798807.

Xiang:2003:RVE

[XL03] Anny Hui Xiang and Bryan Langholz. Robust variance estimation for rate ratio parameter estimates from individually matched casecontrol data. Biometrika, 90(3):741–746, September 2003. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/3/741;
 http://www.jstor.org/stable/30042081.

Xu:2005:SND

[Xu05] Hongquan Xu. Some nonregular designs from the Nordstrom-Robinson code and their statistical properties. Biometrika, 92(2): 385–397, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/2/385;   http://www.jstor.org/stable/20441195.

Xue:2007:ELS

[XZ07] Liugen Xue and Lixing Zhu. Empirical likelihood semiparametric regression analysis for longitudinal data. Biometrika, 94(4):921– 937, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/94/4/921;   http://www.jstor.org/stable/20441426.

Xia:2004:EES

[XZT04] Yingcun Xia, Wenyang Zhang, and Howell Tong. Efficient estimation for semivarying-coefficient models. Biometrika, 91(3): 661–681, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/20441130.

Yang:2005:CSA

[Yan05] Yuhong Yang. Can the strengths of AIC and BIC be shared? A conflict between model indentification and regression estimation. Biometrika, 92(4):937–950, December 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/4/937;
 http://www.jstor.org/stable/20441246.

Yin:2003:ECS

[YC03] Xiangrong Yin and R. Dennis Cook. Estimating central subspaces via inverse third moments. Biometrika, 90(1):113–125, March 2003. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/90/1/113;
 http://www.jstor.org/stable/30042023.

Yin:2004:AHM

[YC04] Guosheng Yin and Jianwen Cai. Additive hazards model with multivariate failure time data. Biometrika, 91(4):801–818, December 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
91/4/801;   http://www.jstor.org/stable/20441145.

Yin:2005:DES

[YC05] Xiangrong Yin and R. Dennis Cook. Direction estimation in singleindex regressions. Biometrika, 92(2):371–384, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/2/371;
 http://www.jstor.org/stable/20441194.

Yoo:2007:OSD

[YC07] Jae Keun Yoo and R. Dennis Cook. Optimal sufficient dimension reduction for the conditional mean in multivariate regression. Biometrika, 94(1):231–242, March 2007. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/1/231;
 http://www.jstor.org/stable/20441365.

Yeo:2000:NFP

[YJ00] In-Kwon Yeo and Richard A. Johnson. A new family of power transformations to improve normality or symmetry. Biometrika, 87(4):954–959, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/87/4/954;   http://
www.jstor.org/stable/2673623.

Yee:2002:AAP

[YJS02] Julie L. Yee, Wesley O. Johnson, and Francisco J. Samaniego. Asymptotic approximations to posterior distributions via conditional moment equations. Biometrika, 89(4):755–767, December 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/4/755;   http://www.jstor.org/stable/4140535.

Yuan:2007:MSE

[YL07] Ming Yuan and Yi Lin. Model selection and estimation in the Gaussian graphical model. Biometrika, 94(1):19–35, March 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
94/1/19;   http://www.jstor.org/stable/20441351.

Ybarra:2008:SAE

[YL08a] Lynn M. R. Ybarra and Sharon L. Lohr. Small area estimation when auxiliary information is measured with error. Biometrika, 95(4):919–931, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/919;   http://
www.jstor.org/stable/20441515.

Yu:2008:NRU

[YL08b] Zhangsheng Yu and Xihong Lin. Nonparametric regression using local kernel estimating equations for correlated failure time data. Biometrika, 95(1):123–137, March 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/1/123;
 http://www.jstor.org/stable/20441447.

Yun:2007:UHL

[YLK07] Sung-Cheol Yun, Youngjo Lee, and Michael G. Kenward. Using hierarchical likelihood for missing data problems. Biometrika, 94(4):905–919, December 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/94/4/905;   http://www.jstor.
org/stable/20441425.

Yang:2005:SAS

[YP05] Song Yang and Ross Prentice. Semiparametric analysis of shortterm and long-term hazard ratios with two-sample survival data. Biometrika, 92(1):1–17, March 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/92/1/1;   http://www.jstor.
org/stable/20441162.

Ye:2006:MCS

[YP06] Huajun Ye and Jianxin Pan. Modelling of covariance structures in generalised estimating equations for longitudinal data. Biometrika, 93(4):927–941, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/4/927;   http://
www.jstor.org/stable/20441336.

Yuen:2006:STI

[YSZ06] Kam-Chuen Yuen, Jian Shi, and Lixing Zhu. A k-sample test with interval censored data. Biometrika, 93(2):315–328, June 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/93/2/315;
 http://www.jstor.org/stable/20441284.

Yang:2001:MPH

[YY01] Yaning Yang and Zhiliang Ying. Marginal proportional hazards models for multiple event-time data. Biometrika, 88(2):581–586, June 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2673502.

Zeng:2009:GFT

[ZCI09] Donglin Zeng, Qingxia Chen, and Joseph G. Ibrahim. Gamma frailty transformation models for multivariate survival times. Biometrika, 96(2):277–291, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/2/277;   http://
www.jstor.org/stable/27798824.

Zeng:2008:DDC

[Zen08] Peng Zeng. Determining the dimension of the central subspace and central mean subspace. Biometrika, 95(2):469–479, June 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/469;
 http://www.jstor.org/stable/20441476.

Zou:2002:NPE

[ZF02] F. Zou and J. P. Fine. A note on a partial empirical likelihood. Biometrika, 89(4):958–961, December 2002. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/89/4/958;
 http://www.jstor.org/stable/4140552.

Zhu:2008:AMR

[ZFH08] Zhongyi Zhu, Wing K. Fung, and Xuming He. On the asymptotics of marginal regression splines with longitudinal data. Biometrika, 95(4):907–917, December 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/95/4/907;   http://
www.jstor.org/stable/20441514.

Zou:2002:ELS

[ZFY02] F. Zou, J. P. Fine, and B. S. Yandell. On empirical likelihood for a semiparametric mixture model. Biometrika, 89(1):61–75, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/89/1/61;   http://www.jstor.org/stable/4140559.

Zhang:2001:IMT

[Zha01] Biao Zhang. An information matrix test for logistic regression models based on case-control data. Biometrika, 88(4):921–932, December 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/88/4/921;   http://www.jstor.org/stable/2673692.

Zhang:2002:SPE

[Zha02] Ying Zhang. A semiparametric pseudolikelihood estimation method for panel count data. Biometrika, 89(1):39–48, March 2002. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
89/1/39;   http://www.jstor.org/stable/4140557.

Zhang:2006:NST

[Zha06] Ying Zhang. Nonparametric k-sample tests with panel count data. Biometrika, 93(4):777–790, December 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/93/4/777;   http://
www.jstor.org/stable/20441327.

Zhou:2008:JMP

[ZHC08] Lan Zhou, Jianhua Z. Huang, and Raymond J. Carroll. Joint modelling of paired sparse functional data using principal components. Biometrika, 95(3):601–619, September 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/3/601;
 http://www.jstor.org/stable/20441489.

Zhao:2009:MPT

[ZHC09] Luping Zhao, Timothy E. Hanson, and Bradley P. Carlin. Mixtures of Polya trees for flexible spatial frailty survival modelling. Biometrika, 96(2):263–276, June 2009. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:/

/biomet.oxfordjournals.org/content/96/2/263;   http://
www.jstor.org/stable/27798823.

Zhou:2005:ELA

[Zho05] Mai Zhou. Empirical likelihood analysis of the rank estimator for the censored accelerated failure time model. Biometrika, 92(2):492– 498, June 2005. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/92/2/492;   http://www.jstor.org/stable/20441203.

Zhu:2006:DAC

[Zhu06] Mu Zhu. Discriminant analysis with common principal components. Biometrika, 93(4):1018–1024, December 2006. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
93/4/1018;   http://www.jstor.org/stable/20441344.

Zhu:2008:DME

[ZITZ08] Hongtu Zhu, Joseph G. Ibrahim, Niansheng Tang, and Heping Zhang. Diagnostic measures for empirical likelihood of general estimating equations. Biometrika, 95(2):489–507, June 2008. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/95/2/489;
 http://www.jstor.org/stable/20441478.

Zhou:2005:ELB

[ZL05] Yong Zhou and Hua Liang. Empirical-likelihood-based semiparametric inference for the treatment effect in the two-sample problem with censoring. Biometrika, 92(2):271–282, June 2005. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/92/2/271;
 http://www.jstor.org/stable/20441188.

Zeng:2006:EES

[ZL06] Donglin Zeng and D. Y. Lin. Efficient estimation of semiparametric transformation models for counting processes. Biometrika, 93(3):627–640, September 2006. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://biomet.

oxfordjournals.org/content/93/3/627;   http://www.jstor.
org/stable/20441312.

Zhang:2007:ALC

[ZL07] Hao Helen Zhang and Wenbin Lu. Adaptive Lasso for Cox’s proportional hazards model. Biometrika, 94(3):691–703, August 2007. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/94/3/691;
 http://www.jstor.org/stable/20441405.

Zhu:2001:CDM

[ZLWZ01] Hongtu Zhu, Sik-Yum Lee, Bo-Cheng Wei, and Julie Zhou. Casedeletion measures for models with incomplete data. Biometrika, 88(3):727–737, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2673442.

Zhang:2001:NTS

[ZLZ01] Ying Zhang, Wei Liu, and Yihui Zhan. A nonparametric twosample test of the failure function with interval censoring case 2. Biometrika, 88(3):677–686, September 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2673438.

Zhu:2000:NMC

[ZN00] Li-Xing Zhu and Georg Neuhaus. Nonparametric Monte Carlo tests for multivariate distributions. Biometrika, 87(4):919–928, December 2000. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://biomet.oxfordjournals.org/
content/87/4/919;   http://www.jstor.org/stable/2673619.

Zou:2008:NPB

[Zou08] Hui Zou. A note on path-based variable selection in the penalized proportional hazards model. Biometrika, 95(1):241–247, March 2008. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://biomet.oxfordjournals.org/content/
95/1/241;   http://www.jstor.org/stable/20441455.

Zhong:2000:ELI

[ZR00] Bob Zhong and J. N. K. Rao. Empirical likelihood inference under stratified random sampling using auxiliary population information. Biometrika, 87(4):929–938, December 2000. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://biomet.oxfordjournals.org/content/87/4/929;
 http://www.jstor.org/stable/2673620.

Zhu:2004:DPB

[ZZ04] Hongtu Zhu and Heping Zhang. A diagnostic procedure based on local influence. Biometrika, 91(3):579–589, September 2004. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/20441124.

Zhang:2005:TRT

[ZZ05] Hao Zhang and Dale L. Zimmerman. Towards reconciling two asymptotic frameworks in spatial statistics. Biometrika, 92(4):921– 936, December 2005. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://biomet.oxfordjournals.
org/content/92/4/921;   http://www.jstor.org/stable/20441245.